A Complete Bibliography of Publications in Journal of Computational and Applied Mathematics (1980–1989)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

04 March 2019 Version 1.02

Title word cross-reference

(A − λM)x = 0 [NNOP86]. (n, p) [Aga88]. (p, n) [Aga88]. $12.95 [Bul88c]. 2 [Han88]. $223.75 [Bul88b]. $29.50 [Rij80]. 3 [SvdV89]. $34.5 [Bul88c]. 4k − 3 [CH87b]. $58.00/Dfl [Van89a]. $79.50 [Ano88f]. $84.25 [Goo88b]. $95.25 [Ano88f]. _{2} [Gre81b]. A [Con87, vdV87]. A_{0} [Tho82a]. a_{n} → 0 [Jac88]. α [Roc88b]. ≈ [Ano89f]. B [Sho89]. C [JW89]. C^{k} [Sab85]. C^{S}[a, b] [Köh89]. Cl_{2}(Θ) [Gro84a, dD84]. d [Kle87, Pre87]. ϵ [Pre87, Tem88]. F [Bic82]. f(A)x = b [vdV87]. f(z) [JW89]. F [α, β] [Bic82]. F_{n}(z) = F_{n−1}(f_{n}(z)) [Gil88]. f_{n}(z) → f(z) [Gil88]. H^{1} [Bia89]. Ĵ [Ran87]. ∫_{0}^{1} x^{ν−1}(1 − x)^{−λ} ln^{m} x dx [Köl87].

∫_{0}^{∞} χ^{v−1}(1 + βχ)^{−λ} ln^{m} χ dχ [Köl86b]. ∫_{0}^{∞} χ^{v−1}(1 + βχ)^{−λ} ln^{m} χ dχ [Köl86a]. J_{ν}(x) [BGV81]. k [BLS84c, Bre89, Lee86]. K(a_{n}1) [Jac88]. K(z_{n}/1) [Waa87b]. K_{n}(z) [DC81]. L^{∞} [Mil89]. L_{2} [Bar83]. l_{p} [Zie84]. LDL^{T} [Han87]. ln^{2}(1 + x) [SK89a, SK89b]. M [Lip81, HO88]. N [GM80a, Heb82]. P [Ana86, CN86, CRN86, Col88, JKG83, Lip81, AEH81]. q [Pre87]. QD [Van87a, Van89c]. QR [Gra86]. r [Bre89, Pre87]. R^{3} [Cad87]. S [SK89a, SK89b, CG89, Pre87]. σ [Gro84b]. T_{+M} [Lem87]. τ [Col87a, Da 87]. × [Rij80]. u_{xx} + (2νx)u_{x} + ϵ^{2}u_{tt} = u_{t} [GL84]. x [Sab83]. y′′ = f(t, y) [Cha86, CR87, CRN86]. y′′ = f(x, y) [FG86]. y′′ = f(x, y, y′) [Ano84f, CS84a, CS84b, KJG83].

-acceptability [Con87]. -algorithm [Tem88, Van87a, Van89c]. -algorithms [Pre87]. -approximation [Zie84]. -convergence [Sho89]. -dimensional [GM80a]. -direction [Lee86]. -error [Mil89]. -fraction [JW89, SK89a, SK89b]. -fractions [Ran87]. -method [Col87a, Da 87]. -norm [Han88, Bia89]. -parametric [AEH81]. -space [Kle87]. -stable [DM82, Ana86, Bic82, CN86, CRN86, Col88, JKG83, Tho82a]. -step [CG89, HO88]. -sums [Gro84b]. -th [BLS84c, Bre89, Heb82]. -type [Roc88b].

0 [Ano88f, Bul88b, Goo88b, Lev88, Van89a]. 0-444-70209-1 [Bul88b]. 0-444-70340-3 [Van89a]. 0-444-70375-6 [Goo88b]. 0-521-33502-7 [Bul88c]. 0-521-33692-9 [Bul88c]. 0-8247-7811-1 [Ano88f]. 0-85274-418-8 [Lev88]. 006 [GM80a].

1 [Ano88f, Bul88b]. 135 [Ano84e]. 1s [Van87b].

2 [BR89]. 205 [NNOP86]. 224 [Ano89k]. 25 [Ano89k].

3 [Bul88a, Van89a]. 3-527-26463-9 [Bul88a].

5 [Ano89f].

6 [Goo88b].

7 [Goo88a].

8 [Ano82, Duf89, Lev88].

9 [Ano84f, Ano84e, Bul88a]. 90 [Ano84f, Goo88a]. 90-6196-292-7 [Goo88a]. 90s [Cas88]. 975 [Ano89f]. 975-429-007-5 [Ano89f].

Abel [CM84, Cap88]. Abel-type [Cap88].

Absolute [CS84a, CS84b, LW88]. absolute-value [LW88]. absorbing [GGD81]. abstract [Van89a]. Academic [Rij80]. Accelerated [LM86b, Mar86, NIS82, aON88, Yey88, Yey89]. Acceleration [GK89, Lem87, Bre85, Dic89, EL88, Jac88, JY88, LB88, Not88, Sab87, Pre87]. accelerometer [Mur87]. acceptability [Con87]. accuracy [Jel85]. accurate [Ana86]. acoustic [Ami87]. Adam [Lev88]. Adams [BGS88, Kho86, Tho82a, Tho82b]. Adams-type [Kho86, Tho82a, Tho82b]. adapting [vdHS86]. Adaptive [Kri82, Ber89a, GM80a, JY88, KR87, SS89a, YaOIN85]. ADI [BH83, De 84, Dyk86, Mam89, vdHdV83]. adjacency [Maa85]. adjoint [Hen87, Lip83a]. adjusting [WO80b]. adjustment [Mur87]. Adomian [Nel88]. advances [Bru82, BOP84, Zla87]. Advantages [HM85]. agglomeration [Spo89]. aggregate [Kiw86]. aggregation [Bla88]. aid [Rou87]. airfoil [Ioa88]. algebra [Dra87, Dra88, HH89, Kos88]. algebraic [BL87a, Gre86a, HC85, Say88]. Algorithm [PB84, AL87, CvD89, Cla81, Cos88, DI89, DL82, Die86, DSV88, Dra87, DKW84, EKL89, FL85, GM80a, Gra86, GM80b, Hof89, HS82, Jab80, JH89, KP82, Kip89, KKK89, LQ86, Lip81, LM84, McC81a, McC81b, Oli80a, PA89, Roo85, Ste85, Tem88, VVH87, VV88, Van89b, Van87a, Van89c, Von88, WKGC87, GM80a]. algorithmic [Lip83a, Sar88, Rij80]. Algorithms [Smy85, Ano83c, BZ89, BS89, Bul80, Dah86, EL89b, Fuk84, HKD86, KTS88, Mac82, MS88, Niz80, Niz84, Pre87, YaOIN85, vdC82, WO80a]. Alliant [Duf89]. allowed [KG87a]. alpha [NSIO82]. alternating [GK89]. Alternative [Hal82]. among [Del87]. amplitude [Gol86]. Amsterdam [Ano86f, Bul88b, Goo88b, Goo88a, Van89a]. analyses [SB88a]. analysis [BW89, Col80, Dic82a, ELS88, FW87, Gwi81, HC85, Kam83, Nii86, Niz80, Rou87, Sar88, Tt89, Twi86, Ano89f]. analytic [EL89a, Eie84, Gil88, Ioa89, JNT83, Kli87a, Kli87b, Let81, Pri80, Son89]. Analytical [CH88, Gre87, GG88, How84]. Announcement [Ano89a]. Announcements [Ano87a, Ano88a]. answer [Axe85]. antennas [SGC85, SG88]. AOR [Yey88, Yey89, KJ88, Mar88]. Appl [Ano82, Ano84f, Ano84e]. application [Aga83, Bog86, DN87, DN88, GLSZ87, Ioa84, Ioa86b, KL87, Nev85, Oli80b, Rei86, Von88]. Applications [EL89b, Son89, Ano82, BDvdM82, Bia89, Gau85, GV87, Goo88a, Ham82, ITH88, Pap89, Pap82, Tt89, Usm80, Waa87a, WO80a, dLM89, Ano86e]. Applied [Ano88m, Ano89k, AHCW86, ACCK88, Ano86j, Ano87n, Ano87o, Ano88f, Ano88n, Ano89o, Ano89p, Cad87, Mie89, Nii86, Puu88a, Sar88, Sny87]. approach [ÁKS88, FK88, GL87a, Gre81a, Gre81b, Ioa81, Kli87a, Sar88, WG81]. approximants [Con87, Cuy86, De 88, Eie84, GV87, GM81, GMS88, Hop82, JM80, Luk82, Mag87, McC81b, Pre83, Sab83, Sid80, Sid81, Wal87]. Approximate [eMH89, AMHAEM89, BTB88, Aga83, BW89, Hem82, Kam83, Lip83a, Ong84a]. approximating [Gre86a, Let87]. Approximation [BGV81, Dun88, Abd81, AHCW86, Bak84, BGN81, Col87a, Cro87, Cuy85, DM84a, Dun80, Dun82, Dun84a, Dun84b, Fuk84, GS86, HO86, IS87, LM82, LM84, Oli80c, Pap82, Pas87, Pri80, Roh87, Smi87a, Soh89, Son89, Van88, Zie84, Ano86d]. approximations [BL87a, Bia89, BZ84, Col87b, GL83, Ioa88, JH89, Luk80, Nak87, PW81b, Usm80, Waa87b, ZH80]. aquifer [SOA82]. arbitrarily [Fab87, Far86]. arbitrary [Cos88, Gre86b]. architectures [BE89]. arcs [Ioa89]. area [BP86]. argument [Sha88a]. arising [Dix87, Gd80, Gla84, Had85, Hüb86b, NR85, SGC85, SG88]. arithmetic [VS88]. Arlington [Bul88b]. ARMA [Bul84a]. Artificial [Goo88b]. Aspects [Whe85, BE89, CSR87, Fre84, Lon84, Sid80]. assignment [Niz84, Ren88]. associated [Gro86, Mur87, Not88, VVH87, Wam80]. Asymptotic [Cam87, Che84, Ano84e, Chi81, GL87a, McC83, Sal83, Van85]. asynchronous [HM85]. atom [Gre86b]. atomic [Van87b]. atoms [Gre81a, Gre81b, Gre87]. attention [DW89]. Author [Ano83a, Ano84a, Ano84b, Ano85a, Ano86a, Ano86b, Ano86c, Ano87e, Ano87b, Ano87c, Ano87d, Ano88d, Ano88e, Ano88b, Ano88c, Ano89b, Ano89c, Ano89d, Ano89e]. autocatalytic [KP89]. Automatic [CH87a, Fis89, GLG85, GSB87, ITH88, RD81, Cal85, Chu83]. Automation [For86a]. auxiliary [De 85]. average [Gre86a]. axisymmetric [Gol86].

B [Ano86d, Ano89f, Goo88b, Bul88c, Neu81, Soh88]. B-splines [Neu81, Soh88]. Banach [Cro87, Yam88]. banded [Pap82]. bandwidth [Smy85]. bandwidths [Lip81]. Bareiss [DI89]. barrier [GGD81, Jel86]. based [ALE84, Cha87, Dic85, Gut86b, Hel82, JV89, NF84, Rab87b, Rab89, Tem88]. Basel [Ano88f]. basic [Fre84, PF84]. basis [EM86, Wan85]. BDF [BGS88]. beamforming [SGC85, SG88]. behavior [MS88]. Benders [BB83]. Bernstein [Ano89k, GN89]. Bessel [Ano84e, BGV81, DC81, Gab80, GL83, ISK87, IS88, Laf86, PB84, Sal83, Tem81]. best [Köh89, LD86]. Between [dB80, AN81, DRV89, GGD81, Gol86, Gro84b, Mur87, RS86]. Bézier [LY89]. Bi [IS87, IN87, KP89]. Bi-orthogonality [IS87, IN87]. Bi-periodicity [KP89]. bibliographies [Ano83d, Gar86b]. Bibliography [Gwi81, Cuy87b, GS80]. bidiagonal [SvdV89]. Bifurcation [MMSY84, Man84b, Cal85, DRV89, DW89, Her83, Roo85, Smi87a, Smi87b]. bifurcations [Chi89, LST89, SCJ89]. biharmonic [RV88]. Biomolecular [ELS88]. Bivariate [dBH83, Lee86]. blending [Cad87]. Block [AP88, DSH89, Axe85, Bul84b, Bul84c, SvdV89]. board [Ano83b, Ano86h, Ano87f, Ano88i, Ano88j, Ano89g, Ano84c, Ano84d, Ano86g, Ano86i, Ano87g, Ano87h, Ano87i, Ano88h, Ano89h, Ano89i, Ano89j]. Böhmer [Ano86e]. Boltzmann [Léc89]. Bolyai [Bul88b]. Book [Ano86d, Ano86e, Ano86f, Ano88f, Ano89f, Bul88a, Bul88c, Bul88b, Goo88b, Goo88a, Lev88, Rij80, Van89a]. bordered [Men87]. bottom [BeMM86]. bound [DKW84, Maa84]. Boundary [AA82, IS85, Aga83, AC84, AU86, Aga87, Aga88, Ami87, AT80, Bia89, CK80, CK82, CS86, CS87, Cha87, CSS88, CS88, De 85, DR87c, ES85, EO82, FVD87, Ioa89, KR86, LM82, Lav84, Lin86, Man84a, MG89, Mu89, Mur82, NR85, Niz80, Özi89, PW81a, PD88, Rei84, RC86, RD87, Ruo87, RS89, Sha88a, Slo86, Sny87, SS88, Tt89, TW86, TT87, Usm80, WO80a, WO80b, ZN84, dG81]. boundary-value [Cha87, CSS88, CS88, LM82, Lav84, Man84a, Sha88a, SS88]. bounded [EHL89, SO87]. boundedness [Gar86a]. Bounds [BLS84a, BLS84c, BLS84b, IS88, Kio86, BF89, DG82, EL88, ISK87, JH89, KG87a, Yam85]. box [CDR89, Lee86, dBH83, Col80]. branch [DKW84]. Branched [Sie80, Cuy87a, CV88, Sie83]. breaking [Bud89]. brief [KY88]. Bristol [Lev88]. buckling [Chi89]. Budapest [Bul88b]. Burry [Ano86d].

C [Rij80]. cable [DM84b]. calculating [Bla87]. Calculation [Pre87, Cal85, Cal86, Cuy86, Gd80, Gro88, HS82, SCJ89]. calculations [PMD87]. calibration [Bog86]. Call [Ano88g]. Callocation [PJ80].

CAM [Ano83c, Ano83d]. Cambridge [Bul88c]. Canada [Ano88f]. Canonical [EL89b]. capacity [Niz84]. Carlo [BC82, Jab80, Lam85]. case [Roc88a]. cases [Zem86]. Catastrophe [Puu88a]. Cauchy [Dic85, Ger86, Hel82, Ioa81, Ioa86b, Ors89, dB80]. Celebi [Ano89f]. cell [Dix87]. census [Von88]. center [Mur82, Son89]. centered [Rok87]. centralization [AZ88]. Centre [Goo88a]. certain [BLS84a, CK87, GL87a, GK86, GV87, GG85, IMT87, Rei86, Tho82a, WP86]. CF [GS86]. CG [YaOIN85]. channel [Wat81]. Characterisation [Col88]. characteristics [PC80]. characterizations [De 89]. characterizing [Fis89]. charged [Fab87]. Chebyshev [Abd81, AEH81, DL82, Dun80, Dun82, Dun84b, HT87, JY88, Kar88a, Kar88b, PR88, SL86, Zie84]. choice [DKW84, NNOP86]. Cholesky [Han87]. circle [Die86]. circuit [AHCW86]. circular [CH87a, Rok87]. class [Ana86, CK80, CS86, Cha86, CS87, Cle81, Col88, FL85, Fuk84, Pap82, Ran87, Sab85, SGC85, SG88, Sug87a, Tt89, WP86]. Classical [ET86, Mu89, Van89b]. Classification [DN87]. Clausen [Gro84a, dD84]. clear [Whe85]. closed [Rab87b, Weg88]. cluster [Rou87]. clustering [Miy89]. cm [Rij80]. Co [RM89]. Co-recursive [RM89]. coal [Spo89]. coaxial [Fab87]. cocircuits [Ano82, Ham82]. code [Duf89]. codes [BGS88, Sha87]. coefficients [BB88, Che84, Heb82, Hop82, Jan87, Kar88a, Kar88b, Luk80, Luk82, Ste87]. coinciding [MPN82]. Collocation [Joe85, Bel84, Rei84, RC86, RS89]. Colloquia [Bul88b]. colored [CC86]. Combined [MP88]. Comment [Ast82]. Comments [Ado89, McN86b]. common [KLT89]. commutative [Dra87, Dra88]. comparative [De 84]. Comparison [McN86a, DM83, EM86, KB88, Mac82, McN88, Roo85]. complements [Tem88]. Complete [BDvdM82]. Completeness [Kli87b]. Complex [Col87a, AL86, BGG88, Col87b, JT85, Mag87, Pet82, PS86, Pri80, Rok87]. Complexity [Kos88, Ros88]. Composite [Sab85, Car89, Mas87, PF84]. Compositions [Gil88]. compressible [Van80]. Comput [Ano82, Ano84e]. Computation [GMM89, GR89b, JM80, Ste87, CH87a, DI89, DF89, FL85, FW87, GLSZ87, Gro84a, Ioa81, ITH88, JT85, PA89, PB84, Roo85, WP86, WKGC87, WG81, vLK88, vdC82, Goo88b]. Computational [Ano88m, Ano89k, DW89, Zab88, Ano86j, Ano87n, Ano87o, Ano88l, Ano88n, Ano89o, Ano89p, Fre84, MS83, NA88]. Computations [Luk80, DSH89, NR85]. Compute [Ano84f]. computed [Kam83]. Computer [Abd81, Goo88b, CK87, ELS88, Goo88a, Gre81a, Gre81b, Hof87, Kli88, Niz84]. computers [HH89, Hak88, RLW89, SvdV89]. Computing [Maa85, VS88, Wal88, Cas88, CvD89, CK82, Cos88, FVD87, Gab80, McC81a, Rok86, SV87, VVH87]. concepts [Roc88a]. concerning [Gro84a, LLM82, LM86b]. condensed [DSH89]. Condition [NR85, Yey89]. Conditioning [BW89, Bul84a, Kam83]. Conditions [Heb82, BP80, CK80, EO82, Ioa89, PW81b, TW86]. conduction [MP88, Mur88]. conferences [Ano87a, Ano88a, Ano89a]. configuration [Day84a, Day86]. configurations [DM84b]. confluent [Tem81]. Conformal [TW86, Ber86, Del87, Gut86b, HO86, Hoi86, Hou85, Hüb86b, KT86, May86, PWH83, PW84, PWH86, PKW87, Pap89, Rei86, Weg86a, Weg86b, Weg88, Zem86]. congress [Ano89n]. conical [KG87a]. Conjugate [vdVD88, Gut86a, Kaa88, Lip84, MS88, MH83, Seg84]. conjugate-gradient [MH83]. conjugation [Gut86b, Hen82]. connected [May86, PW84, Weg86b]. connection [AEH81]. conservative [OS87]. constant [Jan87]. constrained [GMSW88, Seg84, Sny88, Ste85]. constraints [GLG85, KG87a, LW88, Mur87]. constructing [HO86]. Construction [CH87b, FS87, Jan87, SS89a]. Constructive [Gau85, ÁKS88]. contender [LD86]. continual [GW81]. Continuation [CPS89, Gar86b, MM89, ACG89, BM89, HM89, SF84, Slo86]. Continued [BS83, Gd80, BL87b, Bre87b, Bul80, Cuy87a, CV88, JW86, Jac87, Jac88, Lem87, Lem89, Lev86, LB88, McC83, Sie80, Sie83, Waa87a, Waa87b, dB83, vdC82]. continuity [CWW84, WW89]. contour [Ioa89]. contours [DSV88, Ioa89]. Contraction [Aga83]. contractivity [De 89, Spi85]. Control [Sha87, Ano88f, Cas88, FB89, KTS88, LY89, Niz84]. controlled [MW89, Özi89]. convection [Bir81, Gol86, MG80, NSOS88, Van80, dGvV89]. convective [IS85, Wil86]. converge [Dah86]. Convergence [BL87a, Bre85, CM86, Dru84, Ran87, RV88, Sab87, SL86, SB88a, Van89c, Ado84b, AR86, Aga88, AL86, BL87b, Eie84, EL88, Fuk84, GM81, GMS88, Hal84, Jac88, JH89, KL87, LQ86, Lav84, LB88, Mag87, Mar88, Rab87b, Sho89, Yey89, dB83, WO80b]. Convergent [Ado84a, Paa89, TFI88]. Convex [Met83, Met84, Fuk84, Tun87]. convolution [BD87, Soh89, Tem85]. coordinate [Özi89]. corner [PWH86]. corners [BP86, Ruo87]. Cornu [MW89]. Correction [Ano86e, Hem82, MH83, Ten87]. corrections [Sol89]. correctness [Van89a]. corrector [SF86, SF87, Sha81]. correctors [Tho82a]. correlations [DI89, EL89b]. correspondence [Jac87]. corresponding [McC83]. costs [Ano89f]. Coulomb [Bar82]. counterexample [Ado89]. countries [Ano88f]. coupled [JH89, VDD80]. course [Ano88l]. CRAY [HH89]. criss [DM84a]. criss-cross [DM84a]. Criteria [Szu89]. critical [Nel88]. cross [Cam80, DM84a, Tem88]. cross-rule [Tem88]. cross-section [Cam80]. CTR [KTS88]. Cubature [Kea87, CH87a, CH87b, Sol87]. Cubic [KR86, PS87, ALE84, BP80, CS88, Met83, Met84, PW81a, PW81b, Pap82]. curiosity [Gre86b]. curvature [MW89, Ron86]. curve [Kli87b]. curves [BM87, Cad87, DRV89, LY89, MW89, RR81, Ruo87, Weg88]. cusp [GR89b]. cuts [BGN81]. CWI [Goo88a, Van89a]. CYBER [NNOP86]. cylindrical [Dyk86].

D [Bul88b, SvdV89]. DAE [Mie89]. Dahlquist [Jel86]. damped [BM89]. data [DD87, Die86, Far86, FK88, Kle87, LR89, Mur87, Van89a]. Dawson [Col87a]. DE-formula [Sug87c]. DE-rule [Mor85]. decade [Bre85]. decentralization [AZ88]. Decision [AZ88]. decomposition [Ado84b, AR86, AR88, Ado89, AP88, Nel88, WW89]. decompositions [BGG88]. decrease [EL89a]. defect [Hem82, Ano86e]. defective [SB88a]. deficient [Hem82]. defined [Sid87]. defining [Hop82]. definite [BW89, GHLN89, PR87, vdV87]. definition [Jac88]. deflection [AA82]. degree [CH87b, Cos88, Lip84, ST87, Soh89]. Dekker [Ano86f, Ano88f]. delay [AM84, Bel84, Bru89, FB89, Tor89, vdHS84]. dependence [KTS88]. dependent [De 85, HM89, KG87b, NSOS88, ZH80]. depending [Cri86]. derivation [Yam85]. derivative [DC81, Kar88a, Kar88b, PW81b, TW86]. Derivatives [Waa87a, ITH88, Oli80c, PA89, Rok86]. derived [HC85]. deriving [KG87a, Sid81]. Design [CC86, Saa88b]. desired [KPC80]. Detection [Han87]. determinants [WP86].

Determination [Ioa86a, GL87a, PWH83, PM83, Rei84]. determine [Hoi86]. deterministic [Gre87, GLSZ87]. development [Roc88a]. developments [GMSW88]. device [CPS89, WW89]. Dfl [Bul88b, Goo88b, Goo88a]. diagonal [CK82, Lip81, dB80]. diagonally [Paa89, Som87]. difference [AE89, BZ89, Bic82, BC82, CK80, CK82, CS86, Cha87, Dic85, Dic89, FVD87, Jel85, LLM82, LM86b, Mas87, McC81b, Nii86, Pan89, VS85, Vul89, ZH80, vdHS86]. differences [CV88, Ypm87]. differentiable [Gwi81, Kar88a, Kar88b]. Differential [Ano86f, AM84, BL87a, BTB88, Bel84, BZ89, Bic82, Bru82, Bru86, Bru89, Car89, CN86, Col88, Cri86, Da 87, Day86, Dic82b, Dix86, Dix87, FG86, FB89, GR81, Heb82, Her83, Jan87, JV89, JH89, KLS87, Kri81, Kri82, NF84, OP85, PR88, Roc88b, RM89, SB84, Sha88a, SGC85, SG88, Tt89, TBB88, Tor89, VS85, VDD80, Vol88, Ypm86, ZH80]. differential-difference [Bic82]. differentiation [DN87, Fis89, Oli80a]. difficult [Zem86]. diffusion [Bir81, Dix87, EM86, IS85, KP89, MMSY84, Man84b, MG80, MS89, NSOS88, Özi89, SV87, dGvV89]. diffusion-controlled [Özi89]. digital [CC86]. Digitale [Bul88a]. dimension [Wan85]. dimensional [BH83, ES85, GM80a, GG88, How84, KR87, Smi87a, Son88b, Sug87b]. dimensions [AE89, Bog86, De 84, Gro88, Lip83a]. dipole [MSW89]. Direct [Bul84b, Lip83a, Cal86, DL82, LM86a, Lin89, SS89b]. direction [Lee86, dBH83]. directly [SG89]. discharge [AHCW86]. discontinuities [WO80a, WO80b]. discontinuity [Ioa89]. discrepancy [Lam85, Léc89]. Discrete [Weg88, Abd81, Bla88, DL82, Dic82a, Dun80, Dun82, Kle82, MS89, Mur88, Mur87]. discrete-time [Kle82, Mur87]. discretisation [MC85]. discretization [Had85, MG89, Nev85, vdHS86]. discretizations [GR89b]. discriminant [DN87]. Discussion [vdC82, dLM89]. disks [Fab87]. dispersion [VS85, vdHS86]. dispersive [Ado88]. dissipative [Ado88]. distance [VCL87]. distinct [Paa89]. distribution [DC81, EHL89, KKN^{+}88, Van85]. divergence [Wal87]. divergent [Sid87]. diverging [How84]. Division [Bul80]. DM [Bul88a, Ano86e]. Docev [Ano84e, Sal83]. doctoral [Ano87j, Ano87k, Ano88k, Ano89l]. does [JW89]. Domain [WW89, AP88, BC82, Bon87, JT85, Pri80, Wad80]. domains [BP86, Dyk86, PW84]. dominant [Paa89]. dominated [dGvV89]. Dordrecht [Ano86d]. double [AE88, MC85, Sie80]. doubling [EKL89]. doubly [PW84, Weg86b]. doubly-connected [PW84]. Dr [Cuy87b]. drawing [MW89]. drift [GGD81]. Dual [Bul84c, EM87]. ducts [How84]. Dynamic [Mie89, AZ88, NA88]. dynamical [KLT89]. dynamics [ELS88, Zab88].

E.A.D.I. [AEH81]. earthquake [DN85]. East [Ano89f]. ecology [Goo88b]. economic [AZ88]. economical [VDD80]. economics [FH88, Puu88b]. ed [Ano88f, Goo88b]. Edge [RC86]. edited [Ano86d, Ano86e]. Editorial [Ano83b, Ano86h, Ano87f, Ano88i, Ano88j, Ano89g, GGW83, Ano84c, Ano84d, Ano86g, Ano86i, Ano87g, Ano87h, Ano87i, Ano88h, Ano89h, Ano89i, Ano89j]. Eds [Ano89f]. Effect [NNOP86, KB88, SB88b]. effectiveness [JY88]. effects [DN85]. Efficiency [BR89, MS83]. Efficient [Nat86, RWP85, SW89, Sha81, Ypm87, SF87, VVH87]. eigenelements [DR87a]. Eigenfunction [RD87]. Eigenfunctions [CGB84]. eigensolution [Wad80]. eigenvalue [BS89, CW85, DSH89, NNOP86, NSIO82, ST86]. eigenvalues [CK82, FVD87, Gd80, KP82, Paa89]. eigenvectors [FVD87]. eight [CR87]. elasticity [KG87b, Mil89]. electrical [AHCW86]. electrocardiography [BdTL85]. electrochemical [FR80]. electrons [Gre86a, Gre86b]. Electrostatic [Fab87]. element [Ami87, BE89, BB88, DM84a, De 85, DMP87, HSM83, OS87, Ong84b, PMD87, RD87, Ruo87, Son88b, TS86, ZN84, dG81]. Elementary [GL83, Gre86a]. elements [IS85, KN87, PD88, PV85, Sab85]. Elimination [GL87b, BR89, DM84b]. Elliptic [Dic87, AE89, AN81, BB88, BC82, Bla88, BH83, Bud89, Dyk86, EO82, Had85, Lav84, Lip83a, Nak87, Ong84b, Smi87a]. Elsevier [Bul88b]. embedded [DP80, DP86, PD81]. Emerging [ACCK88]. End [BP80, PW81b, Sol89]. energy [AHCW86, Bog86]. Engineering [Cas88, Usm80]. entire [FL85]. entrant [PD88]. entropy [Eri88]. environment [HM85]. epsilon [Dra87]. equation [BTB88, Ber86, Bir81, Cam80, Car86, CH88, Dix87, FH88, Gd80, Hüb86a, Ioa88, Kle82, Léc89, Mu89, PF84, RV88, RM89, RS89, SS89b, Tem85, TW86, VCL87, Wad80, Wil86]. equations [Ado84a, AHCW86, AM84, AE89, BD87, Bar83, Bel84, BZ89, Ber88, Ber89b, Bic82, BC82, BZ84, BH83, Bru82, Bru86, Bru89, Bud89, Cah81, Cal85, CM84, Cap88, Car89, CK87, CN86, Col88, Da 87, Day86, De 84, DMP87, Dic82a, Dic82b, Dic85, Dic87, Dic89, Dix86, ED89, FG86, FB89, FW87, GLG85, Ger86, GR81, GR89b, HLS88, Hal82, HC85, Heb82, Her83, Hou85, Iga85, Ioa82, Ioa84, Ioa86b, ITH88, Jab80, Jan87, JY88, Joe85, Kli87a, KG87b, KLS87, Kri81, Kri82, Lav84, LLM82, LM86b, Lip83b, Lip84, Luk80, Luk82, Mae88, MD86, Mar86, MO87, NF84, NSOS88, NR85, OP85, Pan89, PR88, PJ80, PE87, Rei86, Roc88b, SF87, Say88, SB84, Sha88a, Sha88b, Sid88, Son88b, SGC85, SG88, Tt89, Ten87, TBB88]. equations [Tor89, VS85, VDD80, VFD89, Vol88, Wam80, WW89, Xin88, ZH80, vdHdV83, vdHS84, Ano86f, Lev88]. Equidistributing [HSM83, TS86]. equilibrium [NA88, dLM89]. equivalent [GW81, KPC80]. ergodic [Lam85]. Errata [Ano84f, Ano84e]. Erratum [Ano82, Ano89k]. Error [BF89, CK87, SB84, Ber89a, DM82, Dob88, EL88, Hen82, ITH88, JH89, LS88, Lau85, Let81, LD87, MG89, Mil89, Nii86, Pri87, Sha81, Van89a, Yam85]. Error-free [CK87]. error-state [Van89a]. errors [Wil86]. estimate [Hof89, Sha81]. Estimates [Köh89, Che84, Dob88, Hen82, ITH88, Let81, Mil89, Pri87, RV88]. Estimating [Van87b, Gre86b, NIS82]. Estimation [MG89, Ber89a, Day84b, Lau85, Mur82, NSO80, Son89, Ypm87]. estimator [DM82]. estimators [EHL89, SB84]. Euclidean [Akl81, Gro88, Kle87, Ren88]. Euler [Cuy87a, Dic85, Dic87, Dic89, GLG85]. Europe [Zla85]. Evaluation [AR88, PMD87, Ala83, Bar82, DN85, DL82, GG83, GK86, GG88, Gut86a, Kip89, Luk82, Nat86, Oli80b, Tez89]. evaluations [DM83]. Even [JW86, Son88a, Ten87]. exact [Gri82, Sug87a]. exactly [VS88]. example [Gro85]. examples [Nel88]. Existence [Aga87, CS87, Seg84, Tho82a]. Existence-uniqueness [Aga87]. expansion [GG85, HT87, JW89, MS83, RD87, Ste87, Tem81]. expansions [CV88, GL87a, GL84, JNT83, KK89, LD87, WG81]. expectation [BLS84c]. expected [Akl81]. expensive [DR87a]. explicit [CS84a, CS84b, Cha86, CR86, CR87, Day86, Kle82, VSV89]. exploitation [Zla87]. Exponential [SB88b, Bel86, Col80, Con87, EHL89, Nii86, SH87, Wal88]. Exponentially [Car86]. exponentials [DD87, MR87]. expression [Kle82]. Extended [Mar88, Cha86, Van89b]. extension [Dic85, Sug87b, Yey88]. exterior [Ami87]. extraction [Bre89]. extrapolated [Sar88]. Extrapolation [Sid87, Sid88, Bre83, Had87, MR87, Sha87, SB88a, Sug87c].

F [Bul88a]. factor [LQ86]. Factoring [RLW89]. factorization [Axe85]. factorizations [BW89, Han87, PP89]. failure [DN88, Nel88]. Families [CN86, NF84]. family [DP80, JNT83, PR88, Sab83, Twi86]. Fast [HLS88, Ong84a, Roh87, Cam80, Hüb86b, Lin89, Rei86]. FE [NS82]. FE-method [NS82]. Feasibility [Spi85, Tun87]. Fejér [Rab89]. few [Dun88]. FFT [BD87]. field [Xin88]. Filon [Bel86]. filter [Bul88a]. filtering [Bul84a]. filters [CC86]. finding [Iga85, MW87]. fines [Spo89]. Finite [CK80, Gol86, Ong84b, Vul89, ZN84, Ala83, BE89, BB88, BC82, CK82, CS86, Cha87, DM84a, DMP87, FVD87, HSM83, LLM82, LM86b, Lon84, Mas87, OS87, PV85, Sab85, Son88b, TS86, ZH80, dG81]. Finite-difference [Vul89]. finite-element [BE89, BB88]. first [Cap88, Dic82b, ED89, Gre86a, Gre86b, HC85, Hou85, ISK87, IS88, Rei86, VDD80]. first-order [VDD80]. fit [DD87]. fitted [Car86]. Fitting [Kle87, Die86]. five [CK82]. five-diagonal [CK82]. fixed [Kle82, Sab87]. fixed-lag [Kle82]. Flett [Dör85]. flow [eMH89, AMHAEM89, AN81, BeMM86, Dic87, ES85, ET86, NS82, NR85, Niz84, Wat81]. flows [Oue82]. fluid [AN81, ES85, NS82]. fluids [RS86]. flux [Dic85, Dic89]. flux-difference [Dic85, Dic89]. focal [AU86]. fold [DRV89]. Forecasting [Col80]. form [CvD89, Dic82a, FG86, Gil88, Kip89, KPC80, Sha88b]. formal [PR87]. formation [How84]. forming [Sha84]. forms [DSH89, Rok87]. formula [Ehr88, FG86, Gro88, IMT87, PO87, Son88a, Zem86, Sug87c]. formula-pairs [FG86]. Formulae [Gro84a, Ano89k, CH87a, CH87b, DP80, DP86, EM87, FB89, GN89, PD81, SS89a]. formulas [BF89, GM84, Gri82, Kea87, Köh89, Mor85, Sha81, Sol89]. formulation [NA88, ZN84]. Fortran [Dun80]. Foundations [Goo88a]. four [Wam80, dB83]. four-term [dB83]. Fourier [Ala83, Cam80, Che84, Lon85, Neu81, Soh88]. fourth [AC84, Ano84f, BB88, CK82, CN86, Cha86, CRN86, Cha87, CSS88, CS88, DM82, Day86, KJG83, PW81a, RM89, Tho82b, vLK88, vdHdV83]. fourth-order [BB88, CK82, CRN86, Cha87, CSS88, CS88, Day86, PW81a, RM89]. fraction [CV88, Gd80, JW89, McC83, SK89a, SK89b]. fractions [BS83, BL87b, Bre87b, Bul80, Cuy87a, JW86, Jac87, Jac88, Lem87, Lem89, Lev86, LB88, MS83, McC81a, Ran87, Sie80, Sie83, Waa87a, Waa87b, dB83, vdC82]. frame [VCL87]. framework [Goo88a]. Fredholm [Ber86]. free [CK87, DR87c, ET86, Slo86, Sny88]. free-streamline [ET86]. frequency [Ehr88]. Freud [KL87]. Froude [eMH89]. fully [CH87b]. function [BGV81, Cad87, Cub84, Dör85, DC81, FG89, Gab80, Gut86a, Gut86b, IS88, Kar88a, Kar88b, LW88, McC83, Oli80a, PWH83, PB84, Rok86, Ron86]. function-values [FG89]. functional [GLSZ87, Lev88]. functionals [Köh89, PR87]. functions [Ano89k, BS83, Cle81, EL89a, Eie84, EM86, FL85, GN89, Gil88, GL83, GM81, Gri82, Gro86, HT87, HS82, ISK87, Ioa86a, Ioa89, JNT83, JT85, Kli87b, Laf86, LS88, Let81, Lev88, MR87, Nat86, Pre87, Pri80, Roh87, Sid81, Sie83, Soh89, Sug87a, Sug87c, Tem81, Waa87a, Wal87, Wal88, dBH83, vLK88, Ano86d]. future [Kli88, MP88]. FX [Duf89]. FX/8 [Duf89].

G [Ano86f]. Galerkin [KG87b, MG80, Nak87, Vol88, dGvV89]. gamma [JT85, LS88, McC83]. gate [AMHAEM89]. Gauss [Ano89k, GN89, Hof89, JV89, Let81, PE87].

Gaussian

[BR89, BF89, DN87, GG88, Gro88]. gearlike [BP86]. Gegenbauer [Gro86, Rab87b]. General [Jac87, Cha85, DG82, Gro88, Kar88a, Kar88b, Lin89, NA88, Oue82, RR81, Sho89]. generalised [GM80b, GMS88]. generalization [Cla81, Gab80]. Generalized [Dix86, HKD86, Ano84e, CvD89, Che84, DR87c, Dyk86, EL89b, GR89a, Lev86, Sal83, vdC82]. generated [AG88, dB83]. generating [Gro86, Pre83]. generation [CSR87, Gro85, Joe89, ZH89]. generators [AG88]. Generic [LST89]. geometrical [SGC85, SG88]. geometry [Ano88l]. Germany [Bul88a]. Geronimus [Pre88]. Geronimus-type [Pre88]. given [Ast82, ES85, Mei81]. gives [Sug87a]. Global [Mu89, KN87, MG89, PMD87]. globally [TFI88]. GNLMR [HKD86]. Goel [Ano84f]. Golumbic [Rij80]. good [Sug87b]. Gordon [CH88, Xin88]. gradient [ÁKS88, KN87, Lip84, MS88, MH83, Nak87, vdVD88]. gradients [Kaa88]. grammar [Goo88a]. graph [DC81, Maa84, Rij80]. graphical [Rou87]. graphs [vS86, Rij80]. gravity [AMHAEM89]. Green [Cub84]. Greenspan [Bul88b]. grid [CSR87, HM89, ZH89]. grids [Mas87, MPN82]. Grosswald [Ano84e, Sal83]. guaranteed [Hal84].

H

[Ano84e, Ano86d, Ano86e, Bul88c, Cuy87b]. H/b [Bul88c]. half [ES85]. half-plane [ES85]. Hamburger [Hen87]. Hamming [Bul88a]. Hankel [Bul84b, Bul84c]. harmonic [Son88a, Ste87]. Hartman [Bon87]. heart [MSW89]. heat [MP88, Mur88, PF84]. heated [SR88]. HERMITE [Met83, Met84, AL87, Bak84, BL87a, Pas87, Rab89]. HERMITE-spline [Met83, Met84]. Hermitian [Yey88]. Hessenberg [Gra86]. heteroclinic [DF89]. hierarchical [MS88]. High [AT80, Bar82, Bru86, FR80, JKG83, Lin86, PD81, VS85, AHCW86, Bog86, SS89a]. High-order [Bru86, JKG83, Lin86, SS89a]. High-precision [Bar82]. Hilbert [Hen87, Weg88]. Hilger [Lev88]. Hill [WP86]. holes [Dyk86]. Holland [Ano86f, Bul88b, Goo88b, Van89a]. homogeneous [Jan87, PC80, VCL87, Wad80]. homotopy [Wat81]. Hopf [DRV89, DW89, Roo85, SCJ89]. hopscotch [BH83, Ten87]. Huberman [Goo88b]. hundred [DC81]. Hungary [Bul88b]. hybrid [Dic82b]. Hydrogen [Gre81a]. hyperbolas [Yam88]. hyperbolic [Jel85, Sny87, VS85, vdHS86]. hypercube [BR89, EKL89, GHLN89]. hypergeometric [Sid81, Tem81, Waa87a].

IBM [HH89]. ideal [NS82]. identification [SOA82]. II [AP88, Bul84c, CSR87, CR86, Gre81b, KK89, Lin86, Man84b, SG88, WO80b]. ill [Bul84a]. illustrated [Jac88]. illustrative [Gro85]. imbedding [SOA82]. Implementation [Roc88b, BE89]. implications [PP89]. Implicit [Mas87, Lip84, Mit87, PR88, Sha81, Som87, Spi85]. improved [PW81b]. improvement [AL86]. improvements [PM83, PS86]. impulse [KB88, SB88b]. inclusion [Pet82]. Incomplete [Axe85, JT85]. incompressible [KG87b, Ten87]. increase [Mam89]. Indefinite [HT87, Saa88a]. Index [Ano89d, Ano83a, Ano84a, Ano84b, Ano84g, Ano85a, Ano86a, Ano86b, Ano86c, Ano86k, Ano87e, Ano87b, Ano87c, Ano87d, Ano87p, Ano88d, Ano88e, Ano88b, Ano88c, Ano89b, Ano89c, Ano89e]. Inequalities [Laf86, HM89, LS88, MM89, dLM89]. inequality [Dun84a, Dun84b]. infinite [CS86, GG83, KR86, Lyn85, PD88, SW88, Sid87, Sug87a, TFI88, Tez89, Wad80]. infinitely [Kar88a, Kar88b]. infinity [Gro84b]. Influence [RR81]. information [BdTL85, Kli88, Ros88, Saa88b, vdV87]. Informationstechnologie [Bul88a]. initial [Ana86, BZ89, BOP84, CR84, Cha85, CR86, Chu83, GSB87, JKG83, Jai88, JG88, Kri81, Kri82, Mu89, SB88b, Sny87, Twi86, WO80b]. initial-boundary [Mu89]. initial-value [BZ89, CR84, CR86, Jai88, JG88, SB88b]. Inner [Gri82]. input [CC86]. instability [CH88, GR83]. Institutional [Ano87j, Ano87k, Ano88k, Ano89l]. instrument [Puu88b]. integer [GW81]. Integral [Cub84, Ber86, Ber89b, Bru82, Cah81, Cal85, CM84, Cam87, Cap88, Col87a, Dix86, ED89, Ger86, Gro84a, GG85, HLS88, HC85, Hou85, Hüb86a, Ioa82, Ioa84, Ioa86b, Joe85, KG87a, Köl86b, Köl86a, Köl87, MC85, MO87, NR85, PE87, Rei86, RS89, Sug87a, Tem85, dD84]. integrals [AR88, AE88, Ala83, BLS84a, BLS84b, Ehr88, Gab80, GL87a, GK86, Gla84, GLSZ87, GG88, Gro88, Ioa81, Ioa86b, Ioa89, KG87a, Let87, Ors89, PB84, Rae82, Sid87, WKGC87]. integrands [ALE84]. Integrating [Lyn85]. integration [ALE84, Ana86, BGV81, Ber89b, Bic82, CM84, Car89, CR84, Cha85, CR86, FW87, GLG85, GM80a, HT87, JKG83, JV89, Joe89, KR87, KL87, Lau85, PR88, Rab87b, Rab87a, Rab89, RD81, RWP85, Slo85, SO87, Sug87c, SS89a]. integration/interpolation [Bic82]. integrations [Sug87b]. integro [Bru82, Bru86, Bru89, Dix86, Dix87, Vol88]. integro-differential [Bru82, Bru86, Bru89, Dix86, Dix87, Vol88]. integrodifferential [PJ80]. Intelligence [Goo88b]. Interaction [DRV89]. interactions [LLM82, LM86b]. Interactive [ZH89]. interface [RS86]. interior [Rab87a]. International [Ano88l, Ano89f]. interpolants [DR87b, GMM89, GM84]. interpolating [Cos88, LY89]. interpolation [AL87, Bic82, Bre83, Cad87, CDR89, Dun84a, Far86, FK88, Fre84, GL87b, Hel82, Hen82, KL87, Mac82, McN86a, McN86b, MPN82, Met83, Met84, NSIO82, Oli80c, PS87, Pri87, Rab89, ST87, SH87, Weg88]. interpolation-basic [Fre84]. interpolatory [BP80, BP88, Ioa84, Rab87b]. interpretation [Ioa86b, PR87, Rou87]. interpreting [Maa85]. interval [GS80, Gar86b, Gar86a, Maa84, Sug87a]. intervals [Cha86, KR86]. intrinsic [Kam83]. introduction [ACCK88]. Invariance [Bak84]. Invariant [SOA82]. Inverse [KK89, Kri81, DL82, GMS88, Hem82, Hof89, MP88, Mur88, PP89]. Inverses [Pap82]. inversion [HH84]. invited [Ano85b, Ano87l, Ano89m]. involving [Sha88a]. ions [Gre81a, Gre87]. iPSC [BR89]. iPSC/2 [BR89]. irregular [Bar82]. ISBN [Ano88f, Ano89f, Bul88a, Bul88c, Bul88b, Goo88b, Goo88a, Lev88, Van89a]. isoclinal [Men87]. isoparametric [BM87, DMP87]. isothermal [KP89]. iterated [Vol88]. iteration [AEH81, BPW89, Mie89, PS86, SB88a]. iterations [Bre89, Hal84, McN88]. Iterative [AC84, AU86, Ber88, DMP87, FH88, Miy89, MO87, Sha88a, Aga87, Aga88, Ami87, AL86, CG89, DR87a, Had85, Had87, HKD86, MS86, NSO80, NIS82, Not88, Pet82, RV88, SL86, Sar88, Say88, SW88, Weg86a, Weg86b, vdV87]. ITPACK [KY88].

J [Ano82, Ano86d, Ano86e, Ano86f, Ano88f, Lev88, Van89a]. J. [Ano84f, Ano84e]. Jacobi [CvD89, Deh81, Gro86, Hou85, KP82, Paa89]. Jacobi-like [CvD89, KP82]. Jacobian [DM83, Ypm87]. Jacobians [Day84b]. Jain [Ano84f]. Janos [Bul88b]. Janssen [Goo88a]. Jenkins [Col80]. Johnstone [Bul88c]. Jordan [Hof89]. Journal [Ano83c, Ano83d, Ano89k].

Kaczmarz [MD86]. Kambo [Ano84f]. Karasö [Ano89f]. kernel [Cub84, Ger86, KT86]. kernels [PJ80, Rae82]. Khovanskii [Lem89]. kind [Ber86, CM84, Cap88, ED89, Gro86, HC85, Hou85, Joe85, Rei86, vLK88]. Klein [Xin88, CH88]. knapsack [DKW84]. knots [BP80, CH87a]. known [Ioa89]. Kronrod [Ano89k, GN89]. Krylov [vdV87]. Kutta [Ano86f, CS84a, CS84b, FG86, Som87, CQ82, DP80, DP86, Heb82, LDS87, PD81, Sot84, Spi85, VSV89, VDD80, Wam80].

L [DM82]. lag [CR84, CR86, CRN86, CR87, Kle82, Twi86]. Lagrange [GLG85, KL87]. Lagrangian [McN86a, McN86b]. Lanczos [Col87a, CW85, Deh81]. Laplace [Cam80, HH84, Rei84, Tem85, TW86, ZH80]. Laplacian [Pap89, Rae82]. Large [ACG89, eMH89, CW85, ITH88, Lip84, aON88, Ong84a, RLW89, VDD80, Zla87, dG81]. large-scale [ITH88]. Lattice [Slo85, AG88, Joe89, Sug87b]. Laurent [KK89]. Lauricella [vLK88]. layer [MC85]. layered [LR89]. layers [ZN84]. leads [BdTL85, DRV89]. Leapfrog [Say88]. Least [PP89, VCL87, Bul84a, Far86, NS82, SF84, Sar88, VV88, Van89b]. least-squares [Sar88]. Lecture [Ano88f]. Lee [Ano88f]. Legendre [JV89, Let81]. lenticular [SGC85, SG88]. level [Hun88]. like [BGN81, CvD89, KP82, Ypm86]. limit [BL87b, Bre87b, JW86, Lem87, Zie84]. limits [JM80]. line [Bla87]. Linear [HH89, MR87, Waa87b, Abd81, AL86, BL87b, BV86, Cah81, CK87, CM86, CC86, CG89, DN87, Dun84a, Dun84b, GGD81, GR81, Had85, Had87, HC85, HM85, Hof87, Hüb86b, Jab80, Jan87, JY88, Kli87a, Köh89, KPC80, Kos88, Kri81, KN87, LM82, Lip81, Lip83b, Lip84, LW88, Luk80, Luk82, Mae88, MS86, MG89, Mie89, Mil89, aON88, Ong84a, PW81a, Saa88a, Say88, SvdV89, SW88, Sho89, Sid88, Sny87, SV87, TT87, Van80, Vol88, Yey88, ZH80, vdHS84]. linearization [Bon87]. Linearized [BeMM86, KKK89]. lines [GR81, Sol87]. Liouville [Bar83]. List [Ano83c, Ano83d, Ano85b, Ano85c, Ano87l, Ano87m, Ano89m, Ano89n]. literature [FH88]. Local [Dic82a, Lav84, DR87b, Tt89, Wil86]. locating [Bul84a]. Location [Ioa89, EHL89]. logarithmic [ALE84, MR87, Sab87]. logic [Bul88c]. Lommel [Gro84b]. Low [Léc89, Smi87a, DMP87, Lam85]. Low-dimensional [Smi87a]. low-order [DMP87]. lower [Maa84]. LQP [Mur87]. Lyapunov [KB88].

M [Goo88a]. Macleod [McN86b]. mailing [Ano89f]. maintaining [BdTL85]. manipulation [For86a, For86b]. many [Bar83]. map [Zem86]. mapping [Ber86, Del87, Gut86b, Hou85, Hüb86b, KT86, May86, PWH83, PW84, PWH86, PKW87, Pap89, Rei86, TW86, Weg86a, Weg86b, Weg88]. maps [HO86, Hoi86]. Marcel [Ano88f]. Martin [Rij80]. martingales [DG82]. Master [Ano84g]. matched [Chi81]. Math [Ano82, Ano84f, Ano84e]. Mathematica [Bul88b]. Mathematical [CSR87, MSW89, Zla85, ACCK88, Gla84, Gre86b, Roc88a]. Mathematics [Ano88m, Ano89k, Goo88a, Ano86j, Ano87n, Ano87o, Ano88f, Ano88n, Ano89o, Ano89p, GS80, Gar86b]. matrices [BW89, BGG88, Deh81, DSH89, Gar86a, Gd80, Gra86, Han88, KP82, Men87, Paa89, Pap82, SB88a, Sug87b, Szu89, Ypm87, Zie84]. Matrix [Nev85, PR87, AG88, Ast82, Axe85, Duf89, Hof89, JH89, Mag87, Mei81, Smy85, VVH87, VS88, Wal88, Yey88, Zie84, vdV87]. matroids [Ano82, Ham82]. max [Niz84]. max-min [Niz84]. maximal [Gri82]. Mean [KL87]. means [Jac88, Mar86]. measurement [Oue82]. measurements [Mur82]. measures [KG87a]. mechanical [GLSZ87]. mechanism [Saa88b]. media [Ado88, PF84]. medium [LR89, RS86, SR88, Van80, Whe85]. melting [ÁKS88]. memory [MS88]. meromorphic [Ioa86a, Pre87, Wal87]. mesh [Cha87, Gut86a, Lee86, Ruo87, dBH83, dG81]. mesh-width [dG81]. message [BE89]. Method [Sug87b, AR86, Ami87, Ano84f, AL86, BE89, BB83, BB88, BM89, BPW89, Ber89b, BC82, Bla88, Bla87, Cal86, Cap88, CK82, CR84, CS86, CR86, CR87, Cha87, CSS88, CS88, Chu83, Cle81, Col87a, Cri86, Da 87, DM82, Day86, De 85, DG82, Deh81, Dic85, Dic89, Dun84a, Dun84b, Dyk86, EM86, EO82, EL88, Gab80, GK89, GR81, Gre86a, Hoi86, HH84, HKD86, Hüb86a, Hun87, Ioa88, KG87a, KJG83, Kiw86, KB88, Lav84, LW88, MS86, Mam89, Mar86, Mar88, Mit87, MO87, MH83, NNOP86, Nel88, NSO80, NSIO82, NIS82, OS87, aON88, OP85, Özi89, PW81a, PC80, Pet82, PE87, Rei86, Rok86, RD87, Ruo87, RS89, SF84, SF86, SF87, SW88, Sid81, ST86, Sny87, Sny88, Som87, Son88b, SV87, SS89a, TFI88, Tun87, Vol88]. method [Wat81, Weg86a, Weg86b, WO80b, Xin88, Yam88, Yey89, ZH89, dG81, dGvV89, dB80, vdHdV83, vdV87, NS82]. methodology [MS83]. Methods [Ano86e, Ano86f, Sug87c, AC84, AU86, Aga87, Aga88, AM84, Ana86, Ano88f, Axe85, AP88, Bic82, Bre85, BH83, Bru86, BOP84, CQ82, CM84, Car86, CK80, CS84a, CS84b, Cha85, CM86, CN86, Cha86, CRN86, CG89, Col88, CPS89, Day84a, De 84, Del87, DW89, Dic82b, Dix86, FVD87, Far86, FW87, FS87, Gut86b, Had85, Had87, Heb82, HSM83, Ioa82, JKG83, Jai88, JG88, Joe85, Kho86, KG87b, Kri81, Kri82, KJ88, LB88, Lip83a, Lip84, LDS87, MW87, Mae88, MD86, MG89, Mas87, May86, McN88, Mie89, MG80, Miy89, NF84, NSOS88, NT86, PR88, PJ80, PO87, PM83, Phi87, Roc88b, Roo85, SL86, Say88, Sho89, Sid87, SB88a, Sid88, Slo85, Sot84, Spi85, SS89b, Tho82b, TS86, Tor89, Twi86, TT87, VSV89]. methods [VDD80, VFD89, VCL87, Wam80, Ypm86, vdHS84, vdVD88, Bul88b]. metric [dB80]. microelectronics [RD87]. micropolar [ES85]. Middle [Ano89f]. migration [Wei88]. migratory [Wei88]. Min [Niz84]. Min-max [Niz84]. Minding [Cuy87a]. minimal [CR84, CR86, HKD86, Sol87]. Minimax [EHL89]. minimisation [FS87]. minimization [Kiw86, SF86, Seg84, Sny88, SGC85, SG88]. minimizing [Ste85]. minimum [Day84a, Day86]. miscellaneous [Gro86]. mixed [BB88, BH83, CK80, IS88, KKK89]. MM [KTS88]. MM-algorithms [KTS88]. modal [Dic82a]. model [BV86, DN88, DG82, EM86, Eri88, MS89, MSW89, NA88, Oue82, SV87, Wil86, Zla85]. modeling [Gre81a, Gre87, Puu88b]. modelling [ACCK88, Gre81b, Spo89]. models [DM84b, KKN^{+}88, WW89]. modifications [Bre87b, PP89]. modified [DC81, Mit87, NS82, VS88, Wil86]. modulus [VS88]. molecule [Gre81b]. molecules [Gre81a, Gre81b]. mollification [MP88, Mur88]. mollification-future [MP88]. moment [DR87c, Hen87, Lon84, WKGC87]. Moments [Neu81, Kar88b]. Monographs [Van89a]. Monotone [Aga88, LLM82, LM86b]. Monotonicity [LS88, Hun87]. Montague [Goo88a]. Monte [BC82, Jab80, Lam85]. Montessus [GS86]. Montessus-type [GS86]. most [VDD80]. motion [LST89]. mound [SR88]. moving [Far86, HSM83, Özi89, TS86]. MR [KTS88]. MR- [KTS88]. multi [Aga83, HM89, Sug87b]. multi-dimensional [Sug87b]. multi-grid [HM89]. multi-point [Aga83]. Multiderivative [PO87, TT87]. multidimensional [Ber89a, Cam87, DL82]. Multigrid [Dic89, Dic85, Hun88, Phi87]. multilevel [Bla88]. Multiple [SO87, CM86, DKW84, Joe89, Miy89, Slo85, VS88]. multiple-choice [DKW84]. multiplier [Sny88]. multiply [May86]. Multipoint [WO80a, WO80b, Wal87]. Multiprocessing [Duf89]. multiprocessors [EKL89]. multiregion [BC82]. multistep [CM86, Chu83, Cro87, Kri81, Roc88b, VSV89, vdHS84]. Multivariate [CV88, Far86, CWW84, Cuy85, Cuy86, For86a, For86b, GK86, Wan85]. multivector [MS88].

N [Ano84f]. Navier [Dic87, Ten87]. near [Akl81, BF89, CC86, Han87]. near-optimal [Akl81, CC86]. near-singularity [Han87]. necessary [Yey89]. Nested [Ors89, Kip89]. Netherlands [Goo88a]. network [Niz84]. networks [Cad87, HS82, Maa85]. Neumann [Nev85]. Newbery [Oli80b]. Newsletter [Ano86j, Ano87n, Ano87o, Ano88m, Ano88n, Ano89o, Ano89p]. Newton [BM89, Cle81, FS87, Hüb86a, MO87, Yam85, Ypm86, dB80]. Newton-like [Ypm86]. nodal [Dic82a]. noisy [AR86, DD87, Kle87]. Non [YaOIN85, Cah81, CMM86, Cuy86, Dra87, Dra88, GR81, Gwi81, Jab80, KP82, Kle87, Kli87a, LM82, NSO80, PR87, Sha86, SV87, Tt89, Tho82a, TT87, Van80]. Non-adaptive [YaOIN85]. non-commutative [Dra87, Dra88]. non-definite [PR87]. non-differentiable [Gwi81]. non-existence [Tho82a]. non-linear [Cah81, GR81, Jab80, Kli87a, LM82, SV87, TT87, Van80]. non-local [Tt89]. non-normal [CMM86, Cuy86, Dra88, KP82]. non-smooth [Gwi81]. non-stationary [NSO80]. non-stiff [Sha86]. non-tensor-product [Kle87]. nonautonomous [Day86]. noncentral [EHL89]. nonconvex [Kiw86, Von88]. Nonlinear [Ano86f, MS89, Ado84a, AHCW86, ÁKS88, BTB88, BZ84, CS87, CS88, Dix87, Dun80, Dun82, Dun88, Hal82, HKD86, Iga85, ITH88, KLT89, Lin86, Man84a, MD86, Mar86, MM89, Nii86, OP85, Pan89, RS89, SF84, SF87, Son88b, SGC85, SG88, TFI88, TBB88]. nonsingularity [Szu89]. nonsmooth [Kiw86]. nonspherical [SGC85, SG88]. nonstationary [ES85]. nonstiff [Day86]. nonsymmetric [MS86, Saa88a]. nonuniform [BeMM86]. nonuniformly [FK88]. norm [Bia89, Han88, Hof89, KP82]. norm-reducing [KP82]. normal [Cam87, CMM86, Cuy86, Dra88, GK86, JW89, JV89, KP82]. normalized [Lip81, Lip83b, Lip84]. norms [KKK89]. North [Ano86f, Bul88b, Goo88b, Van89a]. North-Holland [Ano86f, Bul88b, Goo88b, Van89a]. Note [Ano84e, NSO80, Sal83, VSV89, Wad80, Bir81, Cad87, Hem82, Hun87, Kar88a, Kar88b, Lee86, LB88, Luk82, Men87, PM83, Som87, SCJ89]. Notes [Ano88f, Bul88c]. notion [Son89]. Noumerov [CR84, CR86]. Noumerov-type [CR84, CR86]. novel [GL87a, Özi89, WG81]. November [Ano86j, Ano89p, DP88]. nr [Ano86j, Ano89p, DP88]. null [Kam83]. number [eMH89, Akl81, Maa84, Sol87, Van88]. Numerical [Ala83, Cah81, Da 87, DF89, GR83, Gut86b, KT86, LR89, NSOS88, NT86, OP85, PKW87, Pap89, Rab87a, SR88, Sha88b, Slo86, VFD89, Wat81, AM84, Ano89f, AN81, BD87, BGV81, Bir81, Bla87, Bru82, Bru86, Bru89, BOP84, Cal86, Cap88, Car86, Car89, Cha85, Del87, DM84b, Ehr88, ED89, EO82, GM80a, GK86, GR81, Gre87, HLS88, Hak87, HO86, Hel82, Her83, HH84, Hou85, Hüb86b, Ioa82, Ioa84, JKG83, Joe89, Kos88, KLS87, Kri81, Lau85, Mam89, Mur88, Niz80, Oli80a, Oli80c, PR88, PWH83, PW84, PWH86, Pri80, ST86, Smi87a, Smi87b, SGC85, SG88, Sug87b, Sug87c, Tt89, Tez89, Tor89, Van88, Van80, Weg88, ZH80, Bul88b]. Nyström [CS84a, CS84b, FG86, PE87, Som87].

O [Ano89f, Gre81b]. O.D.E. [BOP84]. objective [LW88]. objective-function [LW88]. oblique [TW86]. observations [Let81, WKGC87]. obstacle [ET86]. obtained [Mor85, vdV87]. occurring [AA82]. odd [JW86, Ten87]. odd-even [Ten87]. ODE [DM83, GSB87, KLT89, SG89, Wat83]. ODEs [Dob88, Sha86, Sot84]. oil [Spo89]. omitted [BP86]. One [Bel84, Ioa81, Bla87, Bog86, Car86, Dun88, GG88, How84, JKG83, JV89, LST89, Oli80a, Roh87, Son88b]. one-dimensional [GG88, How84, Son88b]. one-parameter [LST89]. one-sided [Roh87]. One-step [Bel84, Car86, JKG83, JV89]. onto [PKW87, Pap89]. operations [Von88]. operator [Cri86, CWW84, GR89b, MC85, Rei84, Soh89, Ano88f]. operators [Cub84, Hen87]. optics [SGC85, SG88]. Optimal [Ano82, DR87b, Ham82, Rou83, Akl81, Ano88f, BGS88, CC86, CGB84, DM84a, Niz84]. optimisation [FG89]. Optimization [Roc88a, GMSW88, Gwi81, TFI88]. Optimum [Had85, HO88, NSO80, NIS82]. orbital [Van87b]. orbits [CGB84, DF89]. order [AC84, Aga87, AHCW86, Ana86, Ano84f, AEH81, BB88, BGS88, Bre89, Bru86, CK82, CR84, Cha85, CN86, Cha86, CR86, CRN86, CR87, Cha87, CSS88, CS88, Col88, DM82, Day86, DMP87, Dic82b, FG86, FR80, Hal84, Heb82, JKG83, Jai88, JG88, Jel86, KJG83, Kar88a, Kar88b, Kri81, Kri82, Lin86, Man84a, Oli80c, Ong84b, PR88, PW81a, PD81, RM89, Sha87, Son88a, SS89a, SG89, Tho82b, Twi86, TT87, VS85, VDD80, vdHdV83]. orders [Wam80]. ordinary [Car89, Her83, KLS87, LB88, NF84]. Orthogonal [Gau85, Gro85, Gro86, Mar86, PR87, Pre83, RM89, Van85, Van87a, Van89c]. Orthogonality [MH83, Gro84b, IS87, IN87].

Orthonormal [Bog86, WG81]. oscillating [Lyn85, Tez89]. oscillatory [Ehr88, GL87a, HT87, Sid87, Sug87c]. other [Ano88f, Had87, Roo85, Wal88, vS86]. outer [Fuk84]. Overconvergence [BP88]. overcorrection [Bla88]. Overrelaxation [Dic82b, NSO80, NIS82, Yey88, Yey89]. oversampled [Nat86]. overview [BS89]. Oxford [Goo88b]. oxygen [Gre81b].

P [Ano86d, Bul88c, Bul88b]. P.D.E. [EO82, Had85]. P/b [Bul88c]. packages [Hak87]. Padé [Bak84, BL87a, BGN81, Bul80, BV86, Con87, CMM86, CWW84, Cuy85, Cuy86, De 88, Dra88, Eie84, GV87, GM81, GMS88, Hop82, JW89, JM80, JNT83, Luk80, Luk82, Mag87, McC81a, McC81b, Pas87, Pre83, Sab83, Sid80, Sid81, Wal87]. Padé-like [BGN81]. Padé-type [Con87, Eie84, Pre83, Sab83]. padeons [LM86a]. pages [Ano86d, Ano86e, Ano86f, Rij80, Van89a]. pair [FB89]. pairs [FG86]. paper [McN86b]. papers [Ano88g, Ano89n]. parabolic [BH83, De 84, Pan89, vdHdV83]. paradigms [ACCK88]. paradox [dLM89]. Parallel [BZ89, MD86, PP89, BE89, BPW89, BS89, Kip89, MW87, Paa89, SS89b, Tun87, VSV89]. Parameter [Mur88, EHL89, HM89, LST89, NSO80, NIS82, ST86, Sny88]. parameter-dependent [HM89]. parameter-free [Sny88]. parameters [Bar83, Dun88, SOA82, SV87]. parametric [AEH81, DN88]. parametrized [MM89]. part [Bul84c, Goo88a, Bul84b, Gre81a, Gre81b, GG85, LD87]. Partial [MS83, De 88, DI89, Dic82b, GR81, ITH88, Jan87, KLS87, OP85, Pan89, SGC85, SG88, Tt89, VV88, ZH80]. partitioned [Szu89]. parts [JW86]. passing [BE89]. past [Bre85]. path [FW87, GG88, Gro88]. path-integrals [GG88, Gro88]. paths [Rou83, SCJ89]. PDE [Hak87, SW89].

PDEs [Lip83a]. penalties [Mur87]. pencil [CvD89]. perfect [Rij80]. period [MS89]. periodic [Ana86, BL87b, Bre87b, CR84, CR86, EO82, Gut86a, JW86, JKG83, Kri82, Lem87, Soh89, Twi86]. periodicity [Cha86, KP89]. Peripheral [BH83]. Permanence [AS88]. Perron [McC81a]. perturbation [Cri86, FR80, Kam83, Nii86, Özi89, Vul89]. perturbed [Her83, OS87, SS88]. Petron [MG80]. Phase [FW87, Twi86, ÁKS88, CR84, CR86, CRN86, CR87, Wei88]. Phase-lag [Twi86, CR84, CR86, CRN86, CR87]. phenomena [AR86, DRV89]. Philadelphia [Lev88]. philosophy [Goo88a]. physics [Bog86, Deh81, Gla84]. Picard [LM82]. piecewise [Cad87, Joe85, Pap82]. Placement [BGN81]. planar [CH87a, CH87b, DSV88, MW89]. plane [Col87b, ES85]. planes [Gol86]. plate [AA82]. plates [AN81]. point [Aga83, AU86, ÁKS88, AT80, CK80, CK82, CS86, CS87, Cha87, CSS88, CS88, CMM86, Ehr88, FVD87, GV87, JM80, JNT83, LM82, LM86a, Lin86, Man84a, McC81b, Niz80, PW81a, Sab87, Sid80, dG81]. points [Cal85, DW89, Joe89, Nii86, Oli80c, Rei84, Roo85, Seg84, Sug87b]. Pointwise [Hen82, Pri87]. Poisson [Lin89, PD88, SS89b]. Pole [PW84, Con87, GV87, PWH86]. Pole-type [PW84, PWH86]. poles [GMM89, Ioa86a, JM80, PWH83, Pre87]. pollution [Zla85]. polygamma [BS83]. polygon [TW86]. polygonal [ET86]. polygons [Von88]. Polynomial [Bre89, Col87b, Not88, Col87a, Hou85, KK89, Mac82, MW87, McN88, Miy89, Oli80b, Pet82, PM83, PS86, WG81]. polynomials [Ano84e, Bog86, Gau85, Gro84b, Gro85, Gro86, IN87, Joe85, Kio86, Luk80, Luk82, NF84, PR87, Pre83, Pre88, Rab87b, RM89, Sal83, Van85, Van89c]. polytopes [RWP85]. populations [BDvdM82]. porous [RS86, SR88, Wat81]. positive [BW89, GHLN89, IS88, Kio86, vdV87]. posteriori [Dob88]. potential [Mu89]. Pounds [Lev88]. power [For86a, For86b, Gar86a, Hop82, Lon85, Sie80]. power-boundedness [Gar86a]. powers [Ast82, Mei81]. pp [Ano88f, Ano89f, Bul88a, Bul88c, Bul88b, Goo88b, Goo88a, Lev88, dBH83]. Practical [Ber89a, Iga85, Lau85]. pre [DM84b]. pre-elimination [DM84b]. precedence [Ste85]. precision [AT80, Bar82]. Preconditioned [Kaa88]. preconditioners [AE89]. Preconditioning [Saa88a, Axe85, AP88, DMP87, HO88, vdVD88]. predator [LLM82, LM86b]. predicting [PC80]. predictor [SF86, SF87, Sha81]. Preface [Ano89q, Bre87a, BGPW85, BGPW89, GW87, Had88, Lau87, MR89, MP87, Tre86]. preliminary [WKGC87]. prescribed [GMM89]. presence [Rab87a, SB88a]. presented [Ano89n]. preserving [CDR89, Cos88, LY89, SH87]. Press [Rij80, Bul88c]. pressure [Ten87]. previous [Gar86b]. prey [LLM82, LM86b]. prey-predator [LLM82, LM86b]. Price [Ano88f, Ano89f, Bul88a, Bul88c, Bul88b, Goo88b, Goo88a, Lev88, Van89a, Ano86d, Ano86e, Ano86f, NA88, Rij80]. principal [Ioa81, Ors89]. principle [KB88]. principles [HSM83, TS86]. Pringsheim [Lev86, Lev89]. prism [vS86]. probabilities [Bla87, DG82]. probability [Cam87]. problem [AC84, Akl81, Ami87, ÁKS88, AT80, BP86, Chi89, Chi81, DKW84, EM86, Fab87, FR80, GW81, GR83, GR89a, Hel82, Hen87, KLT89, KKK89, Lon84, MP88, Mur88, Mur87, NNOP86, Nii86, NSIO82, Niz80, OS87, PW81a, Ren88, SR88, Seg84, SB88b, Slo86, Smi87a, Sny87, Ste87, SGC85, SG88, Tun87, Usm80, Wal87, dG81, dLM89]. problems [Aga83, AU86, Aga87, Aga88, AA82, ACG89, Ana86, BB88, BZ89, BS89, Bia89, Bla88, Bog86, BOP84, CK80, CK82, CR84, Cha85, CS86, CR86, CS87, Cha87, CSS88, CS88, Chu83, CW85, De 85, DR87c, DW89, Dyk86, FVD87, GLSZ87, Hak87, Her83, IS85, JKG83, Jai88, JG88, Jel85, KR86, Kri81, Kri82, LM82, Lav84, Lin86, LW88, Man84a, MG89, Mil89, MG80, Mu89, Nak87, NS82, NT86, Ong84b, Özi89, Pap89, PD88, Rei84, RD87, SF84, SF86, Sar88, ST86, Sny88, Son89, SS88, Tt89, Twi86, TT87, Vul89, Weg88, WO80a, WO80b, Zla87, dGvV89, Ano88f]. procedure [Cal85, Col80, MP88]. procedures [BM89, CW85]. Proceedings [Ano89f]. process [GGD81, Paa89, Yam85]. processes [AHCW86, DN87, Hem82, Pre87]. processor [MS88]. Product [ALE84, CM84, Rab89, AL87, Bar83, Ber89b, Die86, Dyk86, Gri82, Kip89, Kle87, KL87]. Prof. [Cuy87b]. profile [Smy85]. program [Abd81, Dun80, Van89a]. programming [GW81]. programs [Fuk84]. progressive [AE88]. project [KY88]. Projection [Mae88, Bre83, Sid88]. projections [GK89, Mar86, Tun87]. proof [Jel86, Tem88]. Propagation [Ado88, SU89, Whe85]. properties [AM84, Bak84, BP88, EM87, Gro86, Ran87]. property [Gro84b]. pseudo [AG88, EM86]. pseudo-random [AG88]. pseudo-spectral [EM86]. pseudolinear [Lav84]. Publishers [Bul88b]. pure [Ano88f, vdHS84].

Q.D. [GM80b]. QD [Cla81]. Quadratic [MPN82, SH87, LM84, RLW89]. quadratically [Dah86, Paa89]. Quadrature [Let87, Mor85, AL87, Ano89k, Ber89a, BF89, Dix86, Ehr88, EM87, GN89, Gri82, Ioa82, Ioa84, IMT87, JV89, Let81, LD87, Ors89, Sol89, SS89a]. qualitative [Smi87a]. quantum [GLSZ87, PE87]. Quasi [Lam85, dLM89, CDR89, FS87, Lin89, MO87]. quasi-direct [Lin89]. quasi-interpolation [CDR89]. Quasi-Monte [Lam85]. quasi-Newton [FS87, MO87]. Quasi-variational [dLM89]. quasilinear [Lav84, Vul89]. queueing [Bla87]. quintic [BP88, PS87]. quotient [McC81b, PA89].

R [Ano84f, Bul88a]. radial [Sha88b, VFD89]. radii [Gre86a, Van87b]. radius [Day84b, Gre86b]. ramp [eMH89]. Random [vS86, AG88, BLS84c, DN87, Han88, Rou83, SU89]. range [CS86]. rank [Hem82, Zie84]. ranking [BDvdM82]. Rapid [May86, EL89a]. rate [RV88]. rates [AL86, DM84a]. ratio [McC83]. Rational [BM87, Rok87, Sot84, CQ82, Cle81, Dun84a, Dun84b, GMM89, GM84, IS87, MR87, Rok86, Wam80]. ray [GM81]. Rayleigh [GR83, Gro84b]. re [PD88]. re-entrant [PD88]. reaction [Dix87, EM86, KP89, MMSY84, Man84b, MS89]. real [CW85, Oli80a]. reciprocal [CV88]. reconsideration [DP86]. reconstruction [DSV88]. recovery [LR89]. rectangle [PKW87, Pap89]. rectangular [Cam80, EO82, GM80a, Mae88, Zie84]. Rectified [BZ84]. Recurrence [Pas87, Kip89, PC80, dB83]. recurrent [Kip89]. recursions [Bul84b, Bul84c]. Recursive [Bre83, Bul84b, Bul84c, Jan87, EKL89, Gro85, RM89, Son88a]. Reduced [BdTL85, VS85]. reducible [Dix86]. reducing [KP82]. Reduction [vdHS86, BV86, DSH89, MSW89, Smy85]. refinement [DR87a, Ruo87]. reflection [Sha88a]. refraction [Puu88a]. region [Ado84b, Cam80, EO82, GM80a]. Regions [CS84b, CH87a, CH87b, De 89, Lin89, Mar88, May86, SO87, Weg86b]. regular [Ano82, Bar82, CvD89, Ham82, JW89]. Reidel [Ano86d]. Reinsch [Oli80b]. related [SH87, dD84]. relations [Bul84b, Bul84c, Del87, Gro84b, Kip89, Pas87, PC80, Van87a, dB83]. relationship [Mur87]. Relaxation [Phi87, Had87].

Relaxed [Ypm86]. relevant [GG83]. reliability [DN85]. reliable [VVH87]. remark [Ioa84]. Remarks [GM80a, Jac88]. reparameterisation [Hoi86]. reports [Ano87j, Ano87k, Ano88k, Ano89l]. representation [DR87c]. require [DM83]. residual [HKD86, PMD87]. residue [VS88]. resolution [Cap88, EO82]. respect [Hop82, KG87a]. restricted [Abd81]. restrictions [KLS87]. result [Nev85]. results [AT80, BR89, GG83, LS88, NA88, Van85, dB83, dB80]. retaining [GGD81]. retarded [Bic82]. Review [Ano86d, Ano86e, Ano86f, Ano88f, Ano89f, Bul88a, Bul88c, Bul88b, Goo88b, Goo88a, Lev88, Rij80, Van89a, Cuy85, KY88, MS83, Nel88]. revisited [BB83, Lev89]. revolution [MC85]. Riccati [Car86]. Richardson [Sug87c]. Riemann [Dic85, Weg88]. right [AU86]. rigid [Gol86]. ring [Gre86a, Gre86b]. risk [DG82]. robot [LST89]. Romberg [Sha84]. root [BLS84c, MW87, PS86]. root-finding [MW87]. rootfinding [KLT89]. roots [Bre89, CM86, Iga85, Kio86]. Rosenbrock [DM82, Day84a, Day86, Hun87]. rotating [Gol86, Wat81]. rotation [Van88]. Rounding [BP86, ITH88]. routine [JV89, KR87]. routines [Ber89a]. Row [GMS88]. rows [BL87a]. Rózsa [Bul88b]. ruin [DG82]. rule [AE88, Bel86, EL89a, Iga85, Sug87a, Tem88, Mor85]. rules [Cuy86, Ioa84, KL87, Let87, Ors89, PE87, Rab87b]. Run [Day84b]. Runge [Ano86f, CQ82, CS84a, CS84b, DP80, DP86, FG86, Heb82, LDS87, PD81, Som87, Sot84, Spi85, VSV89, VDD80, Wam80].

S [Ano84f, Ano86d]. salesman [Akl81].

Salzer [Ano84e]. SAOR

[EL88, aON88, YaOIN85]. Sard [Köh89]. scalar [Car86, FW87]. scale [ITH88, Ong84a]. scattered [Kle87]. scattering [Ber89b, PE87]. scheduling [Ste85]. scheme [AN81, Car89, LLM82, LM86b, Nii86, Oli80b, Pan89, Ten87]. schemes [AEH81, AL86, GK89, Jel85, Jel86, Lin86, Not88, Phi87, Sar88, VS85, Vul89, Wil86, vdHS86]. Schrödinger [VFD89, Xin88]. Schroedinger [Gd80]. Schur [CvD89, Tem88]. Schur-complements [Tem88]. Science [Bul88b, Goo88b, Goo88a, Goo88a]. sciences [Kli88]. search [Rou83]. Second [AEH81, Hal84, KB88, AHCW86, Ana86, Ber86, CM84, CR84, Cha85, CN86, CR86, CS88, Col88, Gre86a, Gro86, Jai88, JG88, Jel86, Joe85, Kri81, Kri82, Lip84, Man84a, Oli80c, Ong84b, PR88, SG89, Twi86, TT87]. second-kind [CM84]. Second-order [Hal84, CS88, Jai88, JG88, Man84a, PR88, SG89, TT87]. Secondary [Chi89]. section [Cam80]. sectionally [Ioa86a]. segments [BLS84b]. seismic [DN87]. selecting [LD86]. selection [BM89, GSB87, Mur88, Oli80c]. self [Hen87, Lip83a]. self-adjoint [Hen87, Lip83a]. semantics [Van89a]. semi [AEH81, CS86, Let87, Lip81, PD88, SL86, TFI88, Wad80]. semi-bandwidths [Lip81]. semi-infinite [CS86, PD88, TFI88, Wad80]. semi-integrals [Let87]. semi-iteration [AEH81]. semi-iterative [SL86]. semiconductor [CPS89, WW89]. semiderivatives [Let87]. semidiscrete [vdHdV83]. semigroups [Cro87]. semilinear [Bud89, Smi87a]. sense [Sid87]. sensitivity [Hop82]. separable [GW81, SF84]. sequence [Ast82, Jan87, Mei81]. sequences [Fis89, GM81, Gro86, Lam85, Léc89, Sab87]. sequential [MD86]. Series [GL84, TBB88, Ado84a, Ano84e, Cuy87a, For86a, For86b, GG83, GG85, HT87, Hop82, KK89, Lon85, McC83, Sal83, Sie80, Soh88, Son88a, Tez89]. set [BdTL85, Bul88c, Von88]. setups [Ste85]. seven [Lip81]. seven-diagonal [Lip81]. several [Fab87, Yam85]. Shape [CDR89, LY89, Cos88, Ioa89, SH87]. Shape-preserving [CDR89, Cos88]. shear [ES85]. shock [How84]. shooting [MG89]. sided [Dob88, Roh87]. sieve [RLW89]. significant [BdTL85]. Silhouettes [Rou87]. simple [DN87, DN88, De 89, Kle87, LST89, Rok86]. simply [Weg88]. simulation [Col80, CPS89, Mam89, Whe85]. Simultaneous [De 88, Ber88, Eri88, FH88, Pet82, PM83]. Sinc [Bia89]. Sinc-type [Bia89]. Singh [Ano86d]. Single [Hun88, Con87, GV87, JG88, Mit87, SG89]. Single-level [Hun88]. single-step [JG88, Mit87]. Singular [BGG88, Cuy86, Ger86, ALE84, BS89, Ber89b, CS87, Cha87, CSS88, Dix87, Had85, HLS88, Ioa82, Ioa84, Ioa86b, Kaa88, Nii86, PJ80, Rae82, Ste87, VVH87, Vul89]. singularities [AE88, GR89b, PW84, PWH86, PD88]. singularity [Han87, Rab87a]. singularly [OS87, SS88]. six [CRN86, ST87]. Sixth [Jai88, CR87]. Sixth-order [Jai88, CR87]. size [Bon87, GSB87, Sha87, Wat83]. small [Dix87]. smaller [Zie84]. smallest [VVH87]. Smíal [Lev88]. Smooth [Usm80, Gwi81, dBH83]. smoothing [Col80, DSV88, FR80, Kle82, PV85]. Societatis [Bul88b]. software [Hak87, HH89, KLT89]. solidification [Mam89]. Solitons [LM86a]. Solution [GHLN89, Sha86, Wam80, eMH89, AMHAEM89, Ado84a, Ami87, AL86, BD87, Ber88, Bic82, BeMM86, BZ84, Bru86, BOP84, Cah81, Cam80, Car86, Car89, CK87, Chi81, DN88, DMP87, Dic87, Dob88, DM84b, EKL89, ED89, EO82, FR80, GR81, GR83, GR89a, Had85, Had87, HLS88, Hak87, Hel82, Hüb86b, Ioa82, Ioa84, KLT89, Kli87a, KLS87, Kri81, Lip81, MS86, MO87, MP88, Mur88, NA88, NSIO82, Ong84a, Ong84b, OP85, PW81a, Pap89, SR88, ST86, Slo86, SU89, SV87, SS89b, TW86, Usm80, VDD80, ZH80, Zla87, vdV87]. Solutions [PF84, Ado84b, AR86, Akl81, BTB88, Bar83, CS87, GL84, Hou85, Jan87, KR86, Man84a, MS89, Mu89, RV88, Sha88a, Sha88b, SB88b, TBB88]. solved [Tem85]. solver [GSB87, Lin89, Lip83b, Wat83]. solvers [DM83, SW89]. Solving [Hof87, Mar86, SvdV89, SS88, AU86, AM84, Ber89b, Bla88, Chu83, DKW84, FH88, Hal82, HM85, Hüb86a, Jab80, JY88, JH89, Kaa88, Léc89, Lip83a, Lip84, LW88, Luk80, Luk82, Mae88, MS86, MD86, NNOP86, PE87, Rei84, Rei86, SF84, SF86, SF87, VFD89, Xin88, vdV87]. Some [BE89, De 89, EM87, FVD87, Kip89, KKN^{+}88, Let81, Lon84, MW87, McN86b, Sid80, Van85, WKGC87, DM83, Del87, DP86, Gla84, Lyn85, PM83, PS86, Rei84, Sid81, SB88a, TBB88]. space [AE89, Bar83, De 84, EHL89, FW87, Hen87, Kle87, Lip83a, vdHS86]. spaced [BP80, Far86, FK88]. spaces [Cro87, DM84a, Gri82, Wan85, Yam88]. sparse [ACG89, Duf89, GHLN89, Lip81, Lip83b, Lip84, MS86, Ong84a, Ypm87]. sparsity [Zla87]. spatial [NA88, Puu88b]. speakers [Ano85b]. special [DW89, FVD87, FL85]. Spectral [Xin88, Col80, Day84b, DMP87, EM86, EO82, Hof89, Phi87, PD88]. spectrum [Maa85]. speeding [Dru84]. sphere [Kea87, Mur82, RD81]. Spherical [Fre84]. spirals [MW89]. Spline [Ano86d, PE87, CSS88, CS88, CDR89, DR87b, Fre84, Gut86a, KR86, McN86a, MPN82, Met83, Met84, PW81a, PW81b, PS87, Roh87, Soh89, SS88, Usm80]. Spline-Gauss [PE87]. spline-on-spline [PS87]. splines [ALE84, BP80, BP88, Cos88, Die86, DSV88, Kle87, Lee86, Neu81, ST87, SH87, Soh88, Wan85, dBH83]. split [SU89]. split-step [SU89]. splitting [De 84, Dic85, Dic89]. splittings [Not88]. spontaneous [Smi87b]. Springer [Ano86e].

SQP [TFI88]. square [PS86]. squares [Bul84a, Far86, NS82, PP89, SF84, Sar88, VV88, Van89b]. SSOR [HO88]. Stability [AM84, Dun84b, Jel85, Sho89, Tor89, vdHS84, Bir81, CQ82, CS84a, CS84b, KLS87, Mam89, PC80, Puu88b, SB88a, SB88b, Ano86f]. stabilized [dGvV89]. stable [Ana86, Bic82, CN86, CRN86, Col88, DM82, DI89, JKG83, Kho86, MS89, Tho82a, Tho82b]. standard [Hak87]. Starting [BGS88, Wat83]. state [Bla87, DW89, LLM82, LM86b, Mur87, Van89a]. states [Gre87]. stationary [AL86, NSO80]. statistical [BDvdM82]. statistics [RR81]. steady [DW89, Dic85, Dic87, Dic89, How84, IS85, LLM82, LM86b]. steady-state [LLM82, LM86b]. Stefan [ÁKS88, GR89a]. step [Ana86, Bel84, Car86, Cha85, CN86, Cha86, CRN86, CG89, Eri88, GSB87, HO88, JKG83, Jai88, JG88, JV89, Mit87, Sha87, SU89, Twi86, Wat83]. Stepsize [BM89, FB89, KLS87, PO87]. Stetter [Ano86e]. Stieltjes [GM81, Pre88]. Stieltjes- [Pre88]. Stiff [Ano86f, Chu83, Roc88b, SB84, Sha86, Sot84, SG89, Ypm86, dG81]. Stiffly [Kho86, Tho82b]. Stochastic [Wei88]. Stokes [Dic87, Ten87]. stopping [Iga85]. storage [VDD80]. strategy [HM89]. streamline [ET86]. strength [DN88]. stress [DN88]. stress-strength [DN88]. stresses [ES85]. strong [Hen87]. Structural [Puu88b]. structure [AG88, DN85]. structured [Sha86]. Studies [Goo88b, Gre87, MMSY84, Man84b]. study [Col80, De 84, Dix87, PF84, Smi87b, Van80, Wat81]. studying [Zla85]. Sturm [Bar83]. Subdivision [Dah86]. subgradient [Kiw86, Tun87]. Subject [Ano86k, Ano87p]. subproper [Not88]. subsidence [SR88]. subspace [VVH87, vdV87]. Successive [Bre87b, Ioa88, LM82, Mar86]. Suitability [LDS87]. suitable [Sot84]. sulphur [Zla85]. sum [DD87]. summability [Dru84, Sid87]. summation [Lon85]. sums [Gro84b]. Sung [Ano88f]. Superconvergence [Nak87, KN87]. Superstable [Cha85, JG88, Jai88]. Surface [PV85, DSV88, Kea87, LR89]. surfaces [CDR89, MC85]. survey [Bru82, BV86, CW85, Had87, KJ88, Zla87]. SVD [EL89b, HC85, Kam83]. swallowtail [Cal86]. switching [KTS88]. symmetric [BS89, BGG88, CK82, CG89, CH87a, CH87b, CW85, Lip81, vdV87]. Symmetrical [GM84]. symmetries [DW89]. symmetries-with [DW89]. symmetrizer [VS88]. Symmetry [Bud89]. symposium [Ano89f]. Synergetics [Hak88, Zab88]. Synthesis [KPC80]. system [AZ88, BV86, Hüb86b, KP89, Wat81, Xin88]. systems [AL86, Bla87, Bul84b, Bul84c, Car89, Cas88, CK87, CG89, Dic82a, Dic82b, EKL89, FH88, GHLN89, Gro85, Had85, Had87, Hal82, HM85, Heb82, Hof87, HS82, ITH88, Jab80, JV89, JY88, JH89, Kaa88, KLT89, KKN^{+}88, Kli88, KPC80, KB88, Lip81, Lip84, Luk80, Luk82, Mae88, MS86, Man84a, Mar86, Mie89, aON88, Ong84a, Saa88a, SvdV89, SW88, Sid88, SB88b, Sot84, VDD80, VCL87, Yey88, Ypm86]. Szego [Ano89k, GN89, KT86].

T [Bul88c, Goo88a]. table [Bul80, CMM86, DC81, Dra88, JW89, McC81a, Sha84]. tables [BDvdM82]. tails [Lyn85]. Tainter [AMHAEM89]. talks [Ano85c, Ano87l, Ano87m, Ano89m]. tangent [Yam88]. tanh [AE88]. tau [Cri86, OP85]. Taylor [GR83]. Technical [Ano89f, Bul88b]. technique [DR87a, Mam89, Özi89, WW89]. techniques [BD87, BV86, Dic87, FH88, GL87b, Had87, Hal82, Mag87, PWH86, PKW87, Saa88a]. temperature [Mur82]. temperatures [MP88]. tension [SS88]. tensor [Die86, Dyk86, Kle87]. term [BH83, Cri86, PC80, dB83]. terminal [Mur87]. terminating [McN88]. terms [Tem81]. test [dB80]. Texas [Bul88b]. th [BLS84c, Bre89, Heb82]. their [DC81, GR89b, Hop82, JM80, Kip89, PWH83]. Theodorsen [Hüb86a]. theorem [Bon87, GS86, Lev86, Lev89]. theorems [BL87a, GMS88]. Theory [Ano86d, Ano86e, Gro85, AA82, Bak84, Ber89b, BV86, Cuy85, Fre84, Gau85, Kos88, Puu88a, Rij80, Wei88, Yey88, dB80, Bul88c]. thermal [RD87, ZN84]. thermoelasticity [Son88b]. these [Pre87]. theses [Ano87j, Ano87k, Ano88k, Ano89l]. Thiele [CV88, Sie83]. Thiele-type [Sie83]. third [Aga87, Oli80c]. third-order [Aga87]. three [AE89, Ehr88, Gro88, Lin86, Lip83a, PC80, dBH83]. three-point [Ehr88, Lin86]. three-space [Lip83a]. three-term [PC80]. time [Day84b, De 85, Kle82, KG87b, Mur87, Nev85, NSOS88, ZH80]. time-dependent [NSOS88, ZH80]. Toeplitz [Bul84b, Bul84c, Mag87]. Tokyo [Goo88b]. tool [HO86]. TOR [KJ88]. toroidal [Cam80]. total [VV88, Van89b]. tracing [Kli87b]. tracking [CC86]. Tract [Goo88a]. tracts [Von88]. traffic [Maa85, Puu88a, dLM89]. transcendental [PF84]. transfer [HS82, PF84]. transform [ZH80]. transformation [AE88, Lem87, Lem89, LW88, Mor85, Oli80b, Özi89]. transformations [BM87, Lam85, Tem85]. transformed [IN87]. transforms [Cam80, DL82, GG85, HH84, Neu81]. transistor [Sha88b]. transitions [Wei88]. transitive [vS86]. transmission [Bul84a]. trapezoidal [EL89a, Sug87a]. traveller [Oue82]. travelling [Akl81]. Treating [SG89]. treatment [Bru82, Bru89, Da 87, Her83, How84, PWH86]. triangular [KN87, vS86]. tridiagonal [EKL89, Gd80]. tridiagonalization [Deh81]. trigamma [Ron86]. trigonometric [Hel82, Hen82, NF84, Soh89]. Troesch [Chi81]. Truncation [For86b, SW88].

Tucker [Van89a]. turbulent [Whe85]. turning [Nii86]. Two [CRN86, Dob88, ES85, JNT83, McC81b, Ana86, ÁKS88, AT80, Bog86, BH83, CK80, CK82, Cha85, CS86, CN86, Cha86, CS87, Cha87, CSS88, CS88, CMM86, De 84, DR87c, DC81, Eri88, FVD87, Gar86b, GV87, Gro88, HS82, Jai88, JM80, KR87, LM82, LB88, Man84a, Niz80, PW81a, RS86, Sid80, Sie83, ST86, Twi86, dG81]. Two-dimensional [ES85, KR87]. two-parameter [ST86]. two-phase [ÁKS88]. Two-point [JNT83, McC81b, AT80, CK80, CK82, CS86, CS87, Cha87, CSS88, CS88, CMM86, FVD87, GV87, JM80, LM82, Man84a, PW81a, Sid80, dG81]. Two-sided [Dob88]. Two-step [CRN86, Ana86, Cha85, CN86, Cha86, Eri88, Jai88, Twi86]. two-variable [Sie83]. TWODQD [KR87]. twofold [Gro84b]. type [Ano89k, Bar83, Bia89, BC82, Cap88, CR84, CR86, Con87, Cub84, Eie84, FS87, GN89, GV87, GS86, Ioa86b, Kho86, Nii86, PW84, PWH86, Pre83, Pre88, Roc88b, Sab83, Sie83, Tho82a, Tho82b, vdVD88]. types [GG85, Saa88b, Van89a].

U.S [Ano88f, Van89a]. U.S.$ [Ano89f]. U.S.A [Bul88b]. UK [Lev88]. ultimate [Axe85]. unbounded [Bru89, SO87]. unconstrained [FG89, SF86]. underdetermined [MD86]. unequal [Dic82a, Fab87]. unequally [BP80]. unified [Yam85]. Uniform [HO86, Cha87, Gut86a, How84, OS87]. unilateral [NT86]. Unique [Man84a]. uniqueness [Aga87]. unitary [Gra86]. University [Ano89f, Bul88c, Bul88b]. Unsteady [RS86, AN81, Dic82a]. upon [KTS88]. Upper [DG82, ISK87, JH89]. upwind [Jel86, Pan89]. Upwinding [MG80]. Urabe [Mit87]. US$ [Ano86d, Ano86f]. use [FG89, KLT89, Lip84, MW89, PWH83, Sha81]. using [BLS84b, DN87, Die86, DSV88, FB89, Ioa89, Joe85, Kip89, VS88, vdV87]. Uzawa [LQ86].

V [Goo88a, Van89a]. validation [Rou87]. value [Aga83, AC84, AU86, Aga87, Aga88, AA82, Ana86, AT80, BZ89, BS89, Bia89, BOP84, BGG88, Cam87, CK80, CK82, CR84, Cha85, CS86, CR86, CS87, Cha87, CSS88, CS88, Chu83, FVD87, Ioa81, JKG83, Jai88, JG88, KR86, KTS88, Kri81, Kri82, LM82, Lav84, Lin86, LW88, Man84a, MG89, Mu89, Niz80, Ors89, PW81a, Sha88a, SB88b, Sny87, Sug87a, SS88, Tt89, Twi86, TT87, Usm80, WO80a, WO80b, dG81]. valued [GMS88, Nev85]. values [FG89, Rok86, VVH87]. variable [BB88, BLS84c, Ehr88, HS82, KTS88, LW88, Mor85, Oli80a, PO87, Sie83, dB80]. variables [Rou83]. variants [Say88]. variation [Özi89]. variational [CSR87, HM89, MM89, Sny87, ZH89, dLM89]. vary [KG87a]. VCH [Bul88a]. Vector [SS89b, Van87a, GMS88, Hof87, Kam83, PMD87, RLW89, SvdV89, SB88a, Van89c, Van87a]. vector-valued [GMS88]. Vectorizable [AE89, SW89]. vectors [AG88]. Verlagsgesellschaft [Bul88a]. version [Bel86, GW81, KKK89, Mit87]. versions [Nev85]. versus [RD87]. vertex [vS86]. vertex-transitive [vS86]. Verwer [Ano86f]. very [CW85]. via [Dob88, KT86]. viability [AS88]. view [FH88, LM86a]. viscous [AN81]. visualization [Zab88]. Vol [Ano86j, Ano89p, DP88, Van89a, Ano88f, Bul88b]. Volterra [Bru82, Bru86, Bru89, Cah81, Dix86, Dix87, HLS88, PJ80]. Volume [Ano89d, Ano83a, Ano84a, Ano84b, Ano86a, Ano86b, Ano86c, Ano87e, Ano87b, Ano87c, Ano87d, Ano88d, Ano88e, Ano89b, Ano89c, Ano89e]. volumes [Ano84g, Ano85a]. vortex [Zab88]. vorticity [Mu89]. vs [Niz84, Sid88].

W [Ano86d, Bul88a]. waiting [Bla87]. walks [vS86]. water [KKN^{+}88]. Watson [Ano86d]. wave [FW87, SU89, Whe85]. wavefunctions [Bar82]. waves [How84, RC86]. weakly [Dix87, HLS88]. weight [Ano89k, GN89, Gro86, Mag87]. Weighted [Dun82]. weights [CH87a, KL87]. Weinheim [Bul88a]. Werner [Cuy87b]. where [Jac88]. which [DM83, ES85, Sug87a]. width [dG81]. Wien [Ano86e]. Wiener [GGD81]. wind [DN85]. without [Nii86]. works [Cuy87b].

York [Ano88f, Goo88b, Rij80].

zero [IS88, Kli87b]. zeros [Bul84a, Dör85, DC81, FL85, GL83, ISK87, IN87, Miy89, Pet82, PM83, PS86, Rab87b, Van85, Van89c]. Zucker [Van89a].

References

Agarwall:1982:BVP

[AA82] Ravi P. Agarwall and G. Akrivis.

Boundary value problems occurring in plate deflection theory. Journal of Computational and Applied Mathematics, 8(3):145–154, September 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X82900353.

Abdelmalek:1981:CPD

[Abd81] N. N. Abdelmalek. Computer program for the discrete linear restricted Chebyshev approximation. Journal of Computational and Applied Mathematics, 7(2):141–150, June 1981. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X81900486.

Agarwal:1984:IMF

[AC84] Ravi P. Agarwal and Y. M. Chow. Iterative methods for a fourth order boundary value problem. Journal of Computational and Applied Mathematics, 10(2):203–217, April 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278490058X.

Andersson:1988:EPA

[ACCK88] Åke E. Andersson, J. Casti, Peter Leth Christiansen, and Anders Karlqvist. Emerging paradigms in applied mathematical modelling — an introduction. Journal of Computational and Applied Mathematics, 22(2–3):143–181, June 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903974.

Allgower:1989:LSC

[ACG89] E. L. Allgower, C.-S. Chien, and K. Georg. Large sparse continuation problems. Journal of Computational and Applied Mathematics, 26(1–2):3–21, June 1989. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789901441.

Adomian:1984:CSS

[Ado84a] G. Adomian. Convergent series solution of nonlinear equations. Journal of Computational and Applied Mathematics, 11(2):225–230, October 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900220.

Adomian:1984:CRD

[Ado84b] G. Adomian. On the convergence region for decomposition solutions. Journal of Computational and Applied Mathematics, 11(3): 379–380, December 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/037704278490013X.

Adomian:1988:PDD

[Ado88] G. Adomian. Propagation in dissipative or dispersive media. Journal of Computational and Applied Mathematics, 23(3):395–396, September 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278890009X.

Adomian:1989:CCD

[Ado89] G. Adomian. Comments on a “counterexample” to decomposition. Journal of Computational and Applied Mathematics, 26(3):375–376, July 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789903105.  See

[Nel88].

Aihie:1988:TTD

[AE88] V. U. Aihie and G. A. Evans. The tanh transformation for double integrals with singularities — a progressive rule. Journal of Computational and Applied Mathematics, 24(3):349–354, December 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788902968.

Axelsson:1989:VPE

[AE89] O. Axelsson and V. Eijkhout. Vectorizable preconditioners for elliptic difference equations in three space dimensions. Journal of Computational and Applied Mathematics, 27(1–2):299–321, September 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789903725.

Avdelas:1981:SOC

[AEH81] G. Avdelas, D. J. Evans, and A. Hadjidimos. Second order Chebyshev semiiteration in connection with pparametric E.A.D.I. schemes. Journal of Computational and Applied Mathematics, 7 (4):261–266, December 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0771050X8190053X.

Afflerbach:1988:LSP

[AG88] Lothar Afflerbach and Holger Grothe. The lattice structure of pseudo-random vectors generated by matrix generators. Journal of Computational and Applied Mathematics, 23(1):127–131, July 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278890338X.

Agarwal:1983:CAC

[Aga83] Ravi P. Agarwal. Contraction and approximate contraction with an application to multi-point boundary value problems. Journal of Computational and Applied Mathematics, 9(4):315–325, December 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
0377042783900031.

Agarwal:1987:EUI

[Aga87] Ravi P. Agarwal. Existenceuniqueness and iterative methods for third-order boundary value problems. Journal of Computational and Applied Mathematics, 17(3): 271–289, March 1987. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042787901051.

Agarwal:1988:MCI

[Aga88] Ravi P. Agarwal. Monotone convergence of iterative methods for (n, p) and (p, n) boundary value problems. Journal of Computational and Applied Mathematics, 21(2):223–230, February 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788902701.

Al-Hassani:1986:SOA

[AHCW86] S. T. S. Al-Hassani, M. Can, and E. J. Watson. A second order approximation to nonlinear circuit equations as applied to high energy electrical discharge processes. Journal of Computational and Applied Mathematics, 15(2):175–189, June 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786900257.

Akl:1981:ENO

[Akl81] Selim G. Akl. On the expected number of optimal and near-optimal solutions to the Euclidean travelling salesman problem. Journal of Computational and Applied Mathematics, 7(4):287–288, December 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X81900577.

Ang:1988:NTP

[ÁKS88] ang iñh Áng, Fritz Keinert, and Frank Stenger. A nonlinear two-phase Stefan problem with melting point gradient: a constructive approach. Journal of Computational and Applied Mathematics, 23(2):245–255, August 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278890283X.

Avdelas:1986:MIC

[AL86] G. Avdelas and A. Leontitsis. A method for the improvement of the convergence rates of stationary iterative schemes for the solution of complex linear systems. Journal of Computational and Applied Mathematics, 15(1):1–11, May 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786902347.

Alaylioglu:1987:PQA

[AL87] Ayse Alaylioğlu and D. S. Lubinsky. A product quadrature algorithm by Hermite interpolation. Journal of Computational and Applied Mathematics, 17(3):237–269, March 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704278790104X.

Alaylioglu:1983:NEF

[Ala83] Ayse Alaylioğlu. Numerical evaluation of finite Fourier integrals. Journal of Computational and Applied Mathematics, 9(4):305–313, December 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704278390002X.

Alaylioglu:1984:PIL

[ALE84] Ayse Alaylioğlu, D. S. Lubinsky, and D. Eyre. Product integration of logarithmic singular integrands based on cubic splines. Journal of Computational and Applied Mathematics, 11(3):353–366, December 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900104.

Al-Mutib:1984:SPN

[AM84] A. N. Al-Mutib. Stability properties of numerical methods for solving delay differential equations. Journal of Computational and Applied Mathematics, 10(1):71– 79, February 1984. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042784900712.

Abdel-Malek:1989:ASG

[AMHAEM89] M. N. Abdel-Malek, S. N.

Hanna, and M. B. Abd-El-Malek. Approximate solution for gravity flow under a tainter gate. Journal of Computational and Applied Mathematics, 26(3):271–279, July 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789902999.

Amini:1987:IMB

[Ami87] S. Amini. An iterative method for the boundary element solution of the exterior acoustic problem. Journal of Computational and Applied Mathematics, 20(??):109–117, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787901282.

Aziz:1981:NSU

[AN81] A. Aziz and T. Y. Na. A numerical scheme for unsteady flow of a viscous fluid between elliptic plates. Journal of Computational and Applied Mathematics, 7(2):115–119, June 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X81900449.

Ananthakrishnaiah:1986:CTS

[Ana86] U. Ananthakrishnaiah. A class of two-step P -stable methods for the accurate integration of second order periodic initial value problems. Journal of Computational and Applied Mathematics, 14(3):455–459, March 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786900804.

Anonymous:1982:EOC

[Ano82] Anonymous. Erratum: “Optimal cocircuits in regular matroids and applications” [J. Comput. Appl. Math. 8(3) 171–177 (1982)]. Journal of Computational and Applied Mathematics, 8(4): 308, December 1982. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0771050X82900572.

See [Ham82].

Anonymous:1983:AIV

[Ano83a] Anonymous. Author index to volume 9 (1983). Journal of Computational and Applied Mathematics, 9(4): 391–392, December 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042783900109.

Anonymous:1983:EB

[Ano83b] Anonymous. Editorial board.

Journal of Computational and Applied Mathematics, 9 (1):vi, March 1983. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042783900225.

Anonymous:1983:LAJ

[Ano83c] Anonymous. List of algorithms: (journal CAM, 1975– 1982). Journal of Computational and Applied Mathematics, 9(1):91–92, March 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042783900316.

Anonymous:1983:LBJ

[Ano83d] Anonymous. List of bibliographies: (journal CAM, 1975–1982). Journal of Computational and Applied Mathematics, 9(1):93, March 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042783900328.

Anonymous:1984:AIVa

[Ano84a] Anonymous. Author index to volume 10 (1984). Journal of Computational and Applied Mathematics, 10(3):387–388, June 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042784900475.

Anonymous:1984:AIVb

[Ano84b] Anonymous. Author index to volume 11 (1984). Journal of Computational and Applied Mathematics, 11(3): 385–386, December 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042784900153.

Anonymous:1984:EBa

[Ano84c] Anonymous. Editorial Board. Journal of Computational and Applied Mathematics, 10(1): ii, February 1984. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042784900621.

Anonymous:1984:EBb

[Ano84d] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 11 (1):ii, August 1984. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042784900268.

Anonymous:1984:EJCb

[Ano84e]  Anonymous.  Errata: J.

Comput. Appl. Math. 9: H. E. Salzer, Note on the Dočev–Grosswald asymptotic series for generalized Bessel polynomials, (1983) 131–

135. Journal of Computational and Applied Mathematics, 10(1):133, February 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900773.  See

[Sal83].

Anonymous:1984:EJCa

[Ano84f] Anonymous. Errata: J. Compute. Appl. Math. 9: N. S. Kambo, R. K. Jain and R. Goel, A fourth order method for y′′ = f(x, y, y′), (1983) 81–90. Journal of Computational and Applied Mathematics, 10(1):133, February 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900761.  See

[KJG83].

Anonymous:1984:MIV

[Ano84g] Anonymous. Master index to volumes 1–10. Journal of Computational and Applied Mathematics, 10(3):389–405, June 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900487.

Anonymous:1985:AIV

[Ano85a]  Anonymous.  Author in-

dex to volumes 12 and 13 (1985). Journal of Computational and Applied Mathematics, 12–13(??):667–668, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042785900573.

Anonymous:1985:LIS

[Ano85b] Anonymous. List of invited speakers. Journal of Computational and Applied Mathematics, 12–13(??):viii, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042785900020.

Anonymous:1985:LT

[Ano85c] Anonymous. List of talks.

Journal of Computational and Applied Mathematics, 12–13(??):ix–xiv, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042785900032.

Anonymous:1986:AIVa

[Ano86a]  Anonymous.  Author in-

dex to volume 14 (1986). Journal of Computational and Applied Mathematics, 14(3):473–474, March 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042786900865.

Anonymous:1986:AIVb

[Ano86b] Anonymous. Author index to volume 15 (1986). Journal of Computational and Applied Mathematics, 15(3):401–402, July 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786902323.

Anonymous:1986:AIVc

[Ano86c] Anonymous. Author index to volume 16 (1986). Journal of Computational and Applied Mathematics, 16(3): 397–398, November 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042786900105.

Anonymous:1986:BRA

[Ano86d] Anonymous. Book review: Approximation Theory and Spline Functions: edited by S. P. Singh, J. W. H. Burry and B. Watson, (Reidel, Dordrecht, 1983) 485 pages; price: 69,50 US$. Journal of Computational and Applied Mathematics, 14(3):471–472, March 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786900853.

Anonymous:1986:BRD

[Ano86e] Anonymous. Book review: Defect Correction Methods. Theory and Applications: edited by K. Böhmer and H. J. Stetter, (Springer, Wien, 1984) 243 pages; price: 68.00 DM. Journal of Computational and Applied Mathematics, 14(3):471, March 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786900841.

Anonymous:1986:BRS

[Ano86f]  Anonymous.  Book re-

view: Stability of Runge– Kutta Methods for Stiff Nonlinear Differential Equations: by K. Dekker and J. G. Verwer (North-Holland, Amsterdam, 1984) 308 pages; price: 36.50 US$. Journal of Computational and Applied Mathematics, 14(3):471, March 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278690083X.

Anonymous:1986:EBa

[Ano86g] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 14(1–2):ifc, February 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042786901263.

Anonymous:1986:EBb

[Ano86h] Anonymous. Editorial board.

Journal of Computational and Applied Mathematics, 15 (1):iv, May 1986. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042786902335.

Anonymous:1986:EBc

[Ano86i] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 16(1):vi, September 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042786901688.

Anonymous:1986:NCA

[Ano86j] Anonymous. Newsletter on computational and applied mathematics: Vol. 2, nr. 3, November 1986 applied mathematics. Journal of Computational and Applied Mathematics, 16(3):P1, 1–14, November 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786900117.

Anonymous:1986:SI

[Ano86k]  Anonymous.  Subject in-

dex. Journal of Computational and Applied Mathematics, 14(1–2):267–269, February 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786901433.

Anonymous:1987:AC

[Ano87a] Anonymous. Announcements of conferences. Journal of Computational and Applied Mathematics, 17(3):N4– N16, March 1987. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042787901191.

Anonymous:1987:AIVa

[Ano87b] Anonymous. Author index to volume 17 (1987). Journal of Computational and Applied Mathematics, 17(3): 379, March 1987. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042787901166.

Anonymous:1987:AIVb

[Ano87c] Anonymous. Author index to volume 18 (1987). Journal of Computational and Applied Mathematics, 18(3):403–404, June 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900112.

Anonymous:1987:AIVc

[Ano87d]  Anonymous.  Author in-

dex to volume 19 (1987). Journal of Computational and Applied Mathematics, 19(3):401–402, September 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787902093.

Anonymous:1987:AIV

[Ano87e]  Anonymous.  Author in-

dex to volume 20 (1987). Journal of Computational and Applied Mathematics, 20(??):423–424, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787901622.

Anonymous:1987:EBa

[Ano87f] Anonymous. Editorial board. Journal of Computational and Applied Mathematics, 17(1–2):ii, February 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/037704278790032X.

Anonymous:1987:EBb

[Ano87g] Anonymous. Editorial Board. Journal of Computational and Applied Mathematics, 18 (1):viii, April 1987. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042787900501.

Anonymous:1987:EBc

[Ano87h] Anonymous. Editorial Board. Journal of Computational and Applied Mathematics, 19 (1 (supplement 1)):ii, July 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787901646.

Anonymous:1987:EBd

[Ano87i] Anonymous. Editorial Board. Journal of Computational and Applied Mathematics, 20 (??):ii–iii, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042787901208.

Anonymous:1987:IRDa

[Ano87j]  Anonymous.  Institutional

reports and doctoral theses. Journal of Computational and Applied Mathematics, 17(3):N2–N4, March 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278790118X.

Anonymous:1987:IRDb

[Ano87k]  Anonymous.  Institutional

reports and doctoral theses. Journal of Computational and Applied Mathematics, 20(??):n1–n11, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901634.

Anonymous:1987:LIT

[Ano87l] Anonymous. List of invited talks. Journal of Computational and Applied Mathematics, 20(??):417, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        0377042787901609.

Anonymous:1987:LT

[Ano87m] Anonymous. List of talks. Journal of Computational and Applied Mathematics, 20(??):418–421, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787901610.

Anonymous:1987:NCAa

[Ano87n] Anonymous. Newsletter on computational and applied mathematics. Journal of Computational and Applied Mathematics, 17(3):N1, March 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787901178.

Anonymous:1987:NCAb

[Ano87o]  Anonymous.  Newsletter

on computational and applied mathematics. Journal of Computational and Applied Mathematics, 19(1 (supplement 1)):N1–N9, July 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787901865.

Anonymous:1987:SI

[Ano87p]  Anonymous.  Subject in-

dex. Journal of Computational and Applied Mathematics, 17(1–2):235–236, February 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900495.

Anonymous:1988:AC

[Ano88a] Anonymous. Announcements of conferences. Journal of Computational and Applied Mathematics, 22(2–3):N5–N8, June 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788904189.

Anonymous:1988:AIa

[Ano88b]  Anonymous.  Author in-

dex. Journal of Computational and Applied Mathematics, 23(3):397–398, September 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788900106.

Anonymous:1988:AIb

[Ano88c]  Anonymous.  Author in-

dex. Journal of Computational and Applied Mathematics, 24(3):411–412, December 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903020.

Anonymous:1988:AIVa

[Ano88d]  Anonymous.  Author in-

dex to volume 21 (1988). Journal of Computational and Applied Mathematics, 21(3):387–388, March 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042788903238.

Anonymous:1988:AIVb

[Ano88e] Anonymous. Author index to volume 22 (1988). Journal of Computational and Applied Mathematics, 22(2–3):381, June 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788904153.

Anonymous:1988:BRO

[Ano88f] Anonymous. Book review: Operator methods for optimal control problems: Sung J. Lee (ed.) Lecture Notes in pure and applied mathematics vol. 108, Marcel dekker, Inc., New York, Basel, 1987, 344 pp., Price: $79.50 (U.S. and Canada), $95.25 (all other countries), ISBN: 0-8247-7811-1. Journal of Computational and Applied Mathematics, 24(3):N3, December 1988. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278890307X.

Anonymous:1988:CP

[Ano88g] Anonymous. Call for papers. Journal of Computational and Applied Mathematics, 24(3):N5–N6, December 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903111.

Anonymous:1988:EBa

[Ano88h] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 21(1):ii–iii, January 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042788903822.

Anonymous:1988:EBb

[Ano88i] Anonymous. Editorial board.

Journal of Computational and Applied Mathematics, 22(1): iv–v, April 1988. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042788902841.

Anonymous:1988:EBc

[Ano88j] Anonymous. Editorial board.

Journal of Computational and Applied Mathematics, 23 (1):ii–iii, July 1988. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042788903251.

Anonymous:1988:IRD

[Ano88k]  Anonymous.  Institutional

reports and doctoral theses. Journal of Computational and Applied Mathematics, 22(2–3):N2–N4, June 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788904177.

Anonymous:1988:ICC

[Ano88l]  Anonymous.  International

course on computational geometry. Journal of Computational and Applied Mathematics, 24(3):N6–N8, December 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903123.

Anonymous:1988:NCAa

[Ano88m] Anonymous. Newsletter on computational and applied mathematics. Journal of Computational and Applied Mathematics, 21(3):N1–N11, March 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278890324X.

Anonymous:1988:NCAb

[Ano88n] Anonymous. Newsletter on computational and applied mathematics. Journal of Computational and Applied Mathematics, 22(2–3):N1, June 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788904165.

Anonymous:1989:AC

[Ano89a] Anonymous. Announcement of conferences. Journal of Computational and Applied Mathematics, 26(3):N6–N11, July 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903154.

Anonymous:1989:AIVa

[Ano89b] Anonymous. Author index to volume 25 (1989). Journal of Computational and Applied Mathematics, 25(3):387–388, May 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789900423.

Anonymous:1989:AIVb

[Ano89c] Anonymous. Author index to volume 26 (1989). Journal of Computational and Applied Mathematics, 26(3):377–378, July 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903117.

Anonymous:1989:AIVc

[Ano89d] Anonymous. Author index to volume 27 (1989). Journal of Computational and Applied Mathematics, 27(3): 431–432, November 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042789900290.

Anonymous:1989:AIVd

[Ano89e]  Anonymous.  Author in-

dex to volume 28 (1989). Journal of Computational and Applied Mathematics, 28(??):419–420, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903543.

Anonymous:1989:BRP

[Ano89f] Anonymous. Book review: Proceedings of the international symposium on numerical analysis: O. Celebi and B. Karasö (Eds.) 1989 Middle East Technical University, 350 pp., Price: ≈ 5 U.S.$ + mailing costs, ISBN 975-429-007-5. Journal of Computational and Applied Mathematics, 26(3):N2, July 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903130.

Anonymous:1989:EBa

[Ano89g] Anonymous. Editorial board.

Journal of Computational and Applied Mathematics, 25(1):ii–iii, January 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042789900691.

Anonymous:1989:EBb

[Ano89h] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 26(1– 2):ii–iii, June 1989. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042789901428.

Anonymous:1989:EBc

[Ano89i] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 27 (1–2):ii–iii, September 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042789903567.

Anonymous:1989:EBd

[Ano89j] Anonymous. Editorial Board. Journal of Computational and Applied Mathematics, 28 (??):ii–iii, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042789903166.

Anonymous:1989:EGK

[Ano89k] Anonymous. Erratum: Gauss– Kronrod quadrature formulae for weight functions of Bernstein–Szegő type: Journal of Computational and Applied Mathematics 25 (2) (1989) 199–224. Journal of Computational and Applied Mathematics, 27(3): 429, November 1989. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042789900289.

See [GN89].

Anonymous:1989:IRD

[Ano89l]  Anonymous.  Institutional

reports and doctoral theses. Journal of Computational and Applied Mathematics, 26(3):N3–N5, July 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903142.

Anonymous:1989:LIT

[Ano89m] Anonymous. List of invited talks. Journal of Computational and Applied Mathematics, 28(??):413, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278990352X.

Anonymous:1989:LPP

[Ano89n] Anonymous. List of papers presented at the congress. Journal of Computational and Applied Mathematics, 28(??):414–417, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903531.

Anonymous:1989:NCAa

[Ano89o] Anonymous. Newsletter on computational and applied mathematics. Journal of Computational and Applied Mathematics, 25(2):N1–N18, February 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789900502.

Anonymous:1989:NCAb

[Ano89p]  Anonymous.  Newsletter

on computational and applied mathematics: Vol. 5, nr. 3, November 1989.

Journal of Computational and Applied Mathematics, 28(??):N1–N13, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903555.

Anonymous:1989:P

[Ano89q] Anonymous. Preface. Journal of Computational and Applied Mathematics, 27(1– 2):ix–x, September 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042789903579.

Ohsaki:1988:ASM

[aON88] Ichi ann Ohsaki and Hiroshi Niki. The accelerated SAOR method for large linear systems. Journal of Computational and Applied Mathematics, 24(1–2):277–291, November 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903597.

Axelsson:1988:BPD

[AP88] O. Axelsson and B. Polman. Block preconditioning and domain decomposition methods. II. Journal of Computational and Applied Mathematics, 24(1–2):55–72, November 1988. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903433.

Adomian:1986:NCP

[AR86] G. Adomian and R. Rach. The noisy convergence phenomena in decomposition method solutions. Journal of Computational and Applied Mathematics, 15(3):379–381, July 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786902281.

Adomian:1988:EID

[AR88] G. Adomian and R. Rach.

Evaluation of integrals by decomposition. Journal of Computational and Applied Mathematics, 23(1):99–101, July 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903330.

Aubin:1988:PV

[AS88] Jean-Pierre Aubin and Karl Sigmund. Permanence and viability. Journal of Computational and Applied Mathematics, 22(2–3):203–209, June 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        0377042788904001.

Astin:1982:CSP

[Ast82] J. Astin. Comment “On the sequence of the powers of a given matrix”. Journal of Computational and Applied Mathematics, 8(1):69, March 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X82900109.  See

[Mei81].

Anglesio:1980:HPR

[AT80] J. D. Anglesio and B. A. Troesch. High precision results for a two-point boundary value problem. Journal of Computational and Applied Mathematics, 6(2):99–103, June 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X80900030.

Agarwal:1986:IMS

[AU86] Ravi P. Agarwal and Riaz A. Usmani. Iterative methods for solving right focal point boundary value problems. Journal of Computational and Applied Mathematics, 14(3):371–390, March 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
0377042786900749.

Axelsson:1985:IBM

[Axe85] O. Axelsson. Incomplete block matrix factorization preconditioning methods. The ultimate answer? Journal of Computational and Applied Mathematics, 12–13(??):3–18, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042785900044.

Andersson:1988:DCD

[AZ88] Åke E. Andersson and Wei-Bin Zhang. Decision centralization and decentralization in a dynamic economic system. Journal of Computational and Applied Mathematics, 22(2–3):319–337, June 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788904116.

Baker:1984:IPH

[Bak84] George A. Baker, Jr. Invariance properties in Hermite– Padé approximation theory. Journal of Computational and Applied Mathematics, 11(1):49–56, August 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
0377042784900311.

Barnett:1982:HPE

[Bar82] A. R. Barnett. High-precision evaluation of the regular and irregular Coulomb wavefunctions. Journal of Computational and Applied Mathematics, 8(1):29–33, March 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X82900043.

Barakat:1983:PSS

[Bar83] M. F. A. Barakat. On the product solutions of Sturm– Liouville equations with many parameters in a space of the type L_{2}. Journal of Computational and Applied Mathematics, 9(4):327–331, December 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042783900043.

Balas:1983:BMR

[BB83] Egon Balas and Christian Bergthaller. Benders’s method revisited. Journal of Computational and Applied Mathematics, 9(1):3–12, March 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042783900249.

Balasundaram:1988:MFE

[BB88] S. Balasundaram and P. K.

Bhattacharyya. A mixed finite-element method for fourth-order elliptic problems with variable coefficients. Journal of Computational and Applied Mathematics, 22(1):1–24, April 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788902853.

Bignami:1982:MCM

[BC82] A. Bignami and E. Cupini. A Monte Carlo method for finite difference equations of elliptic type in a multiregion domain. Journal of Computational and Applied Mathematics, 8(2):87–92, June 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X82900602.

Baker:1987:FTN

[BD87] C. T. H. Baker and M. S.

Derakhshan. FFT techniques in the numerical solution of convolution equations. Journal of Computational and Applied Mathematics, 20(??):5–24, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        0377042787901221.

Bozzini:1985:RSL

[BdTL85] M. Bozzini, F. de Tisi, and L. Lenarduzzi. Reduced set of leads for electrocardiography, maintaining significant information. Journal of Computational and Applied Mathematics, 12–13(??):195–206, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042785900160.

Beirlant:1982:CSR

[BDvdM82] J. Beirlant, E. J. Dudewicz, and E. C. van der Meulen. Complete statistical ranking of populations, with tables and applications. Journal of Computational and Applied Mathematics, 8(3): 187–201, September 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0771050X82900419.

Babuska:1989:SAP

[BE89] I. Babuska and H. C. Elman. Some aspects of parallel implementation of the finite-element method on message passing architectures. Journal of Computational and Applied Mathematics, 27(1–2):157–187, September 1989. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903658.

Bellen:1984:OSC

[Bel84] A. Bellen. One-step collocation for delay differential equations. Journal of Computational and Applied Mathematics, 10(3):275–283, June 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900396.

Belward:1986:EVF

[Bel86]  J. A. Belward.  An ex-

ponential version of Filon’s rule. Journal of Computational and Applied Mathematics, 14(3):461–466, March 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786900816.

Boutros:1986:LSF

[BeMM86] Y. Z. Boutros, M. B. Abd el Malek, and S. Z. Masoud. Linearized solution of a flow over a nonuniform bottom. Journal of Computational and Applied Mathematics, 16(1):105–116, September 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        0377042786901779.

Berrut:1986:FIE

[Ber86] Jean-Paul Berrut. A Fredholm integral equation of the second kind for conformal mapping. Journal of Computational and Applied Mathematics, 14(1–2):99–110, February 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786901329.

Berg:1988:ISS

[Ber88] Lothar Berg. Iterative solution of simultaneous equations. Journal of Computational and Applied Mathematics, 24(1–2):257–264, November 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788903573.

Berntsen:1989:PEE

[Ber89a] Jarle Berntsen. Practical error estimation in adaptive multidimensional quadrature routines. Journal of Computational and Applied Mathematics, 25(3):327–340, May 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789900368.

Bertram:1989:PIM

[Ber89b] B. Bertram. On the product integration method for solving singular integral equations in scattering theory. Journal of Computational and Applied Mathematics, 25(1):79–92, January 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042789900770.

Brass:1989:EBQ

[BF89] Helmut Brass and Klaus-Jürgen Förster. Error bounds for quadrature formulas near Gaussian quadrature. Journal of Computational and Applied Mathematics, 28(??):145–154, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903269.

Bunse-Gerstner:1988:SVD

[BGG88] Angelika Bunse-Gerstner and William B. Gragg. Singular value decompositions of complex symmetric matrices. Journal of Computational and Applied Mathematics, 21(1):41–54, January 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278890386X.

Baumel:1981:PCP

[BGN81] R. T. Baumel, J. L. Gammel, and J. Nuttall. Placement of cuts in Padé-like approximation. Journal of Computational and Applied Mathematics, 7(2):135–140, June 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X81900474.

Broeckx:1985:P

[BGPW85] F. Broeckx, M. Goovaerts, R. Piessens, and L. Wuytack. Preface. Journal of Computational and Applied Mathematics, 12–13(??):vii, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042785900019.

Broeckx:1989:P

[BGPW89] F. Broeckx, M. J. Goovaerts, R. Piessens, and L. Wuytack. Preface. Journal of Computational and Applied Mathematics, 28(??):1, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789903178.

Brankin:1988:SBA

[BGS88] R. W. Brankin, I. Gladwell, and L. F. Shampine. Starting BDF and Adams codes at optimal order. Journal of Computational and Applied Mathematics, 21(3):357–368, March 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903196.

Baratella:1981:ABF

[BGV81] P. Baratella, M. Garetto, and G. Vinardi. Approximation of the Bessel function J_{ν}(x) by numerical integration. Journal of Computational and Applied Mathematics, 7(2):87–91, June 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X81900401.

Broom:1983:PAH

[BH83] M. A. Broom and T. R. Hopkins. Peripheral and ADI hopscotch methods for two dimensional parabolic and elliptic equations with a mixed term. Journal of Computational and Applied Mathematics, 9(3):239–245, September 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042783900171.

Bialecki:1989:STA

[Bia89] Bernard Bialecki. Sinc-type approximations in H^{1}-norm with applications to boundary value problems. Journal of Computational and Applied Mathematics, 25(3):289–303, May 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789900332.

Bickart:1982:SSI

[Bic82] Theodore A. Bickart. F - stable and F [α, β]-stable integration/interpolation methods in the solution of retarded differential-difference equations. Journal of Computational and Applied Mathematics, 8(4):279–284, December 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X82900523.

Biringen:1981:NNS

[Bir81]  Sedat Biringen.  A note

on the numerical stability of the convection–diffusion equation. Journal of Computational and Applied Mathematics, 7(1):17–20, March 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X81900024.

Baker:1987:CTR

[BL87a] George A. Baker, Jr. and D. S. Lubinsky. Convergence theorems for rows of differential and algebraic Hermite–Padé approximations. Journal of Computational and Applied Mathematics, 18(1):29–52, April 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900549.

Brezinski:1987:LCL

[BL87b] C. Brezinski and A. Lembarki.

The linear convergence of limit periodic continued fractions. Journal of Computational and Applied Mathematics, 19(1 (supplement 1)):75–77, July 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901749.

Blanc:1987:NMC

[Bla87] J. P. C. Blanc. On a numerical method for calculating state probabilities for queueing systems with more than one waiting line. Journal of Computational and Applied Mathematics, 20(??):119–125, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901294.

Blaheta:1988:MMO

[Bla88] Radim Blaheta. A multilevel method with overcorrection by aggregation for solving discrete elliptic problems. Journal of Computational and Applied Mathematics, 24(1–2):227–239, November 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704278890355X.

Bowman:1984:BCI

[BLS84a] K. O. Bowman, H. K. Lam, and L. R. Shenton. Bounds for certain integrals. Journal of Computational and Applied Mathematics, 10(2):245–254, April 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704278490061X.

Bowman:1984:BIU

[BLS84b] K. O. Bowman, H. K. Lam, and L. R. Shenton. Bounds for integrals using segments. Journal of Computational and Applied Mathematics, 11(1):57–68, August 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042784900323.

Bowman:1984:BER

[BLS84c] K. O. Bowman, H. K. Lam, and L. R. Shenton. Bounds for the expectation of the kth root of a random variable. Journal of Computational and Applied Mathematics, 10(2):229–244, April 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900608.

Baart:1987:RCI

[BM87] M. L. Baart and R. J. Y.

McLeod. Rational curves and isoparametric transformations. Journal of Computational and Applied Mathematics, 18(1):121–128, April 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900604.

Bank:1989:SSC

[BM89] Randolph E. Bank and Hans D. Mittelmann. Stepsize selection in continuation procedures and damped Newton’s method. Journal of Computational and Applied Mathematics, 26(1–2):67–77, June 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789901489.

Bogdanova:1986:OPO

[Bog86] Nina Bogdanova Bogdanova.

Orthonormal polynomials in one and two dimensions: application to calibration problems in high energy physics. Journal of Computational and Applied Mathematics, 14(3):345–351, March 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786900713.

Bonckaert:1987:SDL

[Bon87] Patrick Bonckaert. On the size of the domain of linearization in Hartman’s theorem. Journal of Computational and Applied Mathematics, 19(2):279–282, August 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901968.

Bui:1984:RAM

[BOP84] T. D. Bui, A. K. Oppenheim, and D. T. Pratt. Recent advances in methods for numerical solution of O.D.E. initial value problems. Journal of Computational and Applied Mathematics, 11(3):283–296, December 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900037.

Behforooz:1980:ECI

[BP80] G. H. Behforooz and N. Papamichael. End conditions for interpolatory cubic splines with unequally spaced knots. Journal of Computational and Applied Mathematics, 6(1):59–65, March 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X80900182.

Barnard:1986:RCG

[BP86] Roger W. Barnard and Kent Pearce. Rounding corners of gearlike domains and the omitted area problem. Journal of Computational and Applied Mathematics, 14(1–2):217–226, February 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786901408.

Behforooz:1988:OPQ

[BP88] G. Behforooz and N. Papamichael. Overconvergence properties of quintic interpolatory splines. Journal of Computational and Applied Mathematics, 24(3): 337–347, December 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042788902956.

Barry:1989:MPI

[BPW89] J. M. Barry, J. P. Pollard, and E. W. Wachspress. A method of parallel iteration. Journal of Computational and Applied Mathematics, 28(??):119–127, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789903245.

Beernaert:1989:ERG

[BR89] Lutgarde Beernaert and Dirk Roose. Efficiency results for Gaussian elimination on the iPSC/2 hypercube. Journal of Computational and Applied Mathematics, 28(??):129–143, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789903257.

Brezinski:1983:RIE

[Bre83]  Claude Brezinski.  Recur-

sive interpolation, extrapolation and projection. Journal of Computational and Applied Mathematics, 9(4): 369–376, December 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042783900080.

Brezinski:1985:CAM

[Bre85] Claude Brezinski. Convergence acceleration methods: the past decade. Journal of Computational and Applied Mathematics, 12–13(??): 19–36, May 1985. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042785900056.

Brezinski:1987:P

[Bre87a] C. Brezinski. Preface. Journal of Computational and Applied Mathematics, 19 (1 (supplement 1)):1, July 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901658.

Brezinski:1987:SML

[Bre87b] C. Brezinski. Successive modifications of limit periodic continued fractions. Journal of Computational and Applied Mathematics, 19(1 (supplement 1)):67–74, July 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901737.

Breuer:1989:PIO

[Bre89]  Shlomo Breuer.  Polyno-

mial iterations of r-th order for the extraction of kth roots. Journal of Computational and Applied Mathematics, 25(1):61–68, January 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789900757.

Brunner:1982:SRA

[Bru82] Hermann Brunner. A survey of recent advances in the numerical treatment of Volterra integral and integrodifferential equations. Journal of Computational and Applied Mathematics, 8(3): 213–229, September 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0771050X82900444.

Brunner:1986:HOM

[Bru86] Hermann Brunner. High-order methods for the numerical solution of Volterra integrodifferential equations. Journal of Computational and Applied Mathematics, 15(3):301–309, July 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786902219.

Brunner:1989:NTV

[Bru89] Hermann Brunner. The numerical treatment of Volterra integro-differential equations with unbounded delay. Journal of Computational and Applied Mathematics, 28 (??):5–23, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/037704278990318X.

Bowman:1983:CFP

[BS83] K. O. Bowman and L. R.

Shenton. Continued fractions and the polygamma functions. Journal of Computational and Applied Mathematics, 9(1):29–39, March 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042783900262.

Berry:1989:OPA

[BS89] Michael Berry and Ahmed Sameh. An overview of parallel algorithms for the singular value and symmetric eigenvalue problems. Journal of Computational and Applied Mathematics, 27(1–2):191–213, September 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278990366X.

Balachandran:1988:ASN

[BTB88] K. Balachandran, E. Thandapani, and G. Balasubramanian. Approximate solutions of a nonlinear differential equation. Journal of Computational and Applied Mathematics, 24(3): 399–402, December 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042788903007.

Budd:1989:SBS

[Bud89] Chris Budd. Symmetry breaking and semilinear elliptic equations. Journal of Computational and Applied Mathematics, 26(1–2):79–96, June 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789901490.

Bultheel:1980:DAC

[Bul80] A. Bultheel. Division algorithms for continued fractions and the Padé table. Journal of Computational and Applied Mathematics, 6(4):259–266, December 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X80900340.

Bultheel:1984:ICL

[Bul84a]  A. Bultheel.  On the

ill conditioning of locating transmission zeros in least squares ARMA filtering. Journal of Computational and Applied Mathematics, 11(1):103–118, August 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900360.

Bultheel:1984:RRBa

[Bul84b] A. Bultheel. Recursive relations for block Hankel and Toeplitz systems. Part I: Direct recursions. Journal of Computational and Applied Mathematics, 10(3):301–328, June 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900426.

Bultheel:1984:RRBb

[Bul84c] A. Bultheel. Recursive relations for block Hankel and Toeplitz systems part II: Dual recursions. Journal of Computational and Applied Mathematics, 10(3):329–354, June 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900438.

Bultheel:1988:BRD

[Bul88a]  A. Bultheel.  Book re-

view: Digitale filter: Hamming R. W. Informationstechnologie VCH Verlagsgesellschaft, Weinheim, F. R. Germany, 1987, 275 pp., Price:

DM 64, ISBN 3-527-26463-9. Journal of Computational and Applied Mathematics, 24(3):N3, December 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903068.

Bultheel:1988:BRNb

[Bul88b]  A. Bultheel.  Book re-

view: Numerical methods: D. Greenspan (University of Texas at Arlington, U.S.A.), P. Rózsa (Technical University of Budapest, Hungary), Colloquia Mathematica Societatis Janos Bolyai, vol 50 North-Holland, Elsevier Science Publishers, Amsterdam, 1988, 688 pp., Price: US $223.75 / Dfl. 425.00, ISBN 0-444-70209-1. Journal of Computational and Applied Mathematics, 24(3):N2, December 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903056.

Bultheel:1988:BRNa

[Bul88c] Adhemar Bultheel. Book review: Notes on logic and set theory: P. T. Johnstone, Cambridge University press, Cambridge, 1987, pp., Price: $34.5 H/b (ISBN: 0-521-33502-7), $12.95 P/b (ISBN: 0-521-33692-9). Journal of Computational and Applied Mathematics, 24(3):N2, December 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903044.

Bultheel:1986:PTM

[BV86] A. Bultheel and M. Van Barel.

Padé techniques for model reduction in linear system theory: a survey. Journal of Computational and Applied Mathematics, 14(3):401–438, March 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786900762.

Beauwens:1989:CAP

[BW89] Robert Beauwens and Renaud Wilmet. Conditioning analysis of positive definite matrices by approximate factorizations. Journal of Computational and Applied Mathematics, 26(3):257–269, July 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789902987.

Breuer:1984:RAS

[BZ84] Shlomo Breuer and Gideon Zwas. Rectified approximations for the solution of nonlinear equations. Journal of Computational and Applied Mathematics, 11 (2):251–258, October 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042784900244.

Bellen:1989:PAI

[BZ89] Alfredo Bellen and Marino Zennaro. Parallel algorithms for initial-value problems for difference and differential equations. Journal of Computational and Applied Mathematics, 25(3):341–350, May 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704278990037X.

Cadete:1987:NPB

[Cad87] M. O. Rodrigues Cadete. A note on piecewise blending function interpolation applied to networks of curves in R^{3}. Journal of Computational and Applied Mathematics, 17(3):291–298, March 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787901063.

Cahlon:1981:NSN

[Cah81] Baruch Cahlon. Numerical solution of non-linear Volterra integral equations. Journal of Computational and Applied Mathematics, 7(2):121–128, June 1981. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X81900450.

Caluwaerts:1985:APC

[Cal85]  R. Caluwaerts.  An au-

tomatic procedure for the calculation of bifurcation points of integral equations. Journal of Computational and Applied Mathematics, 12–13(??):207–215, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042785900172.

Caluwaerts:1986:NDM

[Cal86] R. Caluwaerts. A new direct method for the numerical calculation of a swallowtail. Journal of Computational and Applied Mathematics, 15(3):395–396, July 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278690230X.

Campbell:1980:SFF

[Cam80] J. B. Campbell. The solution by fast Fourier transforms of Laplace’s equation in a toroidal region with a rectangular cross-section. Journal of Computational and Applied Mathematics, 6(2):89–92, June 1980. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X80900017.

Campbell:1987:AVM

[Cam87] L. L. Campbell. Asymptotic value of a multidimensional normal probability integral. Journal of Computational and Applied Mathematics, 19(3):287–292, September 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787901981.

Capobianco:1988:MNR

[Cap88] Maria Rosaria Capobianco. A method for the numerical resolution of Abel-type integral equations of the first kind. Journal of Computational and Applied Mathematics, 23(3):281–304, September 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788900039.

Carroll:1986:EFO

[Car86]  John Carroll.  Exponen-

tially fitted one-step methods for the numerical solution of the scalar Riccati equation. Journal of Computational and Applied Mathematics, 16(1):9–25, September 1986. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786901706.

Carroll:1989:CIS

[Car89] John Carroll. A composite integration scheme for the numerical solution of systems of ordinary differential equations. Journal of Computational and Applied Mathematics, 25(1):1–13, January 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789900708.

Casti:1988:ECS

[Cas88] J. Casti. Engineering control systems and computing in the 90s. Journal of Computational and Applied Mathematics, 22(2–3):301–314, June 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788904098.

Chen:1986:DNO

[CC86] Guanrong Chen and Charles K.

Chui. Design of near-optimal linear digital tracking filters with colored input. Journal of Computational and Applied Mathematics, 15(3):353–370, July 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786902268.

Chui:1989:SPQ

[CDR89] Charles K. Chui, Harvey Diamond, and Louise Raphael. Shape-preserving quasi-interpolation and interpolation by box spline surfaces. Journal of Computational and Applied Mathematics, 25(2):169– 198, February 1989. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042789900460.

Chronopoulos:1989:SIM

[CG89] A. T. Chronopoulos and C. W. Gear. s-step iterative methods for symmetric linear systems. Journal of Computational and Applied Mathematics, 25(2):153–168, February 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789900459.

Cuninghame-Green:1984:EOO

[CGB84] R. A. Cuninghame-Green and R. E. Burkard. Eigenfunctions and optimal orbits. Journal of Computational and Applied Mathematics, 11(1):83–94, August 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900347.

Cools:1987:ACK

[CH87a] Ronald Cools and Ann Haegemans. Automatic computation of knots and weights of cubature formulae for circular symmetric planar regions. Journal of Computational and Applied Mathematics, 20(??):153–158, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901324.

Cools:1987:CFS

[CH87b] Ronald Cools and Ann Haegemans. Construction of fully symmetric cubature formulae of degree 4k − 3 for fully symmetric planar regions. Journal of Computational and Applied Mathematics, 17(1–2):173–180, February 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900458.

Cloot:1988:AIK

[CH88] A. H. J. Cloot and B. M.

Herbst. Analytical instability of the Klein–Gordon equation. Journal of Computational and Applied Mathematics, 21(1):17–26, January 1988. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788903846.

Chawla:1985:STS

[Cha85] M. M. Chawla. Superstable two-step methods for the numerical integration of general second order initial value problems. Journal of Computational and Applied Mathematics, 12–13(??):217–220, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042785900184.

Chawla:1986:NCE

[Cha86] M. M. Chawla. A new class of explicit two-step fourth order methods for y′′ = f(t, y) with extended intervals of periodicity. Journal of Computational and Applied Mathematics, 14(3):467–470, March 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786900828.

Chawla:1987:FOF

[Cha87] M. M. Chawla. A fourth-order finite difference method based on uniform mesh for singular two-point boundary-value problems. Journal of Computational and Applied Mathematics, 17(3):359–364, March 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901129.

Chen:1984:AEG

[Che84] T. H. C. Chen. Asymptotic estimates for the generalized Fourier coefficients. Journal of Computational and Applied Mathematics, 10(1):1–5, February 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042784900633.

Chin:1981:MAS

[Chi81] R. C. Y. Chin. On the matched asymptotic solution of the Troesch problem. Journal of Computational and Applied Mathematics, 7(3):181–185, September 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X81900164.

Chien:1989:SBB

[Chi89] C.-S. Chien. Secondary bifurcations in the buckling problem. Journal of Computational and Applied Mathematics, 25(3):277–287, May 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789900320.

Chu:1983:AMM

[Chu83] Moody T. Chu. An automatic multistep method for solving stiff initial value problems. Journal of Computational and Applied Mathematics, 9(3):229–238, September 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278390016X.

Chawla:1980:FDM

[CK80] M. M. Chawla and C. P. Katti. Finite difference methods for a class of twopoint boundary value problems with mixed boundary conditions. Journal of Computational and Applied Mathematics, 6(3):189–196, September 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X8090025X.

Chawla:1982:NSF

[CK82] M. M. Chawla and C. P. Katti. A new symmetric five-diagonal finite difference method for computing eigenvalues of fourth-order twopoint boundary value problems. Journal of Computational and Applied Mathematics, 8(2):135–136, June 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X82900699.

Chang:1987:EFC

[CK87] Stephen F. Chang and William J.

Kennedy. Error-free computer solution of certain systems of linear equations. Journal of Computational and Applied Mathematics, 18(3):279–287, June 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900021.

Claessens:1981:GQA

[Cla81] Guido Claessens. A generalization of the QD algorithm. Journal of Computational and Applied Mathematics, 7(4):237–247, December 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X81900516.

Clegg:1981:NMC

[Cle81] David Benjamin Clegg. On Newton’s method with a class of rational functions. Journal of Computational and Applied Mathematics, 7(2):93–100, June 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        0771050X81900413.

Cameron:1984:PIM

[CM84] R. F. Cameron and S. McKee. Product integration methods for second-kind Abel integral equations. Journal of Computational and Applied Mathematics, 11(1):1–10, August 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704278490027X.

Chawla:1986:CLM

[CM86] M. M. Chawla and S. Mc-Kee. Convergence of linear multistep methods with multiple roots. Journal of Computational and Applied Mathematics, 14(3):451–454, March 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786900798.

Cooper:1986:NNT

[CMM86] S. Clement Cooper, Arne Magnus, and J. H. Mc-Cabe. On the non-normal two-point Padé table. Journal of Computational and Applied Mathematics, 16(3): 371–380, November 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042786900075.

Chawla:1986:FTS

[CN86] M. M. Chawla and Beny Neta.

Families of two-step fourth order P -stable methods for second order differential equations. Journal of Computational and Applied Mathematics, 15(2):213–223, June 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786900282.

Cole:1980:FES

[Col80] F. Cole. Forecasting by exponential smoothing, the Box and Jenkins procedure and spectral analysis. A simulation study. Journal of Computational and Applied Mathematics, 6(1):53–58, March 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X80900170.

Coleman:1987:CPA

[Col87a] John P. Coleman. Complex polynomial approximation by the Lanczos τmethod: Dawson’s integral. Journal of Computational and Applied Mathematics, 20(??):137–151, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
0377042787901312.

Coleman:1987:PAC

[Col87b] John P. Coleman. Polynomial approximations in the complex plane. Journal of Computational and Applied Mathematics, 18(2):193–211, May 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900161.

Coleman:1988:CCS

[Col88] John P. Coleman. Characterisation of a class of P - stable methods for differential equations of second order. Journal of Computational and Applied Mathematics, 22(1):137–141, April 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788902919.

Concepcion:1987:APT

[Con87] Concepción González Concepción. On the A-acceptability of Padé-type approximants to the exponential with a single pole. Journal of Computational and Applied Mathematics, 19(1 (supplement 1)):133–140, July 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
0377042787901816.

Costantini:1988:ACS

[Cos88] Paolo Costantini. An algorithm for computing shapepreserving interpolating splines of arbitrary degree. Journal of Computational and Applied Mathematics, 22(1): 89–136, April 1988. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042788902907.

Coughran:1989:CMS

[CPS89] W. M. Coughran, Jr., M. R.

Pinto, and R. K. Smith. Continuation methods in semiconductor device simulation. Journal of Computational and Applied Mathematics, 26(1– 2):47–65, June 1989. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic). URL http://www.
math.utah.edu/pub/bibnet/
authors/c/coughran-william-
m.bib;   http://www.sciencedirect.
com/science/article/pii/
0377042789901477.

Calvo:1982:SRR

[CQ82] M. Calvo and M. Mar Quemada. On the stability of rational Runge Kutta methods. Journal of Computational and Applied Mathematics, 8(4):289–292, December 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X82900547.

Chawla:1984:NTM

[CR84] M. M. Chawla and P. S. Rao. A Noumerov-type method with minimal phase-lag for the integration of second order periodic initial-value problems. Journal of Computational and Applied Mathematics, 11(3):277–281, December 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042784900025.

Chawla:1986:NTM

[CR86] M. M. Chawla and P. S. Rao. A Noumerov-type method with minimal phase-lag for the integration of second order periodic initial-value problems. II: explicit method. Journal of Computational and Applied Mathematics, 15(3):329–337, July 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786902244.

Chawla:1987:ESO

[CR87] M. M. Chawla and P. S. Rao. An explicit sixthorder method with phaselag of order eight for y′′ = f(t, y). Journal of Computational and Applied Mathematics, 17(3):365–368, March 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901130.

Crisci:1986:TMP

[Cri86] M. R. Crisci. The tau method with perturbation term depending on the differential operator. Journal of Computational and Applied Mathematics, 15(1):123–136, May 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786902451.

Chawla:1986:TSF

[CRN86] M. M. Chawla, P. S. Rao, and Beny Neta. Two-step fourthorder P -stable methods with phase-lag of order six for y′′ = f(t, y). Journal of Computational and Applied Mathematics, 16(2):233–236, October 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786900944.

Crouzeix:1987:MAS

[Cro87] Michel Crouzeix. On multistep approximation of semigroups in Banach spaces. Journal of Computational and Applied Mathematics, 20 (??):25–35, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042787901233.

Chawla:1984:ASE

[CS84a] M. M. Chawla and S. R. Sharma. Absolute stability of explicit Runge–Kutta– Nyström methods for y′′ = f(x, y, y′). Journal of Computational and Applied Mathematics, 10(2):163–168, April 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042784900530.

Chawla:1984:RAS

[CS84b] M. M. Chawla and R. Subramanian. Regions of absolute stability of explicit Runge–Kutta–Nyström methods for y′′ = f(x, y, y′). Journal of Computational and Applied Mathematics, 11 (2):259–266, October 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042784900256.

Chawla:1986:FDM

[CS86] M. M. Chawla and P. N. Shivakumar. A finite difference method for a class of two-point boundary value problems over a semi-infinite range. Journal of Computational and Applied Mathematics, 15(2):251–256, June 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786900312.

Chawla:1987:ESC

[CS87] M. M. Chawla and P. N.

Shivakumar. On the existence of solutions of a class of singular nonlinear twopoint boundary value problems. Journal of Computational and Applied Mathematics, 19(3):379–388, September 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787902068.

Chawla:1988:NFO

[CS88] M. M. Chawla and R. Subramanian. A new fourthorder cubic spline method for second-order nonlinear twopoint boundary-value problems. Journal of Computational and Applied Mathematics, 23(1):1–10, July 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903263.

Castillo:1987:MAV

[CSR87] JoséE. Castillo, Stanly Steinberg, and Patrick J. Roache. Mathematical aspects of variational grid generation II.

Journal of Computational and Applied Mathematics, 20(??):127–135, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787901300.

Chawla:1988:FOS

[CSS88] M. M. Chawla, R. Subramanian, and H. L. Sathi. A fourth-order spline method for singular two-point boundaryvalue problems. Journal of Computational and Applied Mathematics, 21(2):189– 202, February 1988. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042788902671.

Cubillos:1984:IOG

[Cub84] Pedro O. Cubillos. Integral operators with Green’s function type kernel. Journal of Computational and Applied Mathematics, 10(1):25–31, February 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042784900657.

Cuyt:1985:RMP

[Cuy85] Annie A. M. Cuyt. A review of multivariate Padé approximation theory. Journal of Computational and Applied Mathematics, 12–13(??):221–232, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042785900196.

Cuyt:1986:SRC

[Cuy86] A. Cuyt. Singular rules for the calculation of non-normal multivariate Padé approximants. Journal of Computational and Applied Mathematics, 14(3):289–301, March 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786900671.

Cuyt:1987:EMS

[Cuy87a] A. Cuyt. The Euler–Minding series for branched continued fractions. Journal of Computational and Applied Mathematics, 17(3):369–373, March 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901142.

Cuyt:1987:BWP

[Cuy87b] Annie Cuyt. A bibliography of the works of Prof. Dr. H. Werner. Journal of Computational and Applied Mathematics, 19(1 (supplement 1)):3–8, July 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        037704278790166X.

Cuyt:1988:MRD

[CV88] Annie Cuyt and Brigitte Verdonk. Multivariate reciprocal differences for branched Thiele continued fraction expansions. Journal of Computational and Applied Mathematics, 21(2):145–160, February 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788902646.

Charlier:1989:JLA

[CvD89] J.-P. Charlier and P. van Dooren. A Jacobi-like algorithm for computing the generalized Schur form of a regular pencil. Journal of Computational and Applied Mathematics, 27(1– 2):17–36, September 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042789903592.

Cullum:1985:SLP

[CW85] Jane Cullum and Ralph A. Willoughby. A survey of Lanczos procedures for very large real ‘symmetric’ eigenvalue problems. Journal of Computational and Applied Mathematics, 12–13(??):37–60, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042785900068.

Cuyt:1984:CMP

[CWW84] Annie Cuyt, L. Wuytack, and H. Werner. On the continuity of the multivariate Padé operator. Journal of Computational and Applied Mathematics, 11(1):95–102, August 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900359.

DaSilva:1987:NTD

[Da 87] M. R. Da Silva. Numerical treatment of differential equations with the τmethod. Journal of Computational and Applied Mathematics, 20(??):167–173, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901348.

Dahmen:1986:SAC

[Dah86]  Wolfgang Dahmen.  Sub-

division algorithms converge quadratically. Journal of Computational and Applied Mathematics, 16(2):145–158, October 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/0377042786900889.

Day:1984:MCR

[Day84a] J. D. Day. On minimum configuration Rosenbrock methods. Journal of Computational and Applied Mathematics, 11(3):381–383, December 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042784900141.

Day:1984:RTE

[Day84b] J. D. Day. Run time estimation of the spectral radius of Jacobians. Journal of Computational and Applied Mathematics, 11(3): 315–323, December 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042784900062.

Day:1986:MCF

[Day86] J. D. Day. A minimum configuration fourth-order nonautonomous explicit Rosenbrock method for nonstiff differential equations. Journal of Computational and Applied Mathematics, 16(2):247–250, October 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786900981.

denBroeder:1980:BNC

[dB80]  G. den Broeder.  Be-

tween Newton and Cauchy: the diagonal variable metric method; theory and test results. Journal of Computational and Applied Mathematics, 6(4):267–273, December 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X80900352.

deBruin:1983:NCR

[dB83] Marcel G. de Bruin. New convergence results for continued fractions generated by four-term recurrence relations. Journal of Computational and Applied Mathematics, 9(3):271–278, September 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042783900195.

deBoor:1983:BBS

[dBH83] C. de Boor and K. Höllig. Bivariate box splines and smooth pp functions on a three direction mesh. Journal of Computational and Applied Mathematics, 9(1):13–28, March 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042783900250.

Drachman:1981:TTH

[DC81] B. Drachman and C. I.

Chuang. A table of two hundred zeros of the derivative of the modified Bessel function K_{n}(z) and a graph of their distribution. Journal of Computational and Applied Mathematics, 7(3):167–171, September 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X81900140.

deDoelder:1984:CIR

[dD84] P. J. de Doelder. On the Clausen integral Cl_{2}(Θ) and a related integral. Journal of Computational and Applied Mathematics, 11(3): 325–330, December 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042784900074.

DeGroen:1987:FSE

[DD87] P. De Groen and B. De Moor. The fit of a sum of exponentials to noisy data. Journal of Computational and Applied Mathematics, 20(??):175–187, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704278790135X.

DeVries:1984:CSA

[De 84] H. B. De Vries. A comparative study of ADI splitting methods for parabolic equations in two space dimensions. Journal of Computational and Applied Mathematics, 10(2):179–193, April 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900566.

DeMey:1985:ABE

[De 85] G. De Mey. The auxiliary boundary element method for time dependent problems. Journal of Computational and Applied Mathematics, 12–13(??):239–245, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042785900214.

DeBruin:1988:SPP

[De 88] Marcel G. De Bruin. Simultaneous partial Padé approximants. Journal of Computational and Applied Mathematics, 21(3):343–355, March 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903184.

DeOliveira:1989:SSC

[De 89] Paula De Oliveira. Some simple characterizations of contractivity regions. Journal of Computational and Applied Mathematics, 25(1):27– 32, January 1989. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042789900721.

Dehesa:1981:LMT

[Deh81]  Jesus S. Dehesa.  Lanc-

zos method of tridiagonalization, Jacobi matrices and physics. Journal of Computational and Applied Mathematics, 7(4):249–259, December 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X81900528.

Delillo:1987:SRA

[Del87]  Thomas K. Delillo.  On

some relations among numerical conformal mapping methods. Journal of Computational and Applied Mathematics, 19(3):363–377, September 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787902056.

Doedel:1989:NCH

[DF89] Eusebius J. Doedel and Mark J. Friedman. Numerical computation of heteroclinic orbits. Journal of Computational and Applied Mathematics, 26(1–2):155–170, June 1989. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789901532.

deGroen:1981:FEM

[dG81] P. P. N. de Groen. A finite element method with a large mesh-width for a stiff twopoint boundary value problem. Journal of Computational and Applied Mathematics, 7(1):3–15, March 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X81900012.

DeVylder:1982:UBR

[DG82] F. De Vylder and M. J.

Goovaerts. Upper bounds for ruin probabilities in a new general risk model, by the martingales method. Journal of Computational and Applied Mathematics, 8(2):121–126, June 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X82900651.

deGroen:1989:SGM

[dGvV89] P. P. N. de Groen and M. van Veldhuizen. A stabilized Galerkin method for convection dominated diffusion problems. Journal of Computational and Applied Mathematics, 28(??):155–162, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789903270.

Delosme:1989:BAS

[DI89] Jean-Marc Delosme and Ilse C. F. Ipsen. From Bareiss’ algorithm to the stable computation of partial correlations. Journal of Computational and Applied Mathematics, 27(1–2):53–91, September 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789903610.

Dick:1982:LMA

[Dic82a] E. Dick. Local modal analysis of unsteady discrete systems with nodal equations of unequal form. Journal of Computational and Applied Mathematics, 8(1):63–67, March 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X82900092.

Dick:1982:OMH

[Dic82b] E. Dick. Overrelaxation methods for hybrid systems of first order partial differential equations. Journal of Computational and Applied Mathematics, 8(4):267–273, December 1982. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X8290050X.

Dick:1985:MMC

[Dic85]  E. Dick.  A multigrid

method for the Cauchy– Riemann equations based on flux-difference splitting and its extension to the steady Euler equations. Journal of Computational and Applied Mathematics, 12–13(??):247–263, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042785900226.

Dick:1987:EST

[Dic87] E. Dick. Elliptic solution techniques for Euler and Navier– Stokes equations in steady flow. Journal of Computational and Applied Mathematics, 20(??):199–209, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901373.

Dick:1989:MAF

[Dic89] Erik Dick. Multigrid acceleration of a flux-difference splitting method for steady Euler equations. Journal of Computational and Applied Mathematics, 28(??): 173–180, December 1989.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042789903294.

Dierckx:1986:AFD

[Die86] P. Dierckx. An algorithm for fitting data over a circle using tensor product splines. Journal of Computational and Applied Mathematics, 15(2):161–173, June 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786900245.

Dixon:1986:GRQ

[Dix86] Jennifer A. Dixon. Generalized reducible quadrature methods for Volterra integral and integro-differential equations. Journal of Computational and Applied Mathematics, 16(1):27–42, September 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786901718.

Dixon:1987:NWS

[Dix87] Jennifer A. Dixon. A nonlinear weakly singular Volterra integro-differential equation arising from a reaction– diffusion study in a small cell. Journal of Computational and Applied Mathematics, 18(3):289–305, June 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900033.

Dyer:1984:BBA

[DKW84] M. E. Dyer, N. Kayal, and J. Walker. A branch and bound algorithm for solving the multiple-choice knapsack problem. Journal of Computational and Applied Mathematics, 11(2):231–249, October 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900232.

Deville:1982:AEM

[DL82] M. Deville and G. Labrosse.

An algorithm for the evaluation of multidimensional (direct and inverse) discrete Chebyshev transforms. Journal of Computational and Applied Mathematics, 8(4): 293–304, December 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0771050X82900559.

deLuca:1989:QVI

[dLM89] Marino de Luca and Antonino Maugeri. Quasi-variational inequalities and applications to the traffic equilibrium problem: discussion of a paradox. Journal of Computational and Applied Mathematics, 28(??):163–171, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789903282.

Day:1982:SFO

[DM82] J. D. Day and D. N. P. Murthy. An l( a )- stable fourth order Rosenbrock method with error estimator. Journal of Computational and Applied Mathematics, 8(1):21–27, March 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X82900031.

Day:1983:CSO

[DM83] J. D. Day and D. N. P. Murthy. A comparison of some ODE solvers which require Jacobian evaluations. Journal of Computational and Applied Mathematics, 9(3): 213–219, September 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042783900146.

Dahmen:1984:OAR

[DM84a] Wolfgang Dahmen and Charles A. Micchelli. On the optimal approximation rates for criss-cross finite element spaces. Journal of Computational and Applied Mathematics, 10(3):255–273, June 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900384.

Dreyer:1984:NSP

[DM84b] T. P. Dreyer and D. M. Murray. The numerical solution of the pre-elimination models of cable configurations. Journal of Computational and Applied Mathematics, 10(1):81–91, February 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900724.

Deville:1987:ISI

[DMP87] M. O. Deville, E. H. Mund, and A. T. Patera. Iterative solution of isoparametric spectral element equations by loworder finite element preconditioning. Journal of Computational and Applied Mathematics, 20(??):189–197, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901361.

Dargahi-Noubary:1985:EEE

[DN85] G. R. Dargahi-Noubary. On evaluation of the effects of earthquake and wind on reliability of a structure. Journal of Computational and Applied Mathematics, 12–13(??):233–237, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042785900202.

Dargahi-Noubary:1987:CGR

[DN87] G. R. Dargahi-Noubary. Classification of Gaussian random processes using a simple linear discriminant and its application to seismic differentiation. Journal of Computational and Applied Mathematics, 19(2):269–277, August 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901956.

Dargahi-Noubary:1988:PSS

[DN88] G. R. Dargahi-Noubary. A parametric solution for simple stress-strength model of failure with an application. Journal of Computational and Applied Mathematics, 23(2):185–197, August 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788902804.

Dobronec:1988:TSS

[Dob88] B. S. Dobronec. Two-sided solution of ODEs via a posteriori error estimates. Journal of Computational and Applied Mathematics, 23(1):53–61, July 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903305.

Doring:1985:ZFF

[Dör85] B. Döring. On the zeros of Flett’s function. Journal of Computational and Applied Mathematics, 12– 13(??):265–270, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042785900238.

Dormand:1980:FER

[DP80] J. R. Dormand and P. J.

Prince. A family of embedded Runge–Kutta formulae. Journal of Computational and Applied Mathematics, 6(1):19–26, March 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X80900133.

Dormand:1986:RSE

[DP86] J. R. Dormand and P. J.

Prince. A reconsideration of some embedded Runge–Kutta formulae. Journal of Computational and Applied Mathematics, 15(2):203–211, June 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786900270.

Dierckx:1988:VNN

[DP88] P. Dierckx and R. Piessens. Vol. 4, nr. 3, November 1988. Journal of Computational and Applied Mathematics, 24(3):N1, December 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903032.

DAlmeida:1987:IRE

[DR87a] Filomena Dias D’Almeida and Maria João Rodrigues. Is the iterative refinement of eigenelements an expensive technique? Journal of Computational and Applied Mathematics, 20(??): 159–166, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042787901336.

DeVilliers:1987:OLS

[DR87b] J. M. De Villiers and C. H. Rohwer. Optimal local spline interpolants. Journal of Computational and Applied Mathematics, 18(1):107–119, April 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900598.

Deakin:1987:GMR

[DR87c] A. S. Deakin and H. Rasmussen. A generalized moment representation of two free boundary problems. Journal of Computational and Applied Mathematics, 19(3):351–361, September 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787902044.

Draux:1987:EAN

[Dra87]  André Draux.  The ep-

silon algorithm in a noncommutative algebra. Journal of Computational and Applied Mathematics, 19(1 (supplement 1)):9–21, July 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901671.

Draux:1988:NNP

[Dra88]  André Draux.  On the

non-normal Padé table in a non-commutative algebra. Journal of Computational and Applied Mathematics, 21(3):271–288, March 1988. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788903147.

Drummond:1984:CSC

[Dru84]  J. E. Drummond.  Con-

vergence speeding, convergence and summability. Journal of Computational and Applied Mathematics, 11 (2):145–159, October 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042784900177.

DeDier:1989:IBF

[DRV89] Bart De Dier, Dirk Roose, and Paul Van Rompay. Interaction between fold and Hopf curves leads to new bifurcation phenomena. Journal of Computational and Applied Mathematics, 26(1–2):171–186, June 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789901544.

Dongarra:1989:BRM

[DSH89] Jack J. Dongarra, Danny C. Sorensen, and Sven J. Hammarling. Block reduction of matrices to condensed forms for eigenvalue computations. Journal of Computational and Applied Mathematics, 27(1–2):215–227, September 1989. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903671.

Dierckx:1988:ASR

[DSV88] Paul Dierckx, Paul Suetens, and Dirk Vandermeulen. An algorithm for surface reconstruction from planar contours using smoothing splines. Journal of Computational and Applied Mathematics, 23(3):367–388, September 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788900076.

Duff:1989:MSM

[Duf89]  Iain S. Duff.  Multipro-

cessing a sparse matrix code on the Alliant FX/8. Journal of Computational and Applied Mathematics, 27(1–2):229–239, September 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903683.

Dunham:1980:FPD

[Dun80] Charles B. Dunham. A Fortran program for discrete nonlinear Chebyshev approximation. Journal of Computational and Applied Mathematics, 6(3):241–245, September 1980. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X80900315.

Dunham:1982:WDN

[Dun82] Charles B. Dunham. Weighted discrete nonlinear Chebyshev approximation. Journal of Computational and Applied Mathematics, 8(2):133, June 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X82900687.

Dunham:1984:LIM

[Dun84a] Charles B. Dunham. The linear inequality method for rational approximation with interpolation. Journal of Computational and Applied Mathematics, 10(3):383, June 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042784900451.

Dunham:1984:SLI

[Dun84b] Charles B. Dunham. Stability of the linear inequality method for rational Chebyshev approximation. Journal of Computational and Applied Mathematics, 11(2):139–143, October 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
0377042784900165.

Dunham:1988:AOF

[Dun88] Charles B. Dunham. Approximation with one (or few) parameters nonlinear. Journal of Computational and Applied Mathematics, 21 (1):115–118, January 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042788903937.

Dellnitz:1989:CMB

[DW89] Michael Dellnitz and Bodo Werner. Computational methods for bifurcation problems with symmetries-with special attention to steady state and Hopf bifurcation points. Journal of Computational and Applied Mathematics, 26(1–2):97–123, June 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789901507.

Dyksen:1986:TPG

[Dyk86] Wayne R. Dyksen. A tensor product generalized ADI method for elliptic problems on cylindrical domains with holes. Journal of Computational and Applied Mathematics, 16(1):43–58, September 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704278690172X.

Essah:1989:NSF

[ED89] W. A. Essah and L. M. Delves. The numerical solution of first kind integral equations. Journal of Computational and Applied Mathematics, 27(3):363–387, November 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704278990023X.

Eichenauer-Herrmann:1989:MEL

[EHL89] Jürgen Eichenauer-Herrmann and Jürgen Lehn. Minimax estimators for the location parameter of a noncentral exponential distribution when the parameter space is bounded. Journal of Computational and Applied Mathematics, 26(3):333–337, July 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789903051.

Ehrenmark:1988:TPF

[Ehr88] Ulf Torsten Ehrenmark. A three-point formula for numerical quadrature of oscillatory integrals with variable frequency. Journal of Computational and Applied Mathematics, 21(1):87–99, January 1988. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903901.

Eiermann:1984:CPT

[Eie84] Michael Eiermann. On the convergence of Padé-type approximants to analytic functions. Journal of Computational and Applied Mathematics, 10(2):219–227, April 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900591.

Egecioglu:1989:RDA

[EKL89] Ömer Eğecioğlu, Çetin K.

Koc, and Alan J. Laub. A recursive doubling algorithm for solution of tridiagonal systems on hypercube multiprocessors. Journal of Computational and Applied Mathematics, 27(1–2):95–108, September 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903622.

Evans:1988:CSM

[EL88] D. J. Evans and C. Li.

On the convergence of the SAOR method and the error bounds for its acceleration. Journal of Computational and Applied Mathematics, 23(3):267–279, September 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788900027.

Eggert:1989:TRA

[EL89a] Norm Eggert and John Lund. The trapezoidal rule for analytic functions of rapid decrease. Journal of Computational and Applied Mathematics, 27(3):389–406, November 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789900241.

Ewerbring:1989:CCG

[EL89b] L. Magnus Ewerbring and Franklin T. Luk. Canonical correlations and generalized SVD: Applications and new algorithms. Journal of Computational and Applied Mathematics, 27(1– 2):37–52, September 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042789903609.

Eilbeck:1988:BDC

[ELS88] J. C. Eilbeck, P. S. Lomdahl, and A. C. Scott. Biomolecular dynamics by computer analysis. Journal of Computational and Applied Mathematics, 22(2–3):297–299, June 1988. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788904086.

Eilbeck:1986:CBF

[EM86] J. C. Eilbeck and V. S.

Manoranjan. A comparison of basis functions for the pseudo-spectral method for a model reaction–diffusion problem. Journal of Computational and Applied Mathematics, 15(3):371–378, July 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278690227X.

Engels:1987:SPD

[EM87] Hermann Engels and Bernd Anton Merschen. Some properties of dual quadrature formulae. Journal of Computational and Applied Mathematics, 17(1–2):69–77, February 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900392.

Abd-el-Malek:1989:ASF

[eMH89] Mina B. Abd el Malek and Sarwat N. Hanna. Approximate solution of a flow over a ramp for large Froude number. Journal of Computational and Applied Mathematics, 28(??):105–117, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789903233.

Evans:1982:NSE

[EO82] D. J. Evans and S. O. Okolie. The numerical solution of an elliptic P.D.E. with periodic boundary conditions in a rectangular region by the spectral resolution method. Journal of Computational and Applied Mathematics, 8(4):237–241, December 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X82900468.

Eriksson:1988:STS

[Eri88] Jan R. Eriksson. The simultaneous two-step entropy model. Journal of Computational and Applied Mathematics, 22(2–3):271–283, June 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788904062.

El-Sirafy:1985:TDF

[ES85] Ibrahim H. El-Sirafy. Twodimensional flow of a nonstationary micropolar fluid in the half-plane for which the shear stresses are given on the boundary. Journal of Computational and Applied Mathematics, 12–13(??):271–276, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278590024X.

Elcrat:1986:CFS

[ET86] Alan R. Elcrat and Lloyd N.

Trefethen. Classical freestreamline flow over a polygonal obstacle. Journal of Computational and Applied Mathematics, 14(1–2):251– 265, February 1986. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042786901421.

Fabrikant:1987:EPS

[Fab87] V. I. Fabrikant. Electrostatic problem of several arbitrarily charged unequal coaxial disks. Journal of Computational and Applied Mathematics, 18(2):129–147, May 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900124.

Farwig:1986:MIA

[Far86]  Reinhard Farwig.  Multi-

variate interpolation of arbitrarily spaced data by moving least squares methods. Journal of Computational and Applied Mathematics, 16(1):79–93, September 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786901755.

Filippi:1989:SCD

[FB89] S. Filippi and U. Buchacker. Stepsize control for delay differential equations using a pair of formulae. Journal of Computational and Applied Mathematics, 26(3):339–343, July 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789903063.

Filippi:1986:NRK

[FG86] S. Filippi and J. Gräf. New Runge–Kutta–Nyström formula-pairs of order 8(7), 9(8), 10(9) and 11(10) for differential equations of the form y′′ = f(x, y). Journal of Computational and Applied Mathematics, 14(3):361–370, March 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786900737.

Ford:1989:UFV

[FG89] J. A. Ford and R.-A. Ghandhari. On the use of function-values in unconstrained optimisation. Journal of Computational and Applied Mathematics, 28(??): 187–198, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042789903312.

Fisher:1988:ITS

[FH88] P. G. Fisher and A. J. Hughes Hallett. Iterative techniques for solving simultaneous equation systems: a view from the economics literature. Journal of Computational and Applied Mathematics, 24(1–2):241–255, November 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903561.

Fischer:1989:ADC

[Fis89] Herbert Fischer. Automatic differentiation of characterizing sequences. Journal of Computational and Applied Mathematics, 28(??): 181–185, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042789903300.

Feng:1988:AIN

[FK88] Yu Yu Feng and Jernej Kozak.

An approach to the interpolation of nonuniformly spaced data. Journal of Computational and Applied Mathematics, 23(2):169–178, August 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788902786.

Farmer:1985:ACZ

[FL85] M. R. Farmer and G. Loizou. An algorithm for the computation of zeros of a special class of entire functions. Journal of Computational and Applied Mathematics, 12–13(??):433–445, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042785900378.

Forbes:1986:AMM

[For86a] G. W. Forbes. Automation of the manipulation of multivariate power series. Journal of Computational and Applied Mathematics, 15(1):37–58, May 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786902372.

Forbes:1986:TMM

[For86b] G. W. Forbes. Truncation and manipulation of multivariate power series. Journal of Computational and Applied Mathematics, 15(1):27–36, May 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786902360.

Forsyth:1980:HOP

[FR80] P. Forsyth, Jr. and H. Rasmussen. High order perturbation solution of the electrochemical smoothing problem. Journal of Computational and Applied Mathematics, 6(1):37–42, March 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X80900157.

Freeden:1984:SSI

[Fre84]  Willi Freeden.  Spherical

spline interpolation-basic theory and computational aspects. Journal of Computational and Applied Mathematics, 11(3):367–375, December 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900116.

Ford:1987:CMM

[FS87] John A. Ford and Adel F.

Saadallah. On the construction of minimisation methods of quasi-Newton type. Journal of Computational and Applied Mathematics, 20(??):239–246, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787901403.

Fukushima:1984:CCO

[Fuk84]  Masao Fukushima.  On

the convergence of a class of outer approximation algorithms for convex programs. Journal of Computational and Applied Mathematics, 10(2):147–156, April 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042784900517.

Fack:1987:SFD

[FVD87] V. Fack, G. Vanden Berghe, and H. E. De Meyer. Some finite difference methods for computing eigenvalues and eigenvectors of special twopoint boundary value problems. Journal of Computational and Applied Mathematics, 20(??):211–217, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787901385.

Fishman:1987:PSM

[FW87] Louis Fishman and Stephen C. Wales. Phase space methods and path integration: the analysis and computation of scalar wave equations. Journal of Computational and Applied Mathematics, 20(??):219–238, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901397.

Gabutti:1980:GMC

[Gab80] Bruno Gabutti. On the generalization of a method for computing Bessel function integrals. Journal of Computational and Applied Mathematics, 6(2):167–168, June 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X80900108.

Garloff:1986:PBI

[Gar86a]  J. Garloff.  On power-

boundedness of interval matrices. Journal of Computational and Applied Mathematics, 14(3):353–360, March 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786900725.

Garloff:1986:CTP

[Gar86b] Jürgen Garloff. Continuation of two previous bibliographies on interval mathematics. Journal of Computational and Applied Mathematics, 15(2):265, June 1986. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786900348.

Gautschi:1985:OPC

[Gau85] Walter Gautschi. Orthogonal polynomials — constructive theory and applications. Journal of Computational and Applied Mathematics, 12–13(??):61–76, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704278590007X.

Gerck:1980:CFC

[Gd80] Edgardo Gerck and Augusto Brandão d’Oliveira. Continued fraction calculation of the eigenvalues of tridiagonal matrices arising from the Schroedinger equation. Journal of Computational and Applied Mathematics, 6(1):81–82, March 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X80900200.

Gera:1986:SIE

[Ger86] Amos E. Gera. Singular integral equations with a Cauchy kernel. Journal of Computational and Applied Mathematics, 14(3):311–318, March 1986. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786900695.

Garibotti:1983:RRR

[GG83] C. R. Garibotti and F. F.

Grinstein. Recent results relevant to the evaluation of infinite series. Journal of Computational and Applied Mathematics, 9(3):193–200, September 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042783900110.

Grosjean:1985:SEC

[GG85] C. C. Grosjean and M. J.

Goovaerts. On the series expansion of certain types of integral transforms — Part I. Journal of Computational and Applied Mathematics, 12–13(??):277–298, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042785900251.

Grosjean:1988:AEO

[GG88] C. C. Grosjean and M. J.

Goovaerts. The analytical evaluation of one-dimensional Gaussian path-integrals. Journal of Computational and Applied Mathematics, 21(3): 311–331, March 1988. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042788903160.

Gerber:1981:WPD

[GGD81] Hans U. Gerber, Marc Goovaerts, and Nelson DePril. The Wiener process with drift between a linear retaining and an absorbing barrier. Journal of Computational and Applied Mathematics, 7(4):267–269, December 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X81900541.

Goovaerts:1983:E

[GGW83] M. J. Goovaerts, W. B. Gragg, and L. Wuytack. Editorial. Journal of Computational and Applied Mathematics, 9(1):1, March 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042783900237.

George:1989:SSP

[GHLN89] Alan George, Michael Heath, Joseph Liu, and Esmond Ng. Solution of sparse positive definite systems on a hypercube. Journal of Computational and Applied Mathematics, 27(1–2):129–156, September 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903646.

Gill:1988:CAF

[Gil88]  John Gill.  Compositions

of analytic functions of the form F_{n}(z) = F_{n−1}(f_{n}(z)), f_{n}(z) → f(z). Journal of Computational and Applied Mathematics, 23(2):179– 184, August 1988. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042788902798.

Genz:1986:NEC

[GK86] Alan Genz and David K.

Kahaner. The numerical evaluation of certain multivariate normal integrals. Journal of Computational and Applied Mathematics, 16(2):255–258, October 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786901007.

Gearhart:1989:ASM

[GK89] William B. Gearhart and Mathew Koshy. Acceleration schemes for the method of alternating projections. Journal of Computational and Applied Mathematics, 26(3):235–249, July 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789902963.

Giordano:1983:EAZ

[GL83] Carla Giordano and Andrea Laforgia. Elementary approximations for zeros of Bessel functions. Journal of Computational and Applied Mathematics, 9(3):221–228, September 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042783900158.

Givens:1984:SES

[GL84] Clark R. Givens and Chi Yeung Lo. Series expansions of solutions of u_{xx} + (2νx)u_{x} + ϵ^{2}u_{tt} = u_{t}. Journal of Computational and Applied Mathematics, 10(1):7–23, February 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900645.

Gabutti:1987:NAD

[GL87a] Bruno Gabutti and Paolo Lepora. A novel approach for the determination of asymptotic expansions of certain oscillatory integrals. Journal of Computational and Applied Mathematics, 19(2):189–206, August 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901890.

Gasca:1987:ETI

[GL87b] M. Gasca and E. Lebrón.

Elimination techniques and interpolation. Journal of Computational and Applied Mathematics, 19(1 (supplement 1)):125–132, July 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901804.

Glasser:1984:SIA

[Gla84] M. L. Glasser. On some integrals arising in mathematical physics. Journal of Computational and Applied Mathematics, 10(3):293–299, June 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900414.

Gear:1985:AIE

[GLG85] C. W. Gear, B. Leimkuhler, and G. K. Gupta. Automatic integration of Euler– Lagrange equations with constraints. Journal of Computational and Applied Mathematics, 12–13(??):77–90, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
0377042785900081.

Gregus:1987:DCF

[GLSZ87] M. Gregus, Jr., Yu. Yu. Lobanov, O. V. Sidorova, and E. P. Zhidkov. On the deterministic computation of functional integrals in application to quantum mechanical problems. Journal of Computational and Applied Mathematics, 20(??):247–256, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901415.

Genz:1980:RAA

[GM80a] A. C. Genz and A. A. Malik. Remarks on Algorithm 006: an adaptive algorithm for numerical integration over an N-dimensional rectangular region. Journal of Computational and Applied Mathematics, 6(4):295–302, December 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X8090039X.

Graves-Morris:1980:GQD

[GM80b] P. R. Graves-Morris. A generalised Q.D. algorithm. Journal of Computational and Applied Mathematics, 6(3): 247–249, September 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0771050X80900327.

Graves-Morris:1981:CRS

[GM81] P. R. Graves-Morris. The convergence of ray sequences of Padé approximants of Stieltjes functions. Journal of Computational and Applied Mathematics, 7(3):191–201, September 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X81900188.

Graves-Morris:1984:SFR

[GM84] P. R. Graves-Morris. Symmetrical formulas for rational interpolants. Journal of Computational and Applied Mathematics, 10(1):107–111, February 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900748.

Gasca:1989:CRI

[GMM89] M. Gasca, J. J. Martínez, and G. Mühlbach. Computation of rational interpolants with prescribed poles. Journal of Computational and Applied Mathematics, 26(3):297–309, July 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903026.

Graves-Morris:1988:RCT

[GMS88] P. R. Graves-Morris and E. B. Saff. Row convergence theorems for generalised inverse vector-valued Padé approximants. Journal of Computational and Applied Mathematics, 23(1):63–85, July 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788903317.

Gill:1988:RDC

[GMSW88] Philip E. Gill, Walter Murray, Michael A. Saunders, and Margaret H. Wright. Recent developments in constrained optimization. Journal of Computational and Applied Mathematics, 22(2–3):257–270, June 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788904050.

Gautschi:1989:GKQ

[GN89] Walter Gautschi and Sotorios E. Notaris. Gauss– Kronrod quadrature formulae for weight functions of Bernstein–Szegő type. Journal of Computational and Applied Mathematics, 25 (2):199–224, February 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/

science/article/pii/0377042789900472.

See erratum [Ano89k].

Golbabai:1986:FAA

[Gol86] A. Golbabai. Finite amplitude axisymmetric convection between rigid rotating planes. Journal of Computational and Applied Mathematics, 16(3):355–369, November 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786900063.

Goossenaerts:1988:BRF

[Goo88a] Jan Goossenaerts. Book review: Foundations and applications of Montague grammar; part 1: philosophy, framework, computer science: T. M. V. Janssen, CWI Tract 19, Centre for Mathematics and Computer Science, Amsterdam, The Netherlands, 1986, 205 pp., Price: Dfl. 31.30, ISBN: 90-6196-292-

7. Journal of Computational and Applied Mathematics, 24(3):N3–N4, December 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903081.

Goossenaerts:1988:BRE

[Goo88b] Jan Goossenaerts. Book review: The ecology of computation: B. A. Huberman (ed.) Studies in Computer Science and Artificial Intelligence 2, North-Holland, Amsterdam, New York, Oxford, Tokyo, 1988, 342 pp., Price: US $84.25 / Dfl. 160.00, ISBN: 0-444-70375-6. Journal of Computational and Applied Mathematics, 24(3):N4–N5, December 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903093.

Graney:1981:NSN

[GR81] L. Graney and A. A. Richardson. The numerical solution of non-linear partial differential equations by the method of lines. Journal of Computational and Applied Mathematics, 7(4):229–236, December 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X81900504.

Greydanus:1983:NSR

[GR83] J. Greydanus and H. Rasmussen. Numerical solution of a Rayleigh–Taylor instability problem. Journal of Computational and Applied Mathematics, 9(2):97–103, June 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278390033X.

Greydanus:1989:SGS

[GR89a] J. Greydanus and H. Rasmussen. On the solution of a generalized Stefan problem. Journal of Computational and Applied Mathematics, 25(1):47–60, January 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789900745.

Griewank:1989:CCS

[GR89b] Andreas Griewank and George W.

Reddien. Computation of cusp singularities for operator equations and their discretizations. Journal of Computational and Applied Mathematics, 26(1–2):133–153, June 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789901520.

Gragg:1986:AUH

[Gra86] William B. Gragg. The QR algorithm for unitary Hessenberg matrices. Journal of Computational and Applied Mathematics, 16(1):1– 8, September 1986. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/037704278690169X.

Greenspan:1981:NCAa

[Gre81a] Donald Greenspan. A new computer approach to the modeling of atoms, ions and molecules: Part I — hydrogen. Journal of Computational and Applied Mathematics, 7(1):41–49, March 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X81900073.

Greenspan:1981:NCAb

[Gre81b] Donald Greenspan. A new computer approach to the modelling of atoms and molecules. Part II. The oxygen molecule O_{2}. Journal of Computational and Applied Mathematics, 7(2):129–133, June 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X81900462.

Greenspan:1986:EAM

[Gre86a] Donald Greenspan. An elementary algebraic method for approximating average radii of first and second ring electrons. Journal of Computational and Applied Mathematics, 16(1):117–134, September 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
0377042786901780.

Greenspan:1986:MCE

[Gre86b] Donald Greenspan. A mathematical curiosity in estimating the radius of the first ring electrons of an arbitrary atom. Journal of Computational and Applied Mathematics, 15(1):103–108, May 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786902426.

Greenspan:1987:ANS

[Gre87]  Donald Greenspan.  Ana-

lytical and numerical studies in deterministic modeling of the states of ions and atoms. Journal of Computational and Applied Mathematics, 19(2):207–222, August 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901907.

Gribble:1982:IPQ

[Gri82] J. D. Gribble. Inner product quadrature formulas exact on maximal product spaces of functions. Journal of Computational and Applied Mathematics, 8(2):73–79, June 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        0771050X82900584.

Grosjean:1984:FCC

[Gro84a]  C. C. Grosjean.  Formu-

lae concerning the computation of the Clausen integral Cl_{2}(Θ). Journal of Computational and Applied Mathematics, 11(3):331–342, December 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042784900086.

Grosjean:1984:OPL

[Gro84b] C. C. Grosjean. The orthogonality property of the Lommel polynomials and a twofold infinity of relations between Rayleigh’s σ-sums. Journal of Computational and Applied Mathematics, 10(3):355–382, June 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704278490044X.

Grosjean:1985:TRG

[Gro85] C. C. Grosjean. Theory of recursive generation of systems of orthogonal polynomials: an illustrative example. Journal of Computational and Applied Mathematics, 12–13(??):299–318, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
0377042785900263.

Grosjean:1986:WFG

[Gro86]  C. C. Grosjean.  The

weight functions, generating functions and miscellaneous properties of the sequences of orthogonal polynomials of the second kind associated with the Jacobi and the Gegenbauer polynomials. Journal of Computational and Applied Mathematics, 16(3):259–307, November 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786900014.

Grosjean:1988:GFC

[Gro88] C. C. Grosjean. A general formula for the calculation of Gaussian path-integrals in two and three Euclidean dimensions. Journal of Computational and Applied Mathematics, 23(2):199–234, August 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788902816.

Garloff:1980:BIM

[GS80] Jürgen Garloff and Klaus-Peter Schwierz. A bibliography on interval mathematics. Journal of Computational and Applied Mathematics, 6(1):67–79, March 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X80900194.

Gutknecht:1986:MTT

[GS86] Martin H. Gutknecht and Edward B. Saff. A de montessustype theorem for CF approximation. Journal of Computational and Applied Mathematics, 16(2):251–254, October 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786900993.

Gladwell:1987:ASI

[GSB87] I. Gladwell, L. F. Shampine, and R. W. Brankin. Automatic selection of the initial step size for an ODE solver. Journal of Computational and Applied Mathematics, 18(2):175–192, May 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704278790015X.

Gutknecht:1986:ECF

[Gut86a] Martin H. Gutknecht. The evaluation of the conjugate function of a periodic spline on a uniform mesh. Journal of Computational and Applied Mathematics, 16(2):181–201, October 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786900919.

Gutknecht:1986:NCM

[Gut86b] Martin H. Gutknecht. Numerical conformal mapping methods based on function conjugation. Journal of Computational and Applied Mathematics, 14(1–2):31–77, February 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786901305.

Gonzalez-Vera:1987:CAT

[GV87] Pablo González-Vera. On certain applications of twopoint Padé type approximants with a single pole. Journal of Computational and Applied Mathematics, 19(1 (supplement 1)):151–160, July 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278790183X.

Galperin:1981:SIP

[GW81] A. Galperin and Z. Waksman. A separable integer programming problem equivalent to its continual version. Journal of Computational and Applied Mathematics, 7(3):173–179, September 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X81900152.

Goovaerts:1987:P

[GW87] M. Goovaerts and L. Wuytack. Preface. Journal of Computational and Applied Mathematics, 20(??):1, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704278790121X.

Gwinner:1981:BND

[Gwi81] Joachim Gwinner. Bibliography on non-differentiable optimization and non-smooth analysis. Journal of Computational and Applied Mathematics, 7(4):277–285, December 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X81900565.

Hadjidimos:1985:OIM

[Had85] A. Hadjidimos. Optimum iterative methods for the solution of singular linear systems arising from the discretization of elliptic P.D.E.’s. Journal of Computational and Applied Mathematics, 12–13(??):319–329, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042785900275.

Hadjidimos:1987:SIM

[Had87]  A. Hadjidimos.  A sur-

vey of the iterative methods for the solution of linear systems by extrapolation, relaxation and other techniques. Journal of Computational and Applied Mathematics, 20(??):37–51, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901245.

Hadjidimos:1988:P

[Had88] Apostolos Hadjidimos. Preface. Journal of Computational and Applied Mathematics, 24(1–2):1–2, November 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903391.

Hake:1987:NSP

[Hak87] J.-Fr. Hake. The numerical solution of PDE problems with standard software packages. Journal of Computational and Applied Mathematics, 20(??):257–265, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        0377042787901427.

Haken:1988:SC

[Hak88] H. Haken. Synergetics and computers. Journal of Computational and Applied Mathematics, 22(2–3):197–202, June 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788903998.

Hallett:1982:ATS

[Hal82] A. J. Hughes Hallett. Alternative techniques for solving systems of nonlinear equations. Journal of Computational and Applied Mathematics, 8(1):35–48, March 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X82900055.

Hallett:1984:SOI

[Hal84] A. J. Hughes Hallett. Secondorder iterations with guaranteed convergence. Journal of Computational and Applied Mathematics, 10(3):285–291, June 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042784900402.

Hamacher:1982:OCR

[Ham82]  Horst Hamacher.  Op-

timal cocircuits in regular matroids and applications. Journal of Computational and Applied Mathematics, 8(3):171–177, September 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X82900390. See erra-

tum [Ano82].

Hansen:1987:DNS

[Han87] Per Christian Hansen. Detection of near-singularity in Cholesky and LDL^{T} factorizations. Journal of Computational and Applied Mathematics, 19(3):293–299, September 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901993.

Hansen:1988:NRM

[Han88] Per Christian Hansen. The 2-norm of random matrices. Journal of Computational and Applied Mathematics, 23(1):117–120, July 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903366.

Hansen:1985:SAL

[HC85] Per Christian Hansen and Søren Christiansen. An SVD analysis of linear algebraic equations derived from first kind integral equations. Journal of Computational and Applied Mathematics, 12–13(??):341–357, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042785900299.

Hebsaker:1982:CCR

[Heb82] H.-M. Hebsaker. Conditions for the coefficients of Runge– Kutta methods for systems of n-th order differential equations. Journal of Computational and Applied Mathematics, 8(1):3–14, March 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X82900018.

Hellemans:1982:NSC

[Hel82] P. Hellemans. A numerical solution of the Cauchy problem based on trigonometric interpolation. Journal of Computational and Applied Mathematics, 8(4):305–306, December 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
0771050X82900560.

Hemker:1982:NDC

[Hem82] P. W. Hemker. A note on defect correction processes with an approximate inverse of deficient rank. Journal of Computational and Applied Mathematics, 8(2):137–139, June 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X82900705.

Henrici:1982:PEE

[Hen82] Peter Henrici. Pointwise error estimates for trigonometric interpolation and conjugation. Journal of Computational and Applied Mathematics, 8(2):131–132, June 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X82900675.

Hendriksen:1987:SHM

[Hen87] E. Hendriksen. The strong Hamburger moment problem and self-adjoint operators in Hilbert space. Journal of Computational and Applied Mathematics, 19(1 (supplement 1)):79–88, July 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        0377042787901750.

Hermann:1983:NTP

[Her83] Martin Hermann. The numerical treatment of perturbed bifurcation problems in ordinary differential equations. Journal of Computational and Applied Mathematics, 9(1):71–80, March 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042783900298.

Honig:1984:MNI

[HH84] G. Honig and U. Hirdes. A method for the numerical inversion of Laplace transforms. Journal of Computational and Applied Mathematics, 10(1):113–132, February 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704278490075X.

Hake:1989:LAS

[HH89] J.-Fr. Hake and W. Homberg. Linear algebra software on IBM and CRAY computers. Journal of Computational and Applied Mathematics, 26(3):311–325, July 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
0377042789903038.

Huang:1986:GNM

[HKD86] Chung-Yuan Huang, Stephen R.

Kennon, and George S. Dulikravich. Generalized nonlinear minimal residual (GNLMR) method for iterative algorithms. Journal of Computational and Applied Mathematics, 16 (2):215–232, October 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042786900932.

Hairer:1988:FNS

[HLS88] E. Hairer, Ch. Lubich, and M. Schlichte. Fast numerical solution of weakly singular Volterra integral equations. Journal of Computational and Applied Mathematics, 23(1):87–98, July 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903329.

Hatzopoulos:1985:ASL

[HM85] M. Hatzopoulos and N. M.

Missirlis. Advantages for solving linear systems in an asynchronous environment. Journal of Computational and Applied Mathematics, 12–13(??):331–340, May 1985. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042785900287.

Hoppe:1989:MGC

[HM89] Ronald H. W. Hoppe and Hans D. Mittelmann. A multigrid continuation strategy for parameter-dependent variational inequalities. Journal of Computational and Applied Mathematics, 26(1–2):35–46, June 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789901465.

Hartmann:1986:UAN

[HO86] M. Hartmann and G. Opfer. Uniform approximation as a numerical tool for constructing conformal maps. Journal of Computational and Applied Mathematics, 14(1–2):193–206, February 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278690138X.

Harrar:1988:OSS

[HO88] D. L. Harrar II and J. M. Ortega. Optimum mstep SSOR preconditioning. Journal of Computational and Applied Mathematics, 24(1–2):195–198, November 1988. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903524.

Hoffmann:1987:SLS

[Hof87] W. Hoffmann. Solving linear systems on a vector computer. Journal of Computational and Applied Mathematics, 18(3):353–367, June 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900082.

Hoffmann:1989:ESN

[Hof89]  W. Hoffmann.  An esti-

mate for the spectral norm of the inverse of a matrix with the Gauss–Jordan algorithm. Journal of Computational and Applied Mathematics, 27(3):421–428, November 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789900277.

Hoidn:1986:RMD

[Hoi86] Hans-Peter Hoidn. A reparameterisation method to determine conformal maps. Journal of Computational and Applied Mathematics, 14(1–2):155–161, February 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786901366.

Hopkins:1982:SCP

[Hop82] T. R. Hopkins. On the sensitivity of the coefficients of Padé approximants with respect to their defining power series coefficients. Journal of Computational and Applied Mathematics, 8(2):105–109, June 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X82900638.

Hough:1985:JPS

[Hou85]  D. M. Hough.  Jacobi

polynomial solutions of first kind integral equations for numerical conformal mapping. Journal of Computational and Applied Mathematics, 12–13(??):359–369, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042785900305.

Howes:1984:ATF

[How84] F. A. Howes. An analytical treatment of the formation of one-dimensional steady shock waves in uniform and diverging ducts. Journal of Computational and Applied Mathematics, 10(2):195–201, April 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900578.

Hosny:1982:ACT

[HS82] E. A. Hosny and M. I. Sobhy.

An algorithm for the calculation of transfer functions of two variable systems or networks. Journal of Computational and Applied Mathematics, 8(3):159–164, September 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X82900377.

Herbst:1983:EPM

[HSM83] B. M. Herbst, S. W. Schoombie, and A. R. Mitchell. Equidistributing principles in moving finite element methods. Journal of Computational and Applied Mathematics, 9(4):377–389, December 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042783900092.

Hasegawa:1987:IIO

[HT87] Takemitsu Hasegawa and Tatsuo Torii. Indefinite integration of oscillatory functions by the Chebyshev series expansion. Journal of Computational and Applied Mathematics, 17(1–2):21–29, February 1987. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787900355.

Hubner:1986:NMS

[Hüb86a]  O. Hübner.  The New-

ton method for solving the Theodorsen integral equation. Journal of Computational and Applied Mathematics, 14(1–2):19–30, February 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786901299.

Hubner:1986:FSL

[Hüb86b] O. Hübner. On the fast solution of a linear system arising in numerical conformal mapping. Journal of Computational and Applied Mathematics, 16(3):381–386, November 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786900087.

Hundsdorfer:1987:NMR

[Hun87] W. H. Hundsdorfer. A note on monotonicity of a Rosenbrock method. Journal of Computational and Applied Mathematics, 20(??):267–273, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
0377042787901439.

Hunt:1988:SLM

[Hun88] Roland Hunt. Single-level multigrid. Journal of Computational and Applied Mathematics, 23(2):133–139, August 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788902750.

Igarashi:1985:PSR

[Iga85]  Masao Igarashi.  Practi-

cal stopping rule for finding roots of nonlinear equations. Journal of Computational and Applied Mathematics, 12–13(??):371–380, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042785900317.

Iri:1987:CQF

[IMT87] Masao Iri, Sigeiti Moriguti, and Yoshimitsu Takasawa. On a certain quadrature formula. Journal of Computational and Applied Mathematics, 17(1–2):3–20, February 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900343.

Iserles:1987:BOZ

[IN87] A. Iserles and S. P. Nørsett. Bi-orthogonality and zeros of transformed polynomials. Journal of Computational and Applied Mathematics, 19(1 (supplement 1)):39–45, July 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787901695.

Ioakimidis:1981:OMA

[Ioa81] N. I. Ioakimidis. One more approach to the computation of Cauchy principal value integrals. Journal of Computational and Applied Mathematics, 7(4):289–291, December 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X81900589.

Ioakimidis:1982:QMN

[Ioa82] N. I. Ioakimidis. On the quadrature methods for the numerical solution of singular integral equations. Journal of Computational and Applied Mathematics, 8(2):81–86, June 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X82900596.

Ioakimidis:1984:RAI

[Ioa84]  N. I. Ioakimidis.  A re-

mark on the application of interpolatory quadrature rules to the numerical solution of singular integral equations. Journal of Computational and Applied Mathematics, 11(3):267–276, December 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900013.

Ioakimidis:1986:DPS

[Ioa86a]  N. I. Ioakimidis.  Deter-

mination of poles of sectionally meromorphic functions. Journal of Computational and Applied Mathematics, 15(3):323–327, July 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786902232.

Ioakimidis:1986:NIC

[Ioa86b] N. I. Ioakimidis. A new interpretation of Cauchy type singular integrals with an application to singular integral equations. Journal of Computational and Applied Mathematics, 14(3):271–278, March 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
0377042786900658.

Ioakimidis:1988:SAM

[Ioa88]  N. I. Ioakimidis.  The

successive approximations method for the airfoil equation. Journal of Computational and Applied Mathematics, 21(2):231–238, February 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788902713.

Ioakimidis:1989:LBC

[Ioa89] N. I. Ioakimidis. Location of boundary contours and discontinuity arcs, of known shape and conditions, of analytic functions by using contour integrals. Journal of Computational and Applied Mathematics, 25(3):315–326, May 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789900356.

Ikeuchi:1985:BES

[IS85] Masatoshi Ikeuchi and Michio Sakakihara. Boundary elements in steady convective diffusion problems. Journal of Computational and Applied Mathematics, 12– 13(??):381–389, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042785900329.

Iserles:1987:BOR

[IS87] A. Iserles and E. B. Saff.

Bi-orthogonality in rational approximation. Journal of Computational and Applied Mathematics, 19(1 (supplement 1)):47–54, July 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901701.

Ifantis:1988:BFP

[IS88] E. K. Ifantis and P. D. Siafarikas. Bounds for the first positive zero of a mixed Bessel function. Journal of Computational and Applied Mathematics, 21(2):245–249, February 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788902737.

Ifantis:1987:UBF

[ISK87] E. K. Ifantis, P. D. Siafarikas, and C. B. Kouris. Upper bounds for the first zeros of Bessel functions. Journal of Computational and Applied Mathematics, 17(3):355–358, March 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042787901117.

Iri:1988:ACP

[ITH88] Masao Iri, Takashi Tsuchiya, and Mamoru Hoshi. Automatic computation of partial derivatives and rounding error estimates with applications to large-scale systems of nonlinear equations. Journal of Computational and Applied Mathematics, 24(3):365–392, December 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788902981.

Jablonski:1980:MCA

[Jab80]  Aleksander Jablonski.  A

Monte Carlo algorithm for solving systems of non-linear equations. Journal of Computational and Applied Mathematics, 6(3):171–175, September 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X80900224.

Jacobsen:1987:GCC

[Jac87]  Lisa Jacobsen.  General

correspondence for continued fractions. Journal of Computational and Applied Mathematics, 19(1 (supplement 1)):171–177, July 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
0377042787901853.

Jacobsen:1988:RDC

[Jac88]  Lisa Jacobsen.  Remarks

to a definition of convergence acceleration illustrated by means of continued fractions K(a_{n}1) where a_{n} →

0. Journal of Computational and Applied Mathematics, 21(1):101–110, January 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903913.

Jain:1988:SOS

[Jai88]  R. K. Jain.  Sixth-order

superstable two-step methods for second-order initialvalue problems. Journal of Computational and Applied Mathematics, 24(3): 393–397, December 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042788902993.

Janssen:1987:RCS

[Jan87] H. L. Janssen. Recursive construction of a sequence of solutions to homogeneous linear partial differential equations with constant coefficients. Journal of Computational and Applied Mathematics, 20(??):275–287, November 1987. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901440.

Jeltsch:1985:SAD

[Jel85]  Rolf Jeltsch.  Stability

and accuracy of difference schemes for hyperbolic problems. Journal of Computational and Applied Mathematics, 12–13(??):91–108, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042785900093.

Jeltsch:1986:POB

[Jel86] Rolf Jeltsch. A proof of the order barrier for upwind schemes by Dahlquist’s second barrier. Journal of Computational and Applied Mathematics, 15(2):257–259, June 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786900324.

Jain:1988:SSS

[JG88] R. K. Jain and Rakesh Goel. Superstable single-step methods for second-order initialvalue problems. Journal of Computational and Applied Mathematics, 23(1):121–125, July 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903378.

Jodar:1989:ASC

[JH89] Lucas Jódar and Antonio Hervás. An algorithm for solving coupled differential matrix systems: approximations, convergence and upper error bounds. Journal of Computational and Applied Mathematics, 25(3):351–362, May 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789900381.

Jain:1983:HOO

[JKG83] R. K. Jain, N. S. Kambo, and Rakesh Goel. High-order one-step P -stable methods for the numerical integration of periodic initial value problems. Journal of Computational and Applied Mathematics, 9(4):361–367, December 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042783900079.

Jones:1980:CPT

[JM80] William B. Jones and Arne Magnus. Computation of poles of two-point Padé approximants and their limits. Journal of Computational and Applied Mathematics, 6(2):105–119, June 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X80900042.

Jones:1983:TPP

[JNT83] William B. Jones, Olav Njåstad, and W. J. Thron. Two-point Padé expansions for a family of analytic functions. Journal of Computational and Applied Mathematics, 9(2):105–123, June 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042783900341.

Joe:1985:CMU

[Joe85] S. Joe. Collocation methods using piecewise polynomials for second kind integral equations. Journal of Computational and Applied Mathematics, 12–13(??):391–400, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042785900330.

Joe:1989:GLP

[Joe89] S. Joe. The generation of lattice points for numerical multiple integration. Journal of Computational and Applied Mathematics, 26(3):327–331, July 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278990304X.

Jones:1985:CIG

[JT85] William B. Jones and W. J.

Thron. On the computation of incomplete gamma functions in the complex domain. Journal of Computational and Applied Mathematics, 12–13(??):401–417, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042785900342.

Janssen:1989:OSI

[JV89] H. Janssen and R. Van Dooren. A one-step integration routine for normal differential systems, based on Gauss–Legendre quadrature. Journal of Computational and Applied Mathematics, 28(??):207–217, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903336.

Jacobsen:1986:EOP

[JW86] Lisa Jacobsen and Haakon Waadeland. Even and odd parts of limit periodic continued fractions. Journal of Computational and Applied Mathematics, 15(2):225–233, June 1986. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786900294.

Jacobsen:1989:WDR

[JW89] Lisa Jacobsen and Haakon Waadeland. When does f(z) have a regular C-fraction expansion or a normal Padé table? Journal of Computational and Applied Mathematics, 28(??):199–206, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789903324.

Jea:1988:EAC

[JY88] Kang C. Jea and David M. Young. On the effectiveness of adaptive Chebyshev acceleration for solving systems of linear equations. Journal of Computational and Applied Mathematics, 24(1–2):33–54, November 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788903421.

Kaasschieter:1988:PCG

[Kaa88] E. F. Kaasschieter. Preconditioned conjugate gradients for solving singular systems. Journal of Computational and Applied Mathematics, 24(1–2):265–275, November 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903585.

Kammler:1983:PAI

[Kam83] David W. Kammler. A perturbation analysis of the intrinsic conditioning of an approximate null vector computed with a SVD. Journal of Computational and Applied Mathematics, 9(3):201–204, September 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042783900122.

Karageorghis:1988:NCCa

[Kar88a] Andreas Karageorghis. A note on the Chebyshev coefficients of the general order derivative of an infinitely differentiable function. Journal of Computational and Applied Mathematics, 21(1):129–132, January 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903962.

Karageorghis:1988:NCCb

[Kar88b] Andreas Karageorghis. A note on the Chebyshev coefficients of the moments of the general order derivative of an infinitely differentiable function. Journal of Computational and Applied Mathematics, 21(3):383–386, March 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788903226.

Kulev:1988:SML

[KB88] G. K. Kulev and D. D. Bainov. Second method of Lyapunov and comparison principle for systems with impulse effect. Journal of Computational and Applied Mathematics, 23(3):305–321, September 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788900040.

Keast:1987:CFS

[Kea87] Patrick Keast. Cubature formulas for the surface of the sphere. Journal of Computational and Applied Mathematics, 17(1–2):151–172, February 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787900446.

Kaas:1987:NMD

[KG87a] R. Kaas and M. J. Goovaerts. A new method for deriving bounds for integrals with respect to measures allowed to vary under conical and integral constraints.

Journal of Computational and Applied Mathematics, 20(??):289–297, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901452.

Kok:1987:GMT

[KG87b] Barbara Kok and Tunc Geveci. Galerkin methods for the time dependent incompressible elasticity equations. Journal of Computational and Applied Mathematics, 18(1):53–69, April 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900550.

Khouzam:1986:SSA

[Kho86]  Fouad Khouzam.  Stiffly

stable Adams-type methods. Journal of Computational and Applied Mathematics, 16(3):319–331, November 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786900038.

Kioustelidis:1986:BPR

[Kio86] John B. Kioustelidis. Bounds for positive roots of polynomials. Journal of Computational and Applied Mathematics, 16(2):241–244, October 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786900968.

Kiper:1989:SRR

[Kip89] Ayse Kiper. Some recurrence relations and their parallel evaluation using nested recurrent product form algorithm. Journal of Computational and Applied Mathematics, 28(??):231–235, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704278990335X.

Kiwiel:1986:ASM

[Kiw86] Krzysztof C. Kiwiel. An aggregate subgradient method for nonsmooth and nonconvex minimization. Journal of Computational and Applied Mathematics, 14(3): 391–400, March 1986. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042786900750.

Kuang:1988:SAT

[KJ88] Jiaoxun Kuang and Jun Ji. A survey of AOR and TOR methods. Journal of Computational and Applied Mathematics, 24(1–2):3–12, November 1988. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903408.

Kambo:1983:FOM

[KJG83] N. S. Kambo, R. K. Jain, and Rakesh Goel. A fourth order method for y′′ = f(x, y, y′). Journal of Computational and Applied Mathematics, 9(1):81–90, March 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042783900304. See note

[Ano84f].

Knopfmacher:1989:IPE

[KK89] Arnold Knopfmacher and John Knopfmacher. Inverse polynomial expansions of Laurent series, II. Journal of Computational and Applied Mathematics, 28(??): 249–257, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042789903373.

Klempous:1989:LVA

[KKK89] Ryszard Klempous, Jerzy Kotowski, and Mieczyslaw Kuliszewski. The linearized version of an algorithm for the mixed norms problem. Journal of Computational and Applied Mathematics, 28(??):237–247, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789903361.

Klempous:1988:SMW

[KKN^{+}88] Ryszard Klempous, Jerzy Kotowski, Jan Nikodem, Marek Olesiak, and Jedrzej Ulasiewicz. Some models for water distribution systems. Journal of Computational and Applied Mathematics, 21(3):257–269, March 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788903135.

Knopfmacher:1987:MCL

[KL87] Arnold Knopfmacher and D. S. Lubinsky. Mean convergence of Lagrange interpolation for Freud’s weights with application to product integration rules. Journal of Computational and Applied Mathematics, 17(1– 2):79–103, February 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042787900409.

Klein:1982:EED

[Kle82] A. Klein. An explicit expression for the discrete-time fixed-lag smoothing equation. Journal of Computational and Applied Mathematics, 8(2):127–129, June 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X82900663.

Klein:1987:FSN

[Kle87] Gerd H. Klein. Fitting simple non-tensor-product splines to scattered noisy data on Euclidean d-space. Journal of Computational and Applied Mathematics, 18(3):347–352, June 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900070.

Klip:1987:AAS

[Kli87a] Dorothea A. Klip. An analytic approach to the solution of non-linear equations. Journal of Computational and Applied Mathematics, 20(??):299–306, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901464.

Klip:1987:CZC

[Kli87b] Dorothea A. Klip. Completeness of zero curve tracing for analytic functions. Journal of Computational and Applied Mathematics, 19 (2):243–259, August 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042787901932.

Klir:1988:FIC

[Kli88] George J. Klir. The future of information, computer, and systems sciences. Journal of Computational and Applied Mathematics, 22(2–3):183–195, June 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788903986.

Kraaijevanger:1987:SRS

[KLS87] J. F. B. M. Kraaijevanger, H. W. J. Lenferink, and M. N. Spijker. Stepsize restrictions for stability in the numerical solution of ordinary and partial differential equations. Journal of Computational and Applied Mathematics, 20(??):67–81, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787901269.

Kahaner:1989:URO

[KLT89] D. K. Kahaner, W. F. Lawkins, and S. Thompson. On the use of rootfinding ODE software for the solution of a common problem in nonlinear dynamical systems. Journal of Computational and Applied Mathematics, 28(??):219–230, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903348.

Krizek:1987:GSG

[KN87] M. Krízek and P. Neittaanmäki. On a global superconvergence of the gradient of linear triangular elements. Journal of Computational and Applied Mathematics, 18(2):221–233, May 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900185.

Kohler:1989:ESB

[Köh89] Peter Köhler. Estimates for Sard’s best formulas for linear functionals on C^{S}[a, b]. Journal of Computational and Applied Mathematics, 28(??):259–267, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903385.

Kolbig:1986:Ib

[Köl86a] K. S. Kölbig. On the integral ∫_{0}^{∞} χ^{v−1}(1 + βχ)^{−λ} ln^{m} χ dχ. Journal of Computational and Applied Mathematics, 15 (2):267, June 1986. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/037704278690035X.

Kolbig:1986:Ia

[Köl86b] K. S. Kölbig. On the integral ∫_{0}^{∞} χ^{v−1}(1 + βχ)^{−λ} ln^{m} χ dχ. Journal of Computational and Applied Mathematics, 14(3):319–344, March 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042786900701.

Kolbig:1987:I

[Köl87] K. S. Kölbig. On the integral ∫_{0}^{1} x^{ν−1}(1−x)^{−λ} ln^{m} x dx. Journal of Computational and Applied Mathematics, 18(3):369–394, June 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042787900094.

Kostlan:1988:CTN

[Kos88] Eric Kostlan. Complexity theory of numerical linear algebra. Journal of Computational and Applied Mathematics, 22(2–3):219–230, June 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788904025.

Kiessling:1982:NRJ

[KP82] I. Kiessling and A. Paulik.

A norm-reducing Jacobi-like algorithm for the eigenvalues of non-normal matrices. Journal of Computational and Applied Mathematics, 8(3):203–207, September 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X82900420.

Kaas-Petersen:1989:BPI

[KP89] Christian Kaas-Petersen. Biperiodicity in an isothermal autocatalytic reaction– diffusion system. Journal of Computational and Applied Mathematics, 26(1–2): 187–198, June 1989. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042789901556.

Konstantinov:1980:SLS

[KPC80] M. M. Konstantinov, P. Hr.

Petkov, and N. D. Christov. Synthesis of linear systems with desired equivalent form. Journal of Computational and Applied Mathematics, 6(1):27–35, March 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X80900145.

Kadalbajoo:1986:CSS

[KR86] Mohan K. Kadalbajoo and K. Santhana Raman. Cubic spline solutions of boundary value problems over infinite intervals. Journal of Computational and Applied Mathematics, 15(3):283–291, July 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786902190.

Kahaner:1987:TAR

[KR87] David K. Kahaner and Ottis W. Rechard. TWODQD an adaptive routine for two-dimensional integration. Journal of Computational and Applied Mathematics, 17(1–2):215–234, February 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787900483.

Krishnaiah:1981:ILM

[Kri81] U. Anantha Krishnaiah. Inverse linear multistep methods for the numerical solution of initial value problems of second order differential equations. Journal of Computational and Applied Mathematics, 7(2):111–114, June 1981. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X81900437.

Krishnaiah:1982:AMP

[Kri82] U. Anantha Krishnaiah. Adaptive methods for periodic initial value problems of second order differential equations. Journal of Computational and Applied Mathematics, 8(2):101–104, June 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X82900626.

Kerzman:1986:NCM

[KT86] Norberto Kerzman and Manfred R. Trummer. Numerical conformal mapping via the Szegő kernel. Journal of Computational and Applied Mathematics, 14(1– 2):111–123, February 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042786901330.

Kozek:1988:DMM

[KTS88] Andrzej S. Kozek and Anna Trzmielak-Stanislawska. On dependence of MR- and MMalgorithms upon the value of switching control variable CTR. Journal of Computational and Applied Mathematics, 23(1):109–115, July 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788903354.

Kincaid:1988:BRI

[KY88] David R. Kincaid and David M. Young. A brief review of the ITPACK project. Journal of Computational and Applied Mathematics, 24(1–2):121–127, November 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788903470.

Laforgia:1986:IBF

[Laf86] Andrea Laforgia. Inequalities for Bessel functions. Journal of Computational and Applied Mathematics, 15(1):75–81, May 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786902396.

Lambert:1985:QMC

[Lam85] J. P. Lambert. Quasi-Monte Carlo, low discrepancy sequences, and ergodic transformations. Journal of Computational and Applied Mathematics, 12–13(??):419–423, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
0377042785900354.

Laurie:1985:PEE

[Lau85] Dirk P. Laurie. Practical error estimation in numerical integration. Journal of Computational and Applied Mathematics, 12–13(??):425–431, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042785900366.

Laurie:1987:P

[Lau87] D. P. Laurie. Preface. Journal of Computational and Applied Mathematics, 18(1):1, April 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900513.

Lavery:1984:LCM

[Lav84] John E. Lavery. Local convergence of the method of pseudolinear equations for quasilinear elliptic boundary-value problems. Journal of Computational and Applied Mathematics, 11(1):69–82, August 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900335.

Levrie:1988:NTC

[LB88] Paul Levrie and Adhemar Bultheel. A note on two convergence acceleration methods for ordinary continued fractions. Journal of Computational and Applied Mathematics, 24(3):403–409, December 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903019.

Lee:1986:SBC

[LD86] Young Jack Lee and Edward J. Dudewicz. On selecting the best contender. Journal of Computational and Applied Mathematics, 16 (2):135–143, October 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042786900877.

Lyness:1987:QEE

[LD87] J. N. Lyness and Elise De Doncker-Kapenga. On quadrature error expansions. Part I. Journal of Computational and Applied Mathematics, 17(1–2):131–149, February 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900434.

Liu:1987:SRK

[LDS87] M. Z. Liu, K. Dekker, and M. N. Spijker. Suitability of Runge–Kutta methods. Journal of Computational and Applied Mathematics, 20(??):307–315, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901476.

Lecot:1989:LDS

[Léc89] Christian Lécot. Low discrepancy sequences for solving the Boltzmann equation. Journal of Computational and Applied Mathematics, 25(2):237–249, February 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789900496.

Lee:1986:NBB

[Lee86] D. Lee. A note on bivariate box splines on a k-direction mesh. Journal of Computational and Applied Mathematics, 15(1):117–122, May 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278690244X.

Lembarki:1987:ALP

[Lem87]  A. Lembarki.  Accelera-

tion of limit periodic continued fractions by the T_{+M} transformation. Journal of Computational and Applied Mathematics, 19(1 (supplement 1)):109–116, July 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901786.

Lembarki:1989:KTC

[Lem89] A. Lembarki. On a khovanskii transformation for continued fractions. Journal of Computational and Applied Mathematics, 25(2):125–131, February 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789900435.

Lether:1981:SOG

[Let81] Frank G. Lether. Some observations on Gauss–Legendre quadrature error estimates for analytic functions. Journal of Computational and Applied Mathematics, 7(1): 63–66, March 1981. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0771050X81900097.

Lether:1987:QRA

[Let87]  F. G. Lether.  Quadra-

ture rules for approximating semi-integrals and semiderivatives. Journal of Computational and Applied Mathematics, 17(1–2):115–129, February 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900422.

Levrie:1986:PTG

[Lev86] Paul Levrie. Pringsheim’s theorem for generalized continued fractions. Journal of Computational and Applied Mathematics, 14(3):439–445, March 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786900774.

Levrie:1988:BRF

[Lev88]  P. Levrie.  Book review:

On functions and functional equations: J. Smíal, Adam Hilger, Bristol and Philadelphia 1988, 155 pp., Price: UK 12.50 Pounds, ISBN 0-85274-418-8. Journal of Computational and Applied Mathematics, 24(3):N5, December 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278890310X.

Levrie:1989:PTR

[Lev89] Paul Levrie. Pringsheim’s theorem revisited. Journal of Computational and Applied Mathematics, 25(1):93– 104, January 1989. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042789900782.

Lin:1986:HOT

[Lin86] Avi Lin. High-order threepoint schemes for boundary value problems. II: nonlinear problems. Journal of Computational and Applied Mathematics, 15(3):269–282, July 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786902189.

Liniger:1989:QDF

[Lin89] Werner Liniger. A quasi-direct fast Poisson solver for general regions. Journal of Computational and Applied Mathematics, 28(??):25–47, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789903191.

Lipitakis:1981:NAS

[Lip81] Elias A. Lipitakis. A normalized algorithm for the solution of sparse symmetric seven-diagonal linear systems of semi-bandwidths M and P. Journal of Computational and Applied Mathematics, 7(1):77–84, March 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X81900115.

Lipitakis:1983:DAA

[Lip83a] Elias A. Lipitakis. Direct and approximate algorithmic methods for solving selfadjoint elliptic PDEs in threespace dimensions. Journal of Computational and Applied Mathematics, 9(2):167–176, June 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042783900390.

Lipitakis:1983:NSL

[Lip83b] Elias A. Lipitakis. A normalized sparse linear equations solver. Journal of Computational and Applied Mathematics, 9(3):287–298, September 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042783900213.

Lipitakis:1984:USD

[Lip84]  Elias A. Lipitakis.  The

use of second degree normalized implicit conjugate gradient methods for solving large sparse systems of linear equations. Journal of Computational and Applied Mathematics, 11(1):39–48, August 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704278490030X.

Lazer:1982:MSF

[LLM82] Alan C. Lazer, Anthony W. Leung, and Diego A. Murio. Monotone scheme for finite difference equations concerning steady-state prey-predator interactions. Journal of Computational and Applied Mathematics, 8(4):243–252, December 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X8290047X.

Lal:1982:PSA

[LM82] M. Lal and D. Moffatt. Picard’s successive approximation for non-linear twopoint boundary-value problems. Journal of Computational and Applied Mathematics, 8(4):233–236, December 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X82900456.

Loi:1984:AQA

[LM84] S. L. Loi and A. W. McInnes. An algorithm for the quadratic approximation. Journal of Computational and Applied Mathematics, 11(2):161–174, October 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900189.

Lambert:1986:SDP

[LM86a] F. Lambert and M. Musette.

Solitons from a direct point of view: padeons. Journal of Computational and Applied Mathematics, 15(2):235–249, June 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786900300.

Leung:1986:AMS

[LM86b] Anthony W. Leung and Diego A. Murio. Accelerated monotone scheme for finite difference equations concerning steady-state preypredator interactions. Journal of Computational and Applied Mathematics, 16(3): 333–341, November 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/037704278690004X.

Longman:1984:SAF

[Lon84] I. M. Longman. Some aspects of the finite moment problem. Journal of Computational and Applied Mathematics, 10(2):141–146, April 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042784900505.

Longman:1985:SPS

[Lon85]  I. M. Longman.  The

summation of power series and Fourier series. Journal of Computational and Applied Mathematics, 12– 13(??):447–457, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/037704278590038X.

Langer:1986:CFU

[LQ86] U. Langer and W. Queck. On the convergence factor of Uzawa’s algorithm. Journal of Computational and Applied Mathematics, 15(2):191–202, June 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786900269.

Li:1989:NRL

[LR89] Peiqing Li and A. G. Ramm. Numerical recovery of the layered medium from the surface data. Journal of Computational and Applied Mathematics, 25(3):267–276, May 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789900319.

Laforgia:1988:MRI

[LS88] Andrea Laforgia and Silvana Sismondi. Monotonicity results and inequalities for the gamma and error functions. Journal of Computational and Applied Mathematics, 23(1):25–33, July 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903287.

Lindtner:1989:GOP

[LST89] E. Lindtner, A. Steindl, and H. Troger. Generic one-parameter bifurcations in the motion of a simple robot. Journal of Computational and Applied Mathematics, 26(1–2):199–218, June 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789901568.

Luke:1980:CCP

[Luk80] Yudell L. Luke. Computations of coefficients in the polynomials of Padé approximations by solving systems of linear equations. Journal of Computational and Applied Mathematics, 6(3):213–218, September 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X80900285.

Luke:1982:NEC

[Luk82] Yudell L. Luke. A note on evaluation of coefficients in the polynomials of Padé approximants by solving systems of linear equations. Journal of Computational and Applied Mathematics, 8(2):93–99, June 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X82900614.

Lu:1988:VTM

[LW88] Lin Chao Lu and Yan Wang. A variable transformation method for solving linear absolute-value objectivefunction problems with linear constraints. Journal of Computational and Applied Mathematics, 21(1):111–113, January 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788903925.

Lee:1989:SPS

[LY89] Jenn Hua Lee and Shi Nine Yang. Shape preserving and shape control with interpolating Bézier curves. Journal of Computational and Applied Mathematics, 28(??):269–280, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903397.

Lyness:1985:ISI

[Lyn85]  J. N. Lyness.  Integrat-

ing some infinite oscillating tails. Journal of Computational and Applied Mathematics, 12–13(??):109–117, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278590010X.

Maas:1984:LBI

[Maa84] Christoph Maas. A lower bound for the interval number of a graph. Journal of Computational and Applied Mathematics, 10(1):65–69, February 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900700.

Maas:1985:CIA

[Maa85] Christoph Maas. Computing and interpreting the adjacency spectrum of traffic networks. Journal of Computational and Applied Mathematics, 12–13(??):459–466, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        0377042785900391.

Macleod:1982:CAP

[Mac82]  Allan J. Macleod.  A

comparison of algorithms for polynomial interpolation. Journal of Computational and Applied Mathematics, 8(4):275–277, December 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X82900511. See com-

ments [McN86b].

Maess:1988:PMS

[Mae88] Gerhard Maess. Projection methods solving rectangular systems of linear equations. Journal of Computational and Applied Mathematics, 24(1–2):107–119, November 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788903469.

Magnus:1987:TMT

[Mag87] Alphonse P. Magnus. Toeplitz matrix techniques and convergence of complex weight Padé approximants. Journal of Computational and Applied Mathematics, 19(1 (supplement 1)):23–38, July 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901683.

Mampaey:1989:NTI

[Mam89]  F. Mampaey.  A numer-

ical technique to increase the stability of the ADI method in solidification simulation. Journal of Computational and Applied Mathematics, 28(??):297–308, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903415.

Manicke:1984:USS

[Man84a] Robert L. Manicke. Unique solutions of systems of secondorder nonlinear two-point boundary-value problems. Journal of Computational and Applied Mathematics, 10(2): 175–178, April 1984. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042784900554.

Manoranjan:1984:BSRb

[Man84b] V. S. Manoranjan. Bifurcation studies in reaction–diffusion II. Journal of Computational and Applied Mathematics, 11(3):307–314, December 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
0377042784900050.

Martinez:1986:SSN

[Mar86] JoséMário Martinez. Solving systems of nonlinear equations by means of an accelerated successive orthogonal projections method. Journal of Computational and Applied Mathematics, 16 (2):169–179, October 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042786900907.

Martins:1988:ECR

[Mar88] M. Madalena Martins. Extended convergence regions for the AOR method. Journal of Computational and Applied Mathematics, 24(3): 327–336, December 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042788902944.

Mastin:1987:IFD

[Mas87]  C. Wayne Mastin.  Im-

plicit finite difference methods on composite grids. Journal of Computational and Applied Mathematics, 20(??):317–323, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901488.

Mayo:1986:RMC

[May86] Anita Mayo. Rapid methods for the conformal mapping of multiply connected regions. Journal of Computational and Applied Mathematics, 14(1–2):143–153, February 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786901354.

Magnus:1985:DDL

[MC85] Alphonse Magnus and Régis Chauvaux. On the discretisation of the double layer integral operator for surfaces of revolution. Journal of Computational and Applied Mathematics, 12–13(??):467–473, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042785900408.

McCabe:1981:PFA

[McC81a]  J. H. McCabe.  Perron

fractions: an algorithm for computing the Padé table. Journal of Computational and Applied Mathematics, 7(4):271–275, December 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X81900553.

McCabe:1981:TPP

[McC81b]  J. H. McCabe.  Two-

point Padé approximants and the quotient difference algorithm. Journal of Computational and Applied Mathematics, 7(2):151–153, June 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X81900498.

McCabe:1983:ASC

[McC83] J. H. McCabe. On an asymptotic series and corresponding continued fraction for a gamma function ratio. Journal of Computational and Applied Mathematics, 9(2):125–130, June 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042783900353.

McNamee:1986:CSL

[McN86a] John Michael McNamee. Comparison of spline and Lagrangian interpolation. Journal of Computational and Applied Mathematics, 16 (2):237–240, October 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042786900956.

McNamee:1986:SCP

[McN86b] John Michael McNamee. Some comments on a paper by Macleod on Lagrangian interpolation. Journal of Computational and Applied Mathematics, 16(2):245–246, October 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278690097X.  See

[Mac82].

McNamee:1988:CMT

[McN88] J. M. McNamee. A comparison of methods for terminating polynomial iterations. Journal of Computational and Applied Mathematics, 21(2):239–244, February 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788902725.

Martinez:1986:PSK

[MD86] José Mário Martínez and Raimundo J. B. De Sampaio. Parallel and sequential Kaczmarz methods for solving underdetermined nonlinear equations. Journal of Computational and Applied Mathematics, 15(3):311–321, July 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786902220.

Meinardus:1981:SPG

[Mei81] Günter Meinardus. On the sequence of the powers of a given matrix. Journal of Computational and Applied Mathematics, 7(3):225–226, September 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X81900206. See com-

ment [Ast82].

Meneguette:1987:NBI

[Men87]  M. Meneguette.  A note

on bordered isoclinal matrices. Journal of Computational and Applied Mathematics, 17(3):375–377, March 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787901154.

Mettke:1983:CCH

[Met83] Holger Mettke. Convex cubic HERMITE-spline interpolation. Journal of Computational and Applied Mathematics, 9(3):205–211, September 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042783900134.

Mettke:1984:CCH

[Met84] Holger Mettke. Convex cubic HERMITE-spline interpolation. Journal of Computational and Applied Mathematics, 11(3):377–378, December 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900128.

Mitchell:1980:UPG

[MG80] A. R. Mitchell and D. F.

Griffiths. Upwinding by Petron–Galerkin methods in convection–diffusion problems. Journal of Computational and Applied Mathematics, 6(3):219–228, September 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X80900297.

Marzulli:1989:EGD

[MG89] P. Marzulli and G. Gheri. Estimation of the global discretization error in shooting methods for linear boundary value problems. Journal of Computational and Applied Mathematics, 28(??): 309–314, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042789903427.

Murty:1983:OCC

[MH83] B. S. N. Murty and Asghar Husain. Orthogonality correction in the conjugate-gradient method. Journal of Computational and Applied Mathematics, 9(4):299–304, December 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042783900018.

Miekkala:1989:DIM

[Mie89] Ulla Miekkala. Dynamic iteration methods applied to linear DAE systems. Journal of Computational and Applied Mathematics, 25 (2):133–151, February 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042789900447.

Milner:1989:EEL

[Mil89] Fabio A. Milner. L^{∞}-error estimates for linear elasticity problems. Journal of Computational and Applied Mathematics, 25(3):305–313, May 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789900344.

Mitsui:1987:MVU

[Mit87] Takemoto Mitsui. A modified version of Urabe’s implicit single-step method. Journal of Computational and Applied Mathematics, 20(??):325–332, November 1987. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278790149X.

Miyakoda:1989:IMM

[Miy89] Tsuyako Miyakoda. Iterative methods for multiple zeros of a polynomial by clustering. Journal of Computational and Applied Mathematics, 28(??):315–326, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903439.

Miersemann:1989:CPN

[MM89] Erich Miersemann and Hans D.

Mittelmann. Continuation for parametrized nonlinear variational inequalities. Journal of Computational and Applied Mathematics, 26(1–2):23–34, June 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789901453.

Manoranjan:1984:BSRa

[MMSY84] V. S. Manoranjan, A. R.

Mitchell, B. D. Sleeman, and Kuo Pen Yu. Bifurcation studies in reaction– diffusion. Journal of Computational and Applied Mathematics, 11(1):27–38, August 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042784900293.

Moret:1987:ISI

[MO87] Igor Moret and Pierpaolo Omari. Iterative solution of integral equations by a quasi-Newton method. Journal of Computational and Applied Mathematics, 20(??):333–340, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787901506.

Mori:1985:QFO

[Mor85] Masatake Mori. Quadrature formulas obtained by variable transformation and the DErule. Journal of Computational and Applied Mathematics, 12–13(??):119–130, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042785900111.

Mori:1987:P

[MP87] Masatake Mori and Robert Piessens. Preface. Journal of Computational and Applied Mathematics, 17(1– 2):1, February 1987. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042787900331.

Murio:1988:CMF

[MP88] Diego A. Murio and Jorge R.

Paloschi. Combined mollificationfuture temperatures procedure for solution of inverse heat conduction problem. Journal of Computational and Applied Mathematics, 23(2):235–244, August 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788902828.

Mettke:1982:QSI

[MPN82] Holger Mettke, Eckehard Pfeifer, and Edward Neuman. Quadratic spline interpolation with coinciding interpolation and spline grids. Journal of Computational and Applied Mathematics, 8(1):57–62, March 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X82900080.

Muhlbach:1987:LER

[MR87] G. Mühlbach and L. Reimers.

Linear extrapolation by rational functions, exponentials and logarithmic functions. Journal of Computational and Applied Mathematics, 17(3):329–344, March 1987. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787901099.

Mittelmann:1989:P

[MR89] Hans D. Mittelmann and Dirk Roose. Preface. Journal of Computational and Applied Mathematics, 26(1–2):1–2, June 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704278990143X.

Mahoney:1983:PFE

[MS83] J. F. Mahoney and B. D. Sivazlian. Partial fractions expansion: a review of computational methodology and efficiency. Journal of Computational and Applied Mathematics, 9(3):247–269, September 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042783900183.

Makinson:1986:ISM

[MS86] G. J. Makinson and A. A. Shah. An iterative solution method for solving sparse nonsymmetric linear systems. Journal of Computational and Applied Mathematics, 15(3):339–352, July 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786902256.

Meier:1988:BCG

[MS88] Ulrike Meier and Ahmed Sameh. The behavior of conjugate gradient algorithms on a multivector processor with a hierarchical memory. Journal of Computational and Applied Mathematics, 24(1–2):13–32, November 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278890341X.

Mitchell:1989:NDS

[MS89] A. R. Mitchell and S. W.

Schoombie. Nonlinear diffusion and stable period 2 solutions of a discrete reaction– diffusion model. Journal of Computational and Applied Mathematics, 25(3):363–372, May 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789900393.

Morgan:1989:MRH

[MSW89] Alexander P. Morgan, Andrew Sommese, and Layne T. Watson. Mathematical reduction of a heart dipole model. Journal of Computational and Applied Mathematics, 27(3):407–410, November 1989. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789900253.

Mu:1989:GCS

[Mu89]  Mu Mu.  Global clas-

sical solutions to initialboundary value problems for the potential vorticity equation. Journal of Computational and Applied Mathematics, 28(??):327–338, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789903440.

Murio:1982:EBT

[Mur82] Diego A. Murio. On the estimation of the boundary temperature on a sphere from measurements at its center. Journal of Computational and Applied Mathematics, 8(2):111–119, June 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X8290064X.

Murray:1987:RBT

[Mur87] Daniel M. Murray. The relationship between terminal state constraints and penalties for the discrete-time LQP problem associated with the adjustment of accelerometer data. Journal of Computational and Applied Mathematics, 18(1):83–91, April 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900574.

Murio:1988:PSD

[Mur88] Diego A. Murio. Parameter selection by discrete mollification and the numerical solution of the inverse heat conduction problem. Journal of Computational and Applied Mathematics, 22(1):25–34, April 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788902865.

Maeder:1987:SPM

[MW87] A. J. Maeder and S. A. Wynton. Some parallel methods for polynomial rootfinding. Journal of Computational and Applied Mathematics, 18(1):71–81, April 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900562.

Meek:1989:UCS

[MW89] D. S. Meek and D. J. Walton. The use of Cornu spirals in drawing planar curves of controlled curvature. Journal of Computational and Applied Mathematics, 25(1):69–78, January 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789900769.

Nagurney:1988:GDS

[NA88] Anna Nagurney and Jay Aronson. A general dynamic spatial price equilibrium model: formulation, solution, and computational results. Journal of Computational and Applied Mathematics, 22(2–3):339–357, June 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788904128.

Nakao:1987:SGG

[Nak87] Mitsuhiro T. Nakao. Superconvergence of the gradient of Galerkin approximations for elliptic problems. Journal of Computational and Applied Mathematics, 20(??):341–348, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787901518.

Natterer:1986:EEO

[Nat86] Frank Natterer. Efficient evaluation of oversampled functions. Journal of Computational and Applied Mathematics, 14(3):303–309, March 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786900683.

Nelson:1988:AMD

[Nel88]  Paul Nelson.  Adomian’s

method of decomposition: critical review and examples of failure. Journal of Computational and Applied Mathematics, 23(3):389–393, September 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788900088. See com-

ments [Ado89].

Neuman:1981:MFT

[Neu81] Edward Neuman. Moments and Fourier transforms of Bsplines. Journal of Computational and Applied Mathematics, 7(1):51–62, March 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X81900085.

Nevanlinna:1985:MVV

[Nev85] Olavi Nevanlinna. Matrix valued versions of a result of von Neumann with an application to time discretization. Journal of Computational and Applied Mathematics, 12–13(??):475–489, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704278590041X.

Neta:1984:FMO

[NF84] B. Neta and C. H. Ford. Families of methods for ordinary differential equations based on trigonometric polynomials. Journal of Computational and Applied Mathematics, 10(1):33–38, February 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042784900669.

Niijima:1986:EAD

[Nii86] Koichi Niijima. An error analysis for a difference scheme of exponential type applied to a nonlinear singular perturbation problem without turning points. Journal of Computational and Applied Mathematics, 15(1):93–101, May 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786902414.

Niki:1982:AIMb

[NIS82] Hiroshi Niki, Masatoshi Ikeuchi, and Hideo Sawami. On the accelerated iterative method for estimating the optimum overrelaxation parameter. Journal of Computational and Applied Mathematics, 8(4):257–266, December 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X82900493.

Niznik:1980:NAT

[Niz80] Carol A. Niznik. A numerical analysis of two point boundary value problem algorithms. Journal of Computational and Applied Mathematics, 6(3):201–211, September 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X80900273.

Niznik:1984:MMV

[Niz84] Carol A. Niznik. Min-max vs. max-min flow control algorithms for optimal computer network capacity assignment. Journal of Computational and Applied Mathematics, 11(2):209–224, October 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900219.

Natvig:1986:ECC

[NNOP86] J. Natvig, B. Nour-Omid, and B. N. Parlett. Effect of the CYBER 205 on the choice of method for solving the eigenvalue problem (A − λM)x = 0. Journal of Computational and Applied Mathematics, 15(2):137–159, June 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786900233.

Notay:1988:PAI

[Not88] Yvan Notay. Polynomial acceleration of iterative schemes associated with subproper splittings. Journal of Computational and Applied Mathematics, 24(1–2):153–167, November 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788903500.

Niessner:1985:CBI

[NR85] H. Niessner and M. Ribaut. Condition of boundary integral equations arising from flow computations. Journal of Computational and Applied Mathematics, 12–13(??):491–503, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042785900421.

Neittaanmaki:1982:MLS

[NS82] P. Neittaanmäki and J. Saranen. A modified least squares FE-method for ideal fluid flow problems. Journal of Computational and Applied Mathematics, 8(3):165–169, September 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X82900389.

Niki:1982:AIMa

[NSIO82] Hiroshi Niki, Hideo Sawami, Masatoshi Ikeuchi, and Naotaka Okamoto. The alpha interpolation method for the solution of an eigenvalue problem. Journal of Computational and Applied Mathematics, 8(1):15–19, March 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X8290002X.

Niki:1980:NNS

[NSO80] Hiroshi Niki, Hideo Sawami, and Tetsuya Ondoh. Note on the non-stationary iterative method for the estimation of the optimum overrelaxation parameter. Journal of Computational and Applied Mathematics, 6(2):127–131, June 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X80900066.

Ng-Stynes:1988:NMT

[NSOS88] Mary Jeanne Ng-Stynes, Eugene O’Riordan, and Martin Stynes. Numerical methods for time-dependent convection–diffusion equations. Journal of Computational and Applied Mathematics, 21(3):289–310, March 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788903159.

Noor:1986:NMU

[NT86] Muhammad Aslam Noor and S. I. A. Tirmizi. Numerical methods for unilateral problems. Journal of Computational and Applied Mathematics, 16(3):387–395, November 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786900099.

Oliver:1980:AND

[Oli80a] J. Oliver. An algorithm for numerical differentiation of a function of one real variable. Journal of Computational and Applied Mathematics, 6(2):145–160, June 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X8090008X.

Oliver:1980:ANT

[Oli80b] J. Oliver. On the application of Newbery’s transformation to the Reinsch polynomial evaluation scheme. Journal of Computational and Applied Mathematics, 6(1): 43–52, March 1980. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0771050X80900169.

Oliver:1980:SIP

[Oli80c] J. Oliver. The selection of interpolation points in the numerical approximation of second and third order derivatives. Journal of Computational and Applied Mathematics, 6(3):229–239, September 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X80900303.

Ong:1984:FAS

[Ong84a] Hoon Liong Ong. Fast approximate solution of large scale sparse linear systems. Journal of Computational and Applied Mathematics, 10 (1):45–54, February 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042784900682.

Ong:1984:FES

[Ong84b]  Hoon Liong Ong.  Fi-

nite element solution to the second order elliptic problems. Journal of Computational and Applied Mathematics, 10(1):39–43, February 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042784900670.

Ortiz:1985:NSN

[OP85] E. L. Ortiz and K.-S. Pun. Numerical solution of nonlinear partial differential equations with the Tau method. Journal of Computational and Applied Mathematics, 12– 13(??):511–516, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042785900445.

Orsi:1989:NQR

[Ors89] Annamaria Palamara Orsi. Nested quadrature rules for Cauchy principal value integrals. Journal of Computational and Applied Mathematics, 25(3):251–266, May 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789900307.

ORiordan:1987:UFE

[OS87] Eugene O’Riordan and Martin Stynes. A uniform finite element method for a conservative singularly perturbed problem. Journal of Computational and Applied Mathematics, 18(2):163–174, May 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900148.

Ouellet:1982:GMM

[Oue82] Roch Ouellet. A general model for the measurement of traveller flows. Journal of Computational and Applied Mathematics, 8(1):49–50, March 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X82900067.

Ozis:1989:NVC

[Özi89] Turgut Özis. A novel variation of the coordinate transformation method with perturbation technique for diffusioncontrolled moving boundary problems. Journal of Computational and Applied Mathematics, 25(2):225–235, February 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789900484.

Perez-Acosta:1989:ACD

[PA89] Francisco Pérez-Acosta. An algorithm for the computation of the derivatives of a quotient. Journal of Computational and Applied Mathematics, 25(3):379–385, May 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789900411.

Paardekooper:1989:QCP

[Paa89] M. H. C. Paardekooper. A quadratically convergent parallel Jacobi process for diagonally dominant matrices with distinct eigenvalues. Journal of Computational and Applied Mathematics, 27(1–2):3–16, September 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789903580.

Pandian:1989:PUD

[Pan89] Maharaja C. Pandian. A partial upwind difference scheme for nonlinear parabolic equations. Journal of Computational and Applied Mathematics, 26(3):219–233, July 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789902951.

Papatheodorou:1982:ICB

[Pap82] T. S. Papatheodorou. Inverses for a class of banded matrices and applications to piecewise cubic approximation. Journal of Computational and Applied Mathematics, 8(4):285–288, December 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X82900535.

Papamichael:1989:NCM

[Pap89] N. Papamichael. Numerical conformal mapping onto a rectangle with applications to the solution of Laplacian problems. Journal of Computational and Applied Mathematics, 28(??):63–83, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278990321X.

Paszkowski:1987:RRP

[Pas87]  Stefan Paszkowski.  Re-

currence relations in Padé– Hermite approximation. Journal of Computational and Applied Mathematics, 19(1 (supplement 1)):99–107, July 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901774.

Piessens:1984:ACB

[PB84] Robert Piessens and Maria Branders. Algorithm for the computation of Bessel function integrals. Journal of Computational and Applied Mathematics, 11 (1):119–137, August 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042784900372.

Pescatore:1980:MPS

[PC80] C. Pescatore and R. J. Cole. A method for predicting the stability characteristics of threeterm homogeneous recurrence relations. Journal of Computational and Applied Mathematics, 6(3):197–199, September 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X80900261.

Prince:1981:HOE

[PD81] P. J. Prince and J. R. Dormand. High order embedded Runge–Kutta formulae. Journal of Computational and Applied Mathematics, 7(1):67–75, March 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X81900103.

Phillips:1988:SIS

[PD88] T. N. Phillips and A. R.

Davies. On semi-infinite spectral elements for Poisson problems with re-entrant boundary singularities. Journal of Computational and Applied Mathematics, 21(2):173–188, February 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278890266X.

Pretorius:1987:SGR

[PE87] L. Pretorius and D. Eyre.

Spline-Gauss rules and the Nyström method for solving integral equations in quantum scattering. Journal of Computational and Applied Mathematics, 18(2):235–247, May 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900197.

Petkovic:1982:IMS

[Pet82] M. S. Petković. On an iterative method for simultaneous inclusion of polynomial complex zeros. Journal of Computational and Applied Mathematics, 8(1):51–56, March 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X82900079.

Phelps:1984:STE

[PF84] J. S. Phelps III and M. J. Funcheon. Solutions of a transcendental equation basic to the study of composite media heat transfer. Journal of Computational and Applied Mathematics, 10(2):135–139, April 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042784900499.

Phillips:1987:RSS

[Phi87] Timothy N. Phillips. Relaxation schemes for spectral multigrid methods. Journal of Computational and Applied Mathematics, 18(2):149–162, May 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787900136.

Papatheodorou:1980:CMV

[PJ80] T. S. Papatheodorou and M. E. Jesanis. Callocation methods for Volterra integrodifferential equations with singular kernels. Journal of Computational and Applied Mathematics, 6(1):3–8, March 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X8090011X.

Papamichael:1987:NTC

[PKW87] N. Papamichael, C. A. Kokkinos, and M. K. Warby. Numerical techniques for conformal mapping onto a rectangle. Journal of Computational and Applied Mathematics, 20(??):349–358, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278790152X.

Petkovic:1983:NSI

[PM83] M. S. Petković and G. V. Milovanović. A note on some improvements of the simultaneous methods for determination of polynomial zeros. Journal of Computational and Applied Mathematics, 9(1):65–69, March 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042783900286.

Phillips:1987:ERV

[PMD87] C. Phillips, J. L. Mohamed, and L. M. Delves. Evaluation of the residual vector in global element calculations. Journal of Computational and Applied Mathematics, 18(3):331–346, June 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
0377042787900069.

Patricio:1987:MVS

[PO87] Fernanda Patrício and Paula Oliveira. Multiderivative variable stepsize variable formula methods. Journal of Computational and Applied Mathematics, 19(2):261–267, August 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901944.

Pan:1989:LSM

[PP89] C.-T. Pan and R. J. Plemmons. Least squares modifications with inverse factorizations: Parallel implications. Journal of Computational and Applied Mathematics, 27(1–2):109–127, September 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903634.

Pinar:1987:MIF

[PR87] M. Piñar and V. Ramírez. Matrix interpretation of formal orthogonal polynomials for non-definite functionals. Journal of Computational and Applied Mathematics, 18(3):265–277, June 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
037704278790001X.

Panovsky:1988:FIC

[PR88] J. Panovsky and D. L.

Richardson. A family of implicit Chebyshev methods for the numerical integration of second-order differential equations. Journal of Computational and Applied Mathematics, 23(1):35–51, July 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903299.

Prevost:1983:PTA

[Pre83]  Marc Prevost.  Padé-type

approximants with orthogonal generating polynomials. Journal of Computational and Applied Mathematics, 9(4):333–346, December 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042783900055.

Prevost:1987:CPM

[Pre87]  Marc Prevost.  Calcula-

tion of poles of meromorphic functions with q-d, rs and ϵ-algorithms. Acceleration of these processes. Journal of Computational and Applied Mathematics, 19(1 (supplement 1)):89–98, July 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787901762.

Prevost:1988:SGT

[Pre88]  Marc Prevost.  Stieltjes-

and Geronimus-type polynomials. Journal of Computational and Applied Mathematics, 21(2):133–144, February 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788902634.

Price:1980:NAA

[Pri80] Thomas E. Price, Jr. The numerical approximation of analytic functions in the complex domain. Journal of Computational and Applied Mathematics, 6(3):177–182, September 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X80900236.

Price:1987:PEE

[Pri87] Thomas E. Price, Jr. Pointwise error estimates for interpolation. Journal of Computational and Applied Mathematics, 19(3):389–393, September 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704278790207X.

Petkovic:1986:SIS

[PS86] M. S. Petković and L. V. Stefanović. On some improvements of square root iteration for polynomial complex zeros. Journal of Computational and Applied Mathematics, 15(1):13–25, May 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786902359.

Papamichael:1987:CQS

[PS87] N. Papamichael and M. J.

Soares. Cubic and quintic spline-on-spline interpolation. Journal of Computational and Applied Mathematics, 20(??):359–366, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901531.

Puu:1988:CTA

[Puu88a] Tönu Puu. Catastrophe theory applied to the refraction of traffic. Journal of Computational and Applied Mathematics, 22(2–3):315–318, June 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788904104.

Puu:1988:SSM

[Puu88b] Tönu Puu. Structural stability as a modeling instrument in spatial economics. Journal of Computational and Applied Mathematics, 22(2–3):369–379, June 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788904141.

Potier:1985:SSF

[PV85] C. Potier and C. Vercken. Surface smoothing with finite elements. Journal of Computational and Applied Mathematics, 12–13(??):505–509, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042785900433.

Papamichael:1981:CSM

[PW81a] N. Papamichael and A. J. Worsey. A cubic spline method for the solution of a linear fourth-order twopoint boundary value problem. Journal of Computational and Applied Mathematics, 7(3):187–189, September 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X81900176.

Papamichael:1981:ECI

[PW81b] N. Papamichael and A. J.

Worsey. End conditions for improved cubic spline derivative approximations. Journal of Computational and Applied Mathematics, 7(2):101–109, June 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X81900425.

Papamichael:1984:PTS

[PW84] N. Papamichael and M. K.

Warby. Pole-type singularities and the numerical conformal mapping of doubly-connected domains. Journal of Computational and Applied Mathematics, 10(1):93–106, February 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900736.

Papamichael:1983:DPM

[PWH83] N. Papamichael, M. K. Warby, and D. M. Hough. The determination of the poles of the mapping function and their use in numerical conformal mapping. Journal of Computational and Applied Mathematics, 9(2):155–166, June 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
0377042783900389.

Papamichael:1986:TCP

[PWH86] N. Papamichael, M. K. Warby, and D. M. Hough. The treatment of corner and poletype singularities in numerical conformal mapping techniques. Journal of Computational and Applied Mathematics, 14(1–2):163–191, February 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786901378.

Rabinowitz:1987:NIP

[Rab87a] Philip Rabinowitz. Numerical integration in the presence of an interior singularity. Journal of Computational and Applied Mathematics, 17(1–2):31–41, February 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900367.

Rabinowitz:1987:CCI

[Rab87b] Philip Rabinowitz. On the convergence of closed interpolatory integration rules based on the zeros of Gegenbauer polynomials. Journal of Computational and Applied Mathematics, 17(1–2):43–46, February 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900379.

Rabinowitz:1989:PIB

[Rab89]  Philip Rabinowitz.  Prod-

uct integration based on Hermite–Fejér interpolation. Journal of Computational and Applied Mathematics, 28 (??):85–101, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042789903221.

Rae:1982:LIS

[Rae82]  I. C. Rae.  On Lapla-

cian integrals with singular kernels. Journal of Computational and Applied Mathematics, 8(2):141, June 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X82900717.

Ranga:1987:CPC

[Ran87]  A. Sri Ranga.  Conver-

gence properties of a class of Ĵ-fractions. Journal of Computational and Applied Mathematics, 19(3): 331–342, September 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042787902020.

Reichel:1986:EWB

[RC86] L. Reichel and P. Chapman. Edge waves by boundary collocation. Journal of Computational and Applied Mathematics, 15(1):59–73, May 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786902384.

Roose:1981:AIS

[RD81] Dirk Roose and Elise De Doncker. Automatic integration over a sphere. Journal of Computational and Applied Mathematics, 7(3):203–224, September 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X8190019X.

Rottiers:1987:EEV

[RD87] L. Rottiers and G. De Mey. Eigenfunction expansion versus the boundary element method for thermal problems in microelectronics. Journal of Computational and Applied Mathematics, 20(??):367–372, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787901543.

Reichel:1984:DBC

[Rei84] Lothar Reichel. On the determination of boundary collocation points for solving some problems for the Laplace operator. Journal of Computational and Applied Mathematics, 11(2):175–196, October 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900190.

Reichel:1986:FMS

[Rei86] Lothar Reichel. A fast method for solving certain integral equations of the first kind with application to conformal mapping. Journal of Computational and Applied Mathematics, 14(1–2):125–142, February 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786901342.

Rendl:1988:EAP

[Ren88]  Franz Rendl.  On the

Euclidean assignment problem. Journal of Computational and Applied Mathematics, 23(3):257–265, September 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788900015.

Rijckaert:1980:BRA

[Rij80] M. J. Rijckaert. Book review: Algorithmic graph theory and perfect graphs: Martin C. Golumbic. New York, Academic Press, 1980, 24 cm × 16 cm, 284 pages, price $29.50. Journal of Computational and Applied Mathematics, 6(4):274, December 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X80900364.

Riele:1989:FQS

[RLW89] Herman Te Riele, Walter Lioen, and Dik Winter. Factoring with the quadratic sieve on large vector computers. Journal of Computational and Applied Mathematics, 27(1–2):267–278, September 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903701.

Ronveaux:1989:CRO

[RM89] A. Ronveaux and F. Marcellan. Co-recursive orthogonal polynomials and fourth-order differential equation. Journal of Computational and Applied Mathematics, 25(1):105–109, January 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789900794.

Rockafellar:1988:OCD

[Roc88a] R. T. Rockafellar. Optimization: a case for the development of new mathematical concepts. Journal of Computational and Applied Mathematics, 22(2–3):243–255, June 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788904049.

Rockswold:1988:ITM

[Roc88b] Gary K. Rockswold. Implementation of α-type multistep methods for stiff differential equations. Journal of Computational and Applied Mathematics, 22(1):63–69, April 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788902889.

Rohwer:1987:FOS

[Roh87] C. H. Rohwer. Fast one-sided approximation with spline functions. Journal of Computational and Applied Mathematics, 18(1):93–105, April 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900586.

Rokne:1986:SMC

[Rok86] Jon Rokne. A simple method for computing the values of the derivatives of a rational function. Journal of Computational and Applied Mathematics, 14(3):447–449, March 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786900786.

Rokne:1987:RCC

[Rok87] Jon Rokne. Rational circular complex centered forms. Journal of Computational and Applied Mathematics, 17(3):309–327, March 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042787901087.

Ronning:1986:CTF

[Ron86] Gerd Ronning. On the curvature of the trigamma function. Journal of Computational and Applied Mathematics, 15(3):397–399, July 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786902311.

Roose:1985:ACH

[Roo85]  D. Roose.  An algorithm

for the computation of Hopf bifurcation points in comparison with other methods. Journal of Computational and Applied Mathematics, 12–13(??):517–529, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042785900457.

Rosen:1988:CI

[Ros88]  Robert Rosen.  Complex-

ity and information. Journal of Computational and Applied Mathematics, 22(2–3): 211–218, June 1988. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042788904013.

Rousseeuw:1983:OSP

[Rou83] Peter J. Rousseeuw. Optimal search paths for random variables. Journal of Computational and Applied Mathematics, 9(3):279–286, September 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042783900201.

Rousseeuw:1987:SGA

[Rou87] Peter J. Rousseeuw. Silhouettes: a graphical aid to the interpretation and validation of cluster analysis. Journal of Computational and Applied Mathematics, 20(??):53–65, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787901257.

Rousseeuw:1981:ICG

[RR81] Peter J. Rousseeuw and Elvezio Ronchetti. Influence curves of general statistics. Journal of Computational and Applied Mathematics, 7(3):161–166, September 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X81900139.

Rasmussen:1986:UIB

[RS86] H. Rasmussen and S. D. Salhani. Unsteady interface between two fluids in a porous medium. Journal of Computational and Applied Mathematics, 16(2):159–167, October 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786900890.

Ruotsalainen:1989:CMN

[RS89] K. Ruotsalainen and J. Saranen. On the collocation method for a nonlinear boundary integral equation. Journal of Computational and Applied Mathematics, 28(??):339–348, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903452.

Ruotsalainen:1987:BEM

[Ruo87] K. Ruotsalainen. On the boundary element method with mesh refinement on curves with corners. Journal of Computational and Applied Mathematics, 20(??): 373–378, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042787901555.

Rodrigue:1988:CRE

[RV88] Garry Rodrigue and Richard Varga. Convergence rate estimates for iterative solutions of the biharmonic equation. Journal of Computational and Applied Mathematics, 24(1–2):129–146, November 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903482.

Rudd:1985:EIP

[RWP85] W. G. Rudd, Chien J.

Wang, and Wu Peng. Efficient integration over polytopes. Journal of Computational and Applied Mathematics, 12–13(??):531–540, May 1985. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042785900469.

Saad:1988:PTN

[Saa88a] Youcef Saad. Preconditioning techniques for nonsymmetric and indefinite linear systems. Journal of Computational and Applied Mathematics, 24(1–2):89–105, November 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788903457.

Saari:1988:TIM

[Saa88b] Donald G. Saari. On the types of information and mechanism design. Journal of Computational and Applied Mathematics, 22(2–3):231–242, June 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788904037.

Sablonniere:1983:NFP

[Sab83] Paul Sablonniere. A new family of Padé-type approximants in x. Journal of Computational and Applied Mathematics, 9(4):347–359, December 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
0377042783900067.

Sablonniere:1985:CFE

[Sab85] Paul Sablonniere. Composite finite elements of class C^{k}. Journal of Computational and Applied Mathematics, 12–13(??):541–550, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042785900470.

Sablonniere:1987:CAL

[Sab87] Paul Sablonniere. Convergence acceleration of logarithmic fixed point sequences. Journal of Computational and Applied Mathematics, 19(1 (supplement 1)):55–60, July 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901713.

Salzer:1983:NDG

[Sal83] Herbert E. Salzer. Note on the Dočev–Grosswald asymptotic series for generalized Bessel polynomials. Journal of Computational and Applied Mathematics, 9(2):131–135, June 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042783900365. See er-

rata [Ano84e].

Saridakis:1988:AAA

[Sar88] Yiannis G. Saridakis. An algorithmic approach for the analysis of extrapolated iterative schemes applied to least-squares problems. Journal of Computational and Applied Mathematics, 24(1–2):209–225, November 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788903548.

Saylor:1988:LVI

[Say88] Paul E. Saylor. Leapfrog variants of iterative methods for linear algebraic equations. Journal of Computational and Applied Mathematics, 24(1–2):169–193, November 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788903512.

Shampine:1984:EES

[SB84] Lawrence F. Shampine and Lorraine S. Baca. Error estimators for stiff differential equations. Journal of Computational and Applied Mathematics, 11(2):197–207, October 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042784900207.

Sidi:1988:CSA

[SB88a] Avram Sidi and Jacob Bridger.

Convergence and stability analyses for some vector extrapolation methods in the presence of defective iteration matrices. Journal of Computational and Applied Mathematics, 22(1):35–61, April 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788902877.

Simeonov:1988:ESS

[SB88b] P. S. Simeonov and D. D.

Bainov. Exponential stability of the solutions of the initial-value problem for systems with impulse effect. Journal of Computational and Applied Mathematics, 23(3):353–365, September 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788900064.

Spence:1989:NCP

[SCJ89] A. Spence, K. A. Cliffe, and A. D. Jepson. A note on the calculation of paths of Hopf bifurcations. Journal of Computational and Applied Mathematics, 26(1–2):125–131, June 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        0377042789901519.

Segercrantz:1984:ECP

[Seg84] Jerry Segercrantz. On the existence of conjugate points for a constrained minimization problem. Journal of Computational and Applied Mathematics, 10(3):385–386, June 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042784900463.

Sagara:1984:CMS

[SF84] Nobuko Sagara and Masao Fukushima. A continuation method for solving separable nonlinear least squares problems. Journal of Computational and Applied Mathematics, 10(2):157–161, April 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042784900529.

Sagara:1986:NPC

[SF86] Nobuko Sagara and Masao Fukushima. A new predictor– corrector method for solving unconstrained minimization problems. Journal of Computational and Applied Mathematics, 16(3): 343–354, November 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042786900051.

Sagara:1987:EPC

[SF87] Nobuko Sagara and Masao Fukushima. An efficient predictor–corrector method for solving nonlinear equations. Journal of Computational and Applied Mathematics, 19(3):343–349, September 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787902032.

Sternberg:1988:NMP

[SG88] Robert L. Sternberg and Robert B. Grafton. On a numerical minimization problem for a new class of nonlinear partial differential equations arising in nonspherical geometrical optics and beamforming lenticular antennas II. Journal of Computational and Applied Mathematics, 21(2):213–222, February 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788902695.

Suleiman:1989:TSS

[SG89] M. B. Suleiman and C. W.

Gear. Treating a single, stiff, second-order ODE directly. Journal of Computational and Applied Mathematics, 27(3):331–348, November 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789900216.

Sternberg:1985:NMP

[SGC85] Robert L. Sternberg, Robert B. Grafton, and Donald R. Childs. On a numerical minimization problem for a new class of nonlinear partial differential equations arising in nonspherical geometrical optics and beamforming lenticular antennas. Journal of Computational and Applied Mathematics, 12–13(??):591–607, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042785900512.

Schmidt:1987:QRE

[SH87] Jochen W. Schmidt and Walter Hess. Quadratic and related exponential splines in shape preserving interpolation. Journal of Computational and Applied Mathematics, 18(3):321–329, June 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787900057.

Shampine:1981:EUI

[Sha81] Lawrence F. Shampine. Efficient use of implicit formulas with predictor–corrector error estimate. Journal of Computational and Applied Mathematics, 7(1):33–35, March 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X8190005X.

Shanks:1984:FRT

[Sha84] J. A. Shanks. On forming the Romberg table. Journal of Computational and Applied Mathematics, 11(3): 343–351, December 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042784900098.

Shampine:1986:SSN

[Sha86] Lawrence F. Shampine. Solution of structured non-stiff ODEs. Journal of Computational and Applied Mathematics, 15(3):293–300, July 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786902207.

Shampine:1987:CSS

[Sha87] L. F. Shampine. Control of step size and order in extrapolation codes. Journal of Computational and Applied Mathematics, 18(1):3–16, April 1987. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900525.

Sharma:1988:ISB

[Sha88a] Rakesh K. Sharma. Iterative solutions to boundaryvalue differential equations involving reflection of the argument. Journal of Computational and Applied Mathematics, 24(3):319–326, December 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788902932.

Sharma:1988:NSR

[Sha88b]  Rakesh K. Sharma.  Nu-

merical solutions to radial form of transistor equations. Journal of Computational and Applied Mathematics, 21(3):377–381, March 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903214.

Shoufu:1989:SCG

[Sho89] Li Shoufu. Stability and B-convergence of general linear methods. Journal of Computational and Applied Mathematics, 28(??): 281–296, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042789903403.

Sidi:1980:SAT

[Sid80] Avram Sidi. Some aspects of two-point Padé approximants. Journal of Computational and Applied Mathematics, 6(1):9–17, March 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X80900121.

Sidi:1981:NMD

[Sid81] Avram Sidi. A new method for deriving Padé approximants for some hypergeometric functions. Journal of Computational and Applied Mathematics, 7(1):37–40, March 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X81900061.

Sidi:1987:EMD

[Sid87] Avram Sidi. Extrapolation methods for divergent oscillatory infinite integrals that are defined in the sense of summability. Journal of Computational and Applied Mathematics, 17(1–2):105–114, February 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        0377042787900410.

Sidi:1988:EVP

[Sid88]  Avram Sidi.  Extrapola-

tion vs. projection methods for linear systems of equations. Journal of Computational and Applied Mathematics, 22(1):71–88, April 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788902890.

Siemaszko:1980:BCF

[Sie80] Wojciech Siemaszko. Branched continued fractions for double power series. Journal of Computational and Applied Mathematics, 6(2):121–125, June 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X80900054.

Siemaszko:1983:TTB

[Sie83] Wojciech Siemaszko. Thieletype branched continued fractions for two-variable functions. Journal of Computational and Applied Mathematics, 9(2):137–153, June 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042783900377.

Shenton:1989:Fa

[SK89a] L. R. Shenton and A. W.

Kemp. An S-fraction and ln^{2}(1 + x). Journal of Computational and Applied Mathematics, 25(1):121–124, January 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789900812.

Shenton:1989:Fb

[SK89b] L. R. Shenton and A. W.

Kemp. An S-fraction and ln^{2}(1 + x). Journal of Computational and Applied Mathematics, 26(3):367–370, July 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903087.

Santos:1986:CCS

[SL86] Newton R. Santos and Odelar L. Linhares. Convergence of Chebyshev semi-iterative methods. Journal of Computational and Applied Mathematics, 16(1):59–68, September 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786901731.

Sloan:1985:LMM

[Slo85] Ian H. Sloan. Lattice methods for multiple integration. Journal of Computational and Applied Mathematics, 12–13(??):131–143, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042785900123.

Sloan:1986:NSF

[Slo86]  D. M. Sloan.  Numerical

solution of a free boundary problem by continuation. Journal of Computational and Applied Mathematics, 14(3):279–288, March 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278690066X.

Smiley:1987:LDQ

[Smi87a]  M. W. Smiley.  Low-

dimensional qualitative and numerical approximation of bifurcation in a semilinear elliptic problem. Journal of Computational and Applied Mathematics, 19(3): 301–312, September 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042787902007.

Smiley:1987:NSS

[Smi87b] M. W. Smiley. A numerical study of spontaneous bifurcation. Journal of Computational and Applied Mathematics, 19(2):179–188, August 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901889.

Smyth:1985:ARM

[Smy85] W. F. Smyth. Algorithms for the reduction of matrix bandwidth and profile. Journal of Computational and Applied Mathematics, 12–13(??):551–561, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042785900482.

Snyman:1987:VMA

[Sny87] I. M. Snyman. A variational method applied to a linear hyperbolic initial boundary value problem. Journal of Computational and Applied Mathematics, 17(3):345–354, March 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901105.

Snyman:1988:PFM

[Sny88] J. A. Snyman. A parameterfree multiplier method for constrained minimization problems. Journal of Computational and Applied Mathematics, 23(2):155–168, August 1988. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788902774.

Sloan:1987:MIB

[SO87] I. H. Sloan and T. R. Osborn. Multiple integration over bounded and unbounded regions. Journal of Computational and Applied Mathematics, 17(1–2):181–196, February 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        037704278790046X.

Sagara:1982:III

[SOA82] Nobuko Sagara, Toshio Odanaka, and Tsutomu Arimizu. Invariant imbedding and the identification of aquifer parameters. Journal of Computational and Applied Mathematics, 8(3):209–212, September 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X82900432.

Soh:1988:FSB

[Soh88] B. C. Soh. Fourier series of B-splines. Journal of Computational and Applied Mathematics, 21(1):125–127, January 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
0377042788903950.

Soh:1989:DAP

[Soh89] B. C. Soh. On the degree of approximation to periodic functions by a trigonometric spline convolution operator. Journal of Computational and Applied Mathematics, 26(3):251–255, July 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789902975.

Soler:1987:CMN

[Sol87]  J. Soler.  On cubature

with a minimal number of lines. Journal of Computational and Applied Mathematics, 19(2):223–230, August 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901919.

Solak:1989:QFE

[Sol89] W. Solak. On quadrature formulas with end corrections. Journal of Computational and Applied Mathematics, 25(3):373–377, May 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278990040X.

Sommeijer:1987:NDI

[Som87] B. P. Sommeijer. A note on a diagonally implicit Runge– Kutta–Nyström method. Journal of Computational and Applied Mathematics, 19(3): 395–399, September 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042787902081.

Song:1988:RFE

[Son88a] Iickho Song. A recursive formula for even order harmonic series. Journal of Computational and Applied Mathematics, 21(2):251–256, February 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788902749.

Song:1988:FEM

[Son88b] Jiang Song. A finite element method for equations of one-dimensional nonlinear thermoelasticity. Journal of Computational and Applied Mathematics, 21(1):1–16, January 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788903834.

Sonnevend:1989:ANA

[Son89]  György Sonnevend.  Ap-

plications of the notion of analytic center in approximation (estimation) problems. Journal of Computational and Applied Mathematics, 28(??):349–358, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903464.

Sottas:1984:RRK

[Sot84] G. Sottas. Rational Runge– Kutta methods are not suitable for stiff systems of ODEs. Journal of Computational and Applied Mathematics, 10(2):169–174, April 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900542.

Spijker:1985:FCI

[Spi85]  M. N. Spijker.  Feasibil-

ity and contractivity in implicit Runge–Kutta methods. Journal of Computational and Applied Mathematics, 12–13(??):563–578, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042785900494.

Spoelstra:1989:MOA

[Spo89] J. Spoelstra. The modelling of oil agglomeration of coal fines. Journal of Computational and Applied Mathematics, 28(??):359–366, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789903476.

Schuck:1988:NSH

[SR88] P. L. Schuck and H. Rasmussen. Numerical solution of a heated subsidence mound problem in a porous medium. Journal of Computational and Applied Mathematics, 24(3):293–318, December 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788902920.

Surla:1988:SSP

[SS88] K. Surla and M. Stojanović. Solving singularly perturbed boundary-value problems by spline in tension. Journal of Computational and Applied Mathematics, 24(3): 355–363, December 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/037704278890297X.

Sugiura:1989:CHO

[SS89a] Hiroshi Sugiura and Tetsuya Sakurai. On the construction of high-order integration formulae for the adaptive quadrature method.

Journal of Computational and Applied Mathematics, 28(??):367–381, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903488.

Swarztrauber:1989:VPM

[SS89b] Paul N. Swarztrauber and Roland A. Sweet. Vector and parallel methods for the direct solution of Poisson’s equation. Journal of Computational and Applied Mathematics, 27(1–2):241–263, September 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903695.

Slivnik:1986:NMS

[ST86] Tomaz Slivnik and Gabrijel Tomsic. A numerical method for the solution of twoparameter eigenvalue problems. Journal of Computational and Applied Mathematics, 15(1):109–115, May 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786902438.

Saxena:1987:ISD

[ST87] R. B. Saxena and H. C. Tripathi. (0, 2, 3) and (0, 1,

3) interpolation by six degree splines. Journal of Computational and Applied Mathematics, 18(3):395–401, June 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787900100.

Steiner:1985:AMS

[Ste85] George Steiner. An algorithm for minimizing setups in precedence constrained scheduling. Journal of Computational and Applied Mathematics, 12–13(??):579–589, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042785900500.

Steinberg:1987:CEC

[Ste87] J. Steinberg. Computation of expansion coefficients in a singular harmonic problem. Journal of Computational and Applied Mathematics, 18(2):213–220, May 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787900173.

Spivack:1989:SSS

[SU89] M. Spivack and B. J. Uscinski. The split-step solution in random wave propagation. Journal of Computational and Applied Mathematics, 27(3):349–361, November 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789900228.

Sugihara:1987:CFW

[Sug87a] Masaaki Sugihara. A class of functions for which the trapezoidal rule gives the exact value of integral over the infinite interval. Journal of Computational and Applied Mathematics, 20(??): 387–392, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042787901579.

Sugihara:1987:MGM

[Sug87b] Masaaki Sugihara. Method of good matrices for multidimensional numerical integrations — an extension of the method of good lattice points. Journal of Computational and Applied Mathematics, 17(1–2):197–213, February 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900471.

Sugihara:1987:MNI

[Sug87c] Masaaki Sugihara. Methods of numerical integration of oscillatory functions by the DE-formula with the Richardson extrapolation. Journal of Computational and Applied Mathematics, 17(1–2):47–68, February 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787900380.

Spoelstra:1987:MSN

[SV87] J. Spoelstra and D. J. Van Wyk. A method of solution for a non-linear diffusion model and for computing the parameters in the model. Journal of Computational and Applied Mathematics, 20(??):379–385, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787901567.

Schlichting:1989:SBB

[SvdV89] J. J. F. M. Schlichting and H. A. van der Vorst. Solving 3D block bidiagonal linear systems on vector computers. Journal of Computational and Applied Mathematics, 27(1–2):323–330, September 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789903737.

Shivakumar:1988:IMT

[SW88] P. N. Shivakumar and J. J.

Williams. An iterative method with truncation for infinite linear systems. Journal of Computational and Applied Mathematics, 24(1–2): 199–207, November 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042788903536.

Schonauer:1989:EVP

[SW89] W. Schönauer and R. Weiß.

Efficient vectorizable PDE solvers. Journal of Computational and Applied Mathematics, 27(1–2):279–297, September 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903713.

Szulc:1989:CNP

[Szu89] Tomasz Szulc. Criteria for the nonsingularity of partitioned matrices. Journal of Computational and Applied Mathematics, 26(3):289–295, July 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903014.

Thandapani:1988:SSS

[TBB88] E. Thandapani, K. Balachandran, and G. Balasubramanian. Series solutions of some nonlinear differential equations. Journal of Computational and Applied Mathematics, 23(1):103–107, July 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788903342.

Temme:1981:ECH

[Tem81] N. M. Temme. On the expansion of confluent hypergeometric functions in terms of Bessel functions. Journal of Computational and Applied Mathematics, 7(1):27–32, March 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X81900048.

Temme:1985:CIE

[Tem85]  N. M. Temme.  A con-

volution integral equation solved by Laplace transformations. Journal of Computational and Applied Mathematics, 12–13(??):609–613, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042785900524.

Tempelmeier:1988:NPC

[Tem88] Uwe Tempelmeier. A new proof of the cross-rule for the ϵ-algorithm based on Schur-complements. Journal of Computational and Applied Mathematics, 21(1):

55–61, January 1988. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042788903871.

TenThijeBoonkkamp:1987:OEH

[Ten87] J. H. M. Ten Thije Boonkkamp.

The odd-even hopscotch pressure correction scheme for the incompressible Navier– Stokes equations. Journal of Computational and Applied Mathematics, 20(??): 393–402, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042787901580.

Tezer:1989:NEO

[Tez89] Münevver Tezer. On the numerical evaluation of an oscillating infinite series. Journal of Computational and Applied Mathematics, 28(??):383–390, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278990349X.

Tanaka:1988:GCS

[TFI88] Yoshihiro Tanaka, Masao Fukushima, and Toshihide Ibaraki. A globally convergent SQP method for semiinfinite nonlinear optimization. Journal of Computational and Applied Mathematics, 23(2):141–153, August 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788902762.

Thompson:1982:NEC

[Tho82a]  S. Thompson.  The non-

existence of certain A_{0}- stable Adams-type correctors. Journal of Computational and Applied Mathematics, 8(3):155–157, September 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X82900365.

Thompson:1982:SSF

[Tho82b] S. Thompson. Stiffly stable fourth order Adams-type methods. Journal of Computational and Applied Mathematics, 8(4):253–256, December 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X82900481.

Torelli:1989:SNM

[Tor89] Lucio Torelli. Stability of numerical methods for delay differential equations. Journal of Computational and Applied Mathematics, 25(1): 15–26, January 1989. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/037704278990071X.

Trefethen:1986:P

[Tre86] Lloyd N. Trefethen. Preface. Journal of Computational and Applied Mathematics, 14(1–2):1–6, February 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786901275.

Thrasher:1986:EPM

[TS86] Richard Thrasher and Kamy Sepehrnoori. On equidistributing principles in moving finite element methods. Journal of Computational and Applied Mathematics, 16(3):309–318, November 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786900026.

Twizell:1987:MMN

[TT87] E. H. Twizell and S. I. A. Tirmizi. Multiderivative methods for non-linear secondorder boundary value problems. Journal of Computational and Applied Mathematics, 17(3):299–307, March 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901075.

Ta-tsien:1989:CNL

[Tt89] Li Ta-tsien. A class of nonlocal boundary value problems for partial differential equations and its applications in numerical analysis. Journal of Computational and Applied Mathematics, 28(??):49–62, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903208.

TunesDosSantos:1987:PSP

[Tun87] Lucio Tunes Dos Santos. A parallel subgradient projections method for the convex feasibility problem. Journal of Computational and Applied Mathematics, 18(3):307–320, June 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900045.

Trefethen:1986:CMS

[TW86] Lloyd N. Trefethen and Ruth J. Williams. Conformal mapping solution of Laplace’s equation on a polygon with oblique derivative boundary conditions. Journal of Computational and Applied Mathematics, 14(1–2):227–249, February 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278690141X.

Twizell:1986:PLA

[Twi86] E. H. Twizell. Phase-lag analysis for a family of twostep methods for second order periodic initial value problems. Journal of Computational and Applied Mathematics, 15(2):261–263, June 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786900336.

Usmani:1980:SSA

[Usm80] Riaz A. Usmani. Smooth spline approximations for the solution of a boundary value problem with engineering applications. Journal of Computational and Applied Mathematics, 6(2):93–98, June 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X80900029.

VanderBorght:1980:NSN

[Van80] R. Van der Borght. A numerical study of non-linear convection in a compressible medium. Journal of Computational and Applied Mathematics, 6(4):283–294, December 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X80900388.

VanAssche:1985:SRA

[Van85] Walter Van Assche. Some results on the asymptotic distribution of the zeros of orthogonal polynomials. Journal of Computational and Applied Mathematics, 12– 13(??):615–623, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042785900536.

VanIseghem:1987:VOR

[Van87a]  J. Van Iseghem.  Vector

orthogonal relations. Vector QD-algorithm. Journal of Computational and Applied Mathematics, 19(1 (supplement 1)):141–150, July 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787901828.

VanSiclen:1987:EAO

[Van87b] C. Dew. Van Siclen. Estimating atomic 1s orbital radii. Journal of Computational and Applied Mathematics, 19(2):283–286, August 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278790197X.

VanVeldhuizen:1988:NAR

[Van88]  M. Van Veldhuizen.  On

the numerical approximation of the rotation number. Journal of Computational and Applied Mathematics, 21(2):203–212, February 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788902683.

VanGestel:1989:BRP

[Van89a] E. Van Gestel. Book review:

Program correctness over abstract data types, with errorstate semantics: J. V. Tucker and J. I. Zucker, CWI Monographs, Vol. 6, North-Holland, Amsterdam, 1988, 212 pages, Price: U.S. $58.00/Dfl. 110.00, ISBN 0-444-70340-3. Journal of Computational and Applied Mathematics, 26(3):N2, July 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903129.

VanHuffel:1989:ECT

[Van89b] Sabine Van Huffel. The extended classical total least squares algorithm. Journal of Computational and Applied Mathematics, 25 (1):111–119, January 1989.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042789900800.

VanIseghem:1989:CVA

[Van89c] Jeannette Van Iseghem. Convergence of the vector QDalgorithm. Zeros of vector orthogonal polynomials. Journal of Computational and Applied Mathematics, 25(1):33–46, January 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789900733.

VonHohenbalken:1987:LDM

[VCL87] Balder Von Hohenbalken, Bruce L. Clarke, and James E. Lewis. Least distance methods for the frame of homogeneous equation systems. Journal of Computational and Applied Mathematics, 19(2):231–241, August 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787901920.

vanderCruyssen:1982:DAC

[vdC82] P. van der Cruyssen. Discussion of algorithms for the computation of generalized continued fractions. Journal of Computational and Applied Mathematics, 8(3):179–186, September 1982. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X82900407.

VanDyck:1980:MSE

[VDD80] D. Van Dyck, R. De Ridder, and J. De Sitter. The most storage economical Runge– Kutta methods for the solution of large systems of coupled first-order differential equations. Journal of Computational and Applied Mathematics, 6(1):83–85, March 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X80900212.

vanderHouwen:1983:FOA

[vdHdV83] P. J. van der Houwen and H. B. de Vries. A fourth order ADI method for semidiscrete parabolic equations. Journal of Computational and Applied Mathematics, 9(1): 41–63, March 1983. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042783900274.

vanderHouwen:1984:SLM

[vdHS84] P. J. van der Houwen and B. P. Sommeijer. Stability in linear multistep methods for pure delay equations. Journal of Computational and Applied Mathematics, 10(1):55–63, February 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042784900694.

vanderHouwen:1986:RDH

[vdHS86] P. J. van der Houwen and B. P. Sommeijer. Reduction of dispersion in hyperbolic difference schemes by adapting the space discretization. Journal of Computational and Applied Mathematics, 16(2):203–214, October 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786900920.

vanderVorst:1987:ISM

[vdV87] H. A. van der Vorst. An iterative solution method for solving f(A)x = b, using Krylov subspace information obtained for the symmetric positive definite matrix A. Journal of Computational and Applied Mathematics, 18(2):249–263, May 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787900203.

vanderVorst:1988:CGT

[vdVD88] Henk A. van der Vorst and Kees Dekker. Conjugate gradient type methods and preconditioning. Journal of Computational and Applied Mathematics, 24(1– 2):73–87, November 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042788903445.

VandenBerghe:1989:NMS

[VFD89] G. Vanden Berghe, V. Fack, and H. E. De Meyer. Numerical methods for solving radial Schrödinger equations. Journal of Computational and Applied Mathematics, 28(??):391–401, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903506.

vanLaarhoven:1988:CLF

[vLK88] Peter J. M. van Laarhoven and Ton A. C. M. Kalker. On the computation of Lauricella functions of the fourth kind. Journal of Computational and Applied Mathematics, 21(3):369–375, March 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903202.

Volk:1988:IGM

[Vol88] Wolfgang Volk. The iterated Galerkin method for linear integro-differential equations. Journal of Computational and Applied Mathematics, 21(1):63–74, January 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788903883.

VonHohenbalken:1988:ASO

[Von88] Balder Von Hohenbalken. An algorithm for set operations on nonconvex polygons (with an application to census tracts). Journal of Computational and Applied Mathematics, 23(1):11–24, July 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788903275.

VanDerHouwen:1985:HOD

[VS85] P. J. Van Der Houwen and B. S. Sommeijer. High order difference schemes with reduced dispersion for hyperbolic differential equations. Journal of Computational and Applied Mathematics, 12–13(??):145–161, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042785900135.

vanSlijpe:1986:RWT

[vS86] A. R. D. van Slijpe. Random walks on the triangular prism and other vertextransitive graphs. Journal of Computational and Applied Mathematics, 15(3):383–394, July 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786902293.

Venkaiah:1988:CMS

[VS88] V. Ch Venkaiah and S. K.

Sen. Computing a matrix symmetrizer exactly using modified multiple modulus residue arithmetic. Journal of Computational and Applied Mathematics, 21(1):27–40, January 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903858.

VanDerHouwen:1989:NEP

[VSV89] P. J. Van Der Houwen, B. P.

Sommeijer, and P. A. Van Mourik. Note on explicit parallel multistep Runge– Kutta methods. Journal of Computational and Applied Mathematics, 27(3): 411–420, November 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042789900265.

Vulanovic:1989:FDS

[Vul89] R. Vulanović. Finite-difference schemes for quasilinear singular perturbation problems. Journal of Computational and Applied Mathematics, 26(3):345–365, July 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903075.

VanHuffel:1988:PTL

[VV88] Sabine Van Huffel and Joos Vandewalle. The partial total least squares algorithm. Journal of Computational and Applied Mathematics, 21(3):333–341, March 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903172.

VanHuffel:1987:ERA

[VVH87] Sabine Van Huffel, Joos Vandewalle, and Ann Haegemans. An efficient and reliable algorithm for computing the singular subspace of a matrix, associated with its smallest singular values. Journal of Computational and Applied Mathematics, 19(3):313–330, September 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787902019.

Waadeland:1987:DCF

[Waa87a] Haakon Waadeland. Derivatives of continued fractions with applications to hypergeometric functions. Journal of Computational and Applied Mathematics, 19(1 (supplement 1)):161–169, July 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901841.

Waadeland:1987:LAC

[Waa87b]  Haakon Waadeland.  Lin-

ear approximations to continued fractions K(z_{n}/1). Journal of Computational and Applied Mathematics, 20(??):403–415, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901592.

Wadia:1980:NEH

[Wad80] Aspi Rustom Wadia. Note on the eigensolution of a homogeneous equation with semiinfinite domain. Journal of Computational and Applied Mathematics, 6(2):161–165, June 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X80900091.

Wallin:1987:DPM

[Wal87] Hans Wallin. The divergence problem for multipoint Padé approximants of meromorphic functions. Journal of Computational and Applied Mathematics, 19(1 (supplement 1)):61–65, July 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787901725.

Walz:1988:CME

[Wal88]  Guido Walz.  Comput-

ing the matrix exponential and other matrix functions. Journal of Computational and Applied Mathematics, 21(1):119–123, January 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903949.

Wambecq:1980:SEA

[Wam80]  A. Wambecq.  Solution

of the equations associated with rational Runge–Kutta methods of orders up to four. Journal of Computational and Applied Mathematics, 6(4):275–281, December 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X80900376.

Wang:1985:DBS

[Wan85] Ren-Hong Wang. The dimension and basis of spaces of multivariate splines. Journal of Computational and Applied Mathematics, 12– 13(??):163–177, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042785900147.

Watson:1981:NSP

[Wat81] Layne T. Watson. Numerical study of porous channel flow in a rotating system by a homotopy method. Journal of Computational and Applied Mathematics, 7(1):21–26, March 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X81900036.

Watts:1983:SSS

[Wat83] H. A. Watts. Starting step size for an ODE solver. Journal of Computational and Applied Mathematics, 9(2):177–191, June 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042783900407.

Wegmann:1986:IMCa

[Weg86a] Rudolf Wegmann. An iterative method for conformal mapping. Journal of Computational and Applied Mathematics, 14(1–2):7–18, February 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786901287.

Wegmann:1986:IMCb

[Weg86b] Rudolf Wegmann. An iterative method for the conformal mapping of doubly connected regions. Journal of Computational and Applied Mathematics, 14(1–2):79–98, February 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786901317.

Wegmann:1988:DRH

[Weg88]  Rudolf Wegmann.  Dis-

crete Riemann–Hilbert problems, interpolation of simply closed curves, and numerical conformal mapping. Journal of Computational and Applied Mathematics, 23(3):323–352, September 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788900052.

Weidlich:1988:SMT

[Wei88] Wolfgang Weidlich. Stochastic migration theory and migratory phase transitions.

Journal of Computational and Applied Mathematics, 22 (2–3):359–368, June 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/037704278890413X.

Wise:1981:NAC

[WG81] Gary L. Wise and Neal C.

Gallagher. A novel approach for the computation of orthonormal polynomial expansions. Journal of Computational and Applied Mathematics, 7(3):157–160, September 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X81900127.

Wheeler:1985:ASW

[Whe85] Frederick P. Wheeler. Aspects of the simulation of wave propagation through a clear turbulent medium. Journal of Computational and Applied Mathematics, 12–13(??):625–633, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042785900548.

Wilders:1986:MEM

[Wil86]  P. Wilders.  The modi-

fied equation as a model of local errors in convective schemes. Journal of Computational and Applied Mathematics, 16(1):69–77, September 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786901743.

Wimp:1987:SPO

[WKGC87] Jet Wimp, Bob Kline, Andrew Galardi, and David Colton. Some preliminary observations on an algorithm for the computation of moment integrals. Journal of Computational and Applied Mathematics, 19(1 (supplement 1)):117–124, July 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787901798.

Welsh:1980:MBVa

[WO80a] Wayne Welsh and Takeo Ojika. Multipoint boundary value problems with discontinuities. I. Algorithms and applications. Journal of Computational and Applied Mathematics, 6(2):133–143, June 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0771050X80900078.

Welsh:1980:MBVb

[WO80b] Wayne Welsh and Takeo Ojika. Multipoint boundary value problems with discontinuities II. Convergence of the initial value adjusting method. Journal of Computational and Applied Mathematics, 6(3):183–187, September 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X80900248.

Wille:1986:CCC

[WP86] L. Wille and P. Phariseau.

On the computation of a certain class of Hill determinants. Journal of Computational and Applied Mathematics, 15(1):83–91, May 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786902402.

Wu:1989:DDT

[WW89] Changhui Wu and Song Wang.

Domain decomposition technique for the continuity equations of semiconductor device models. Journal of Computational and Applied Mathematics, 28(??):403–412, December 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903518.

Xinmin:1988:SMS

[Xin88]  Xiang Xinmin.  Spectral

method for solving the system of equations of Schrödinger– Klein–Gordon field. Journal of Computational and Applied Mathematics, 21 (2):161–171, February 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042788902658.

Yamamoto:1985:UDS

[Yam85] T. Yamamoto. A unified derivation of several error bounds for Newton’s process. Journal of Computational and Applied Mathematics, 12–13(??):179–191, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042785900159.

Yamamoto:1988:MTH

[Yam88]  Tetsuro Yamamoto.  On

the method of tangent hyperbolas in Banach spaces. Journal of Computational and Applied Mathematics, 21(1):75–86, January 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042788903895.

Yamada:1985:NAA

[YaOIN85] Seiji Yamada, Ichi ann Ohsaki, Masatoshi Ikeuchi, and Hiroshi Niki. Non-adaptive and adaptive SAOR–CG algorithms. Journal of Computational and Applied Mathematics, 12–13(??):635–650, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704278590055X.

Yeyios:1988:EAO

[Yey88] A. K. Yeyios. An extension of the accelerated overrelaxation (AOR) theory for linear systems with Hermitian matrix. Journal of Computational and Applied Mathematics, 24(1–2):147–151, November 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903494.

Yeyios:1989:NCC

[Yey89] A. K. Yeyios. A necessary condition for the convergence of the accelerated overrelaxation (AOR) method. Journal of Computational and Applied Mathematics, 26(3):371–373, July 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903099.

Ypma:1986:RNL

[Ypm86] T. J. Ypma. Relaxed Newtonlike methods for stiff differential systems. Journal of Computational and Applied Mathematics, 16(1):95–103, September 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042786901767.

Ypma:1987:EES

[Ypm87]  T. J. Ypma.  Efficient

estimation of sparse Jacobian matrices by differences. Journal of Computational and Applied Mathematics, 18(1):17–28, April 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787900537.

Zabusky:1988:CSV

[Zab88] Norman J. Zabusky. Computational synergetics: visualization and vortex dynamics. Journal of Computational and Applied Mathematics, 22(2–3):285–295, June 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042788904074.

Zemach:1986:CMF

[Zem86] Charles Zemach. A conformal map formula for difficult cases. Journal of Computational and Applied Mathematics, 14(1–2):207–215, February 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042786901391.

Zinober:1980:NSL

[ZH80] A. S. I. Zinober and E. Huntley. The numerical solution of linear time-dependent partial differential equations by the Laplace transform and finite difference approximations. Journal of Computational and Applied Mathematics, 6(4):253–258, December 1980. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0771050X80900339.

Zochowski:1989:IMV

[ZH89] A. Zochowski and P. Holnicki.

Interactive method of variational grid generation. Journal of Computational and Applied Mathematics, 26(3):281–287, July 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042789903002.

Zietak:1984:CAR

[Zie84] K. Zietak. The Chebyshev approximation of a rectangular matrix by matrices of smaller rank as the limit of l_{p}-approximation. Journal of Computational and Applied Mathematics, 11(3): 297–305, December 1984.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0377042784900049.

Zlatev:1985:MMS

[Zla85] Zahari Zlatev. Mathematical model for studying the sulphur pollution over Europe. Journal of Computational and Applied Mathematics, 12–13(??):651–666, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042785900561.

Zlatev:1987:SAE

[Zla87]  Zahari Zlatev.  A survey

of the advances in the exploitation of the sparsity in the solution of large problems. Journal of Computational and Applied Mathematics, 20(??):83–105, November 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042787901270.

Zamani:1984:FEF

[ZN84] Nader Zamani and Farhad Nafeiy. Finite element formulation of thermal boundary layers. Journal of Computational and Applied Mathematics, 11(1):11–26, August 1984. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042784900281.